The L^{A}TEX2HTML Translator

January 2, 2021

This document accompanies L^{A}TEX2HTML version 2021.

History Nikos Drakos’ original manuscript was updated for version v96.1 by Herb Swan and converted for L^{A}TEX 2ε by Michel Goossens. Extensive revisions were made by Ross Moore for v96.1 rev-f, incorporating also suggestions from Michel Goossens. Another major revision was required to adequately describe the new features made possible with HTML 3.2 , and recent developments in image-generation and macro-handling. This work was done by Ross Moore, as were most of the revisions for v98.1, v98.2 and v99.1.

Portability for non-Unix systems has been achieved due to work done mainly by Marek Rouchal, Uli Wortmann, Fabrice Popineau and Daniel Taupin.

Abstract

L^{A}TEX2HTML is a conversion tool that allows documents written in L^{A}TEX to become part of the World-Wide Web. In addition, it offers an easy migration path towards authoring complex hyper-media documents using familiar word-processing concepts, including the power of a L^{A}TEX-like macro language capable of producing correctly structured HTML tags.

L^{A}TEX2HTML replicates the basic structure of a L^{A}TEX document as a set of interconnected HTML files which can be explored using automatically generated navigation panels. The cross-references, citations, footnotes, the table-of-contents and the lists of figures and tables, are also translated into hypertext links. Formatting information which has equivalent “tags” in HTML (lists, quotes, paragraph-breaks, type-styles, etc.) is also converted appropriately. The remaining heavily formatted items such as mathematical equations, pictures etc. are converted to images which are placed automatically at the correct position in the final HTML document.

L^{A}TEX2HTML extends L^{A}TEX by supporting arbitrary hypertext links and symbolic cross-references between evolving remote documents. It also allows the specification of conditional text and the inclusion of raw HTML commands. These hyper-media extensions to L^{A}TEX are available as new commands and environments from within a L^{A}TEX document.

This document presents the main features of L^{A}TEX2HTML and describes how to obtain and install it, and how to use it effectively.

Contents

1 Overview 1 1.1 List of Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 Installation and Further Support 4 2.1 Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.2 Installation on Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.3 Getting L^{A}TEX2HTML . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.4 Installing L^{A}TEX2HTML . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.5 Getting Support and More Information . . . . . . . . . . . . . . . . . . . . . 11

3 Special Features 11 3.1 Variation with HTML Versions . . . . . . . . . . . . . . . . . . . . . . . . . . 12 3.2 Internationalisation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

3.2.1 Alternate Character Encodings . . . . . . . . . . . . . . . . . . . . . . 13 3.2.2 Multi-lingual documents, using Images . . . . . . . . . . . . . . . . . . 13

3.3 Mathematics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.4 Figures and Image Conversion . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

3.4.1 An Embedded Image Example . . . . . . . . . . . . . . . . . . . . . . 20 3.4.2 Image Sharing and Recycling . . . . . . . . . . . . . . . . . . . . . . . 21 3.4.3 Quality of Printed Images . . . . . . . . . . . . . . . . . . . . . . . . . 22

3.5 Figures, Tables and Arbitrary Images . . . . . . . . . . . . . . . . . . . . . . 22 3.6 Document Classes and Options . . . . . . . . . . . . . . . . . . . . . . . . . . 24 3.7 Packages and Style-Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

3.7.1 Fancy List-Markers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.7.2 Support for FoilTEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 3.7.3 Indicating Differences between Document Versions . . . . . . . . . . . 28

3.8 Indexing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 3.8.1 Integrated Glossary and Index . . . . . . . . . . . . . . . . . . . . . . 30

4 Hypertext Extensions to L^{A}TEX 32 4.1 Hyper-links in L^{A}TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 4.2 Including Arbitrary HTML Mark-up and Comments . . . . . . . . . . . . . . 37 4.3 Arbitrary Tags and Attributes . . . . . . . . . . . . . . . . . . . . . . . . . . 38 4.4 Conditional Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 4.5 Symbolic References shown as Hyperized Text . . . . . . . . . . . . . . . . . . 43 4.6 Hypertext Links in Bibliographic References (Citations) . . . . . . . . . . . . 45 4.7 Symbolic References between Living Documents . . . . . . . . . . . . . . . . . 46

4.7.1 Cross-Referencing Example . . . . . . . . . . . . . . . . . . . . . . . . 47 4.8 Miscellaneous commands for HTML effects . . . . . . . . . . . . . . . . . . . . . 48 4.9 Active Image Maps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.10 Document Segmentation^{1} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

4.10.1 A Segmentation Example . . . . . . . . . . . . . . . . . . . . . . . . . 53 5 User Manual 56

1This feature is supported only for users of L^{A}TEX 2ε.

5.1 Developing Documents using L^{A}TEX2HTML . . . . . . . . . . . . . . . . . . . . 56 5.2 Command-Line Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

5.2.1 Options controlling Titles, File-Names and Sectioning . . . . . . . . . 58 5.2.2 Options controlling Extensions and Special Features . . . . . . . . . . 60 5.2.3 Switches controlling Image Generation . . . . . . . . . . . . . . . . . . 63 5.2.4 Switches controlling Navigation Panels . . . . . . . . . . . . . . . . . . 65 5.2.5 Switches for Linking to other documents . . . . . . . . . . . . . . . . . 66 5.2.6 Switches for Help and Tracing . . . . . . . . . . . . . . . . . . . . . . 67 5.2.7 Other Configuration Variables, without switches . . . . . . . . . . . . 68

5.3 Extending the Translator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 5.3.1 Asking the Translator to Ignore Commands . . . . . . . . . . . . . . . 73 5.3.2 Asking the Translator to Pass Commands to L^{A}TEX . . . . . . . . . . . 74 5.3.3 Handling “order-sensitive” Commands . . . . . . . . . . . . . . . . . . 74

5.4 Customising the Navigation Panels . . . . . . . . . . . . . . . . . . . . . . . . 74

6 Known Problems 76 6.1 Troubleshooting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

7 Credits 84

List of Figures

1 Images of equation displays, at normal screen resolution . . . . . . . . . . . . 17 2 A sample figure showing part of a page generated by L^{A}TEX2HTML containing

a customised navigation panel (from the CSEP project). . . . . . . . . . . . . 21 3 Displayed math environments with extra-scale of 1.5 . . . . . . . . . . . . . . 22 4 Displayed math environments with extra-scale of 2.0 . . . . . . . . . . . . . . 23 5 An electronic form. In the online version the form would be active. . . . . . . 38 6 Example use of macros for raw HTML code. . . . . . . . . . . . . . . . . . . . . 39

List of Tables

1 Supported Font-encodings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2 Mathematics translation strategies, for HTML versions 3.0 and 3.2,

using <SUP> and <SUB> tags and <TABLE>s . . . . . . . . . . . . . . . . . . . 15 3 Mathematics translation strategies, for HTML version 2.0 . . . . . . . . . . . . 16 4 A sample table taken from [1] . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 5 Alternate view of the table from [1] . . . . . . . . . . . . . . . . . . . . . . . . 23 6 Supported L^{A}TEX2HTML packages and style-files. . . . . . . . . . . . . . . . . . 26

1 Overview

This manual describes the L^{A}TEX2HTML translator which is used to create Web pages from document source written for the L^{A}TEX typesetting system, or simply containing L^{A}TEX commands.

To use L^{A}TEX2HTML to translate a file <file>.tex containing L^{A}TEX commands, simply type:

     latex2html <file>.tex

This will create a new directory called <file> which will contain the generated HTML files, some log files and possibly some images.

Basically the translator reads the source document and creates a linked set of HTML pages, displaying the information it contains. The L^{A}TEX commands and environments that are found are interpreted either as “markup” instructions, or as macros expanding into more text or markup commands. Where such markup corresponds to the intended use for markup tags in the HTML language, a direct translation is made. If there is no natural way to present the information using simple text embellished with HTML markup tags, then an image is generated, using L^{A}TEX itself to interpret the portion of code.

Of course this is a drastically over-simplified description of what L^{A}TEX2HTML actually does. Many questions spring readily to mind. The answers to these and the options available to handle particular situations are discussed elsewhere in this manual.

• What does “natural way to present the information” really mean?

Text and paragraphing clearly should appear as such, whether printed or on-screen. Different font sizes and styles such as “bold-face” or “italic” are generally rendered accordingly. However, whereas L^{A}TEX has access to appropriate fonts for specialised purposes such as mathematical symbols, these cannot be guaranteed to be available with all Web-browsers. So for information requiring such things, L^{A}TEX2HTML will generally resort to making an image, using L^{A}TEX itself to typeset the material required for that image.

Section 1.1 contains a brief overview of how L^{A}TEX’s standard environments are handled within L^{A}TEX2HTML. It also mentions some of the extra features that are available. In general L^{A}TEX2HTML attempts to use textual constructions to represent the required information. Generation of an image is done only when there is no adequate textual construction with the required version of HTML, or when specifically requested to do so. Various extensions, to cope with the different HTML versions and extra features, are discussed in Section 3. That describes what to expect on the HTML pages, with little or no changes required to the L^{A}TEX source.

Just as L^{A}TEX has various packages which can be used to present specific types of information in appropriate ways, so is L^{A}TEX2HTML capable of handling the commands from many of these packages. See Table 6 for a listing of those packages which currently have special support.

• Some features of HTML have no direct counterpart in a L^{A}TEX typeset document. Can such features be used with L^{A}TEX2HTML?

Any effect currently available with any version of the HTML standard can be specified for a document processed by L^{A}TEX2HTML. New L^{A}TEX commands are defined in the html.sty package; the features that these commands allow are the subject of Section 4 in this manual. Some of the new commands provide improved strategies for effects already existing in L^{A}TEX; e.g. cross-references and citations. To use these effectively requires only small changes to the L^{A}TEX source.

Other commands define new environments which are completely ignored when processed by L^{A}TEX. Indeed the full scope of HTML 3.2 is available, using L^{A}TEX-like macros to help structure the source, reduce the tedium of repetitious use of tags, and ensure that all appropriate tags are correctly closed.

• What determines the amount of information that goes onto a single HTML page? How are different pages linked?

The HTML pages can contain whole chapters, sections, (sub)subsections or (sub)paragraphs. This is fully customisable using the command-line options discussed in detail in Section 5.2 of this manual.

• Does the original document have to be a valid L^{A}TEX document, typesetting without errors? If not, does it help if it is?

In fact any document can be fed to the L^{A}TEX2HTML processor, but it is designed specifically to recognise and sensibly translate the intentions expressed by L^{A}TEX markup commands. Although sensible results can be obtained even when the L^{A}TEX source is not valid, the most reliable translations are obtained when it is. Relevant issues are discussed in Section 5.1.

• When developing a document which contains special HTML features, is it best to regularly test it in L^{A}TEX or with L^{A}TEX2HTML?

The answer to such a question changes as the developer gains more experience with the available tools. Some aspects to be considered are discussed in Section 5.1 of this manual.

Information relevant to obtaining the latest version of L^{A}TEX2HTML, installation within the local environment, and where to look for help when things do not go as expected, can be found in Section 2.

What follows next is a brief summary of the features supported within L^{A}TEX2HTML.

1.1 List of Features

Following is a listing of the main features of the translator; more specific details on these is given elsewhere in this manual. The L^{A}TEX2HTML translator …

• breaks up a document into one or more components as specified by the user^{2};

• provides optional, customisable iconic navigation panels on every page which contain links to other parts of the document, or other documents;

• handles inlined equations ( ∑_{i=1}^{n} x_{i} = ∫_{0}^{1} f), handles equation alignment (A_{B_{C+D}}), right-justified numbered equations (see equation 1), tables (see Table 4), figures (see Figure 2), and any arbitrary environment. Either the complete environment or subparts thereof are passed to L^{A}TEX for conversion to images, which are then either included in the document or are made available through hypertext links.

2The user can specify the depth at which the document should not be broken up any further.

• figures or tables can be arbitrarily scaled and oriented, and shown either as inlined images or “thumbnail” sketches or their contents displayed within a table constructed using the <TABLE> tags of HTML 3.2.

• theorem-like environments are supported, along with automatic numbering and counter dependencies.

• can produce output suitable for browsers that support inlined images or characterbased browsers (as specified by the user). In particular the TEX or L^{A}TEX code for mathematical expressions and formulas will be displayed in character-based browsers, such as lynx.

• coloured text and/or background is fully supported, as is the ability to use an image to create a tiled backdrop.

• handles definitions of new commands, environments and counters even when these are defined in external files for input^{3};

• handles footnotes^{4}, tables of contents, lists of figures and tables, bibliographies and can generate an index. By including hyperlinks between index entries, simple navigation aids can be built into the index, for easy browsing.

• automatically translates cross-references and citations into hyper-links, and extends the L^{A}TEX cross-referencing mechanism to work not just within a document but between documents which may reside in remote locations;

• translates L^{A}TEX accent and special character commands (e.g. Â Ø ö £ © ¶) to the equivalent ISO–Latin–1 or Unicode character set, else an image can be created;

• recognises hypertext links (to multi-media resources or arbitrary Internet services such as sound, video, ftp, http, news) and links which invoke arbitrary program scripts—all expressed as L^{A}TEX commands;

• recognises conditional text which is intended only for the hypertext version, or only for the paper (.dvi) version;

• can include raw HTML in a L^{A}TEX document (e.g. in order to specify interactive forms);

• can deal sensibly with virtually all of the concepts and commands described in the L^{A}TEX blue book , where there is a meaningful interpretation appropriate to an HTML document. Also many other L^{A}TEX constructions are handled, including many described in the L^{A}TEX Companion[2] and L^{A}TEX Graphics Companion[3, XY-pic];

• can be configured to translate equations either as GIF images or as HTML 3.0 mark-up (as browsers become available which are suitable for the task), or by making images of subparts of equations, as required.

• links symbolic references across document segments which have been independently processed;

• will try to translate any document with embedded L^{A}TEX commands, irrespective of whether it is complete or syntactically legal.

3This allows the definition of HTML macros in L^{A}TEX !

4Like this!

2 Installation and Further Support

2.1 Requirements

The translator makes use of several utilities all of which are freely available on most platforms.

You can choose between two ways to do the installation of the required tools: either go the convenient way and install binary distributions (no compilation required, just install out of the box), or get and compile a source code distribution. You will stick to the latter in case you have a special kind of operating system or want to make customisations prior to compilation such as applying source level patches. Windows users will probably want to read the section about installation on Windows.

For the best use of L^{A}TEX2HTML you want to get the latest versions of all the utilities that it uses. (It will still work with earlier versions, but some special effects may not be possible. The specific requirements are discussed below.)

• Perl version 5.003, or later (check with perl -v);

• L^{A}TEX, meaning L^{A}TEX 2ε dated <1995/06/01>, or later;

• dvips or dvipsk, at version 5.58 or later;

• Ghostscript at version 4.02 or later;

• netpbm library of graphics utilities, version 1-mar-94

     (check with pnmcrop ‘ -version ’).

• pdftocairo, if you want to produce SVG images (available through the poppler-utils package)

More specific requirements for using L^{A}TEX2HTML depend on the kind of translation you would like to perform, as follows:

1. L^{A}TEX commands but without equations, figures, tables, etc.

• Perl Note: L^{A}TEX2HTML requires Perl 5 to operate.

• DBM or NDBM, the Unix DataBase Management system, or GDBM, the GNU database manager. Note: Some systems lack any DBM support. Perl 5 comes with its own database system SDBM, but it is sometimes not part of some Perl distributions. The installation script install-test will check that for you. If no database system is found, you will have to install Perl properly.

2. L^{A}TEX commands with equations, figures, tables, etc. As above plus …

• latex (version 2e recommended but 2.09 will work — with reduced ability to support styles and packages);

• dvips (version 5.516 or later) or dvipsk Version 5.62 or higher enhances the performance of image creation with a significant speed-up. See latex2html.config for this after you are done with the installation. Do not use the ’dvips -E’ feature unless you have 5.62, else you will get broken images.

• gs Ghostscript (version 4.03 or later); with the ppmraw device driver, or even better pnmraw. Upgrade to 5.10 or later if you want to go sure about seldom problems with 4.03 to avoid (yet unclarified).

• The netpbm library of graphics utilities; netpbm dated 1 March 1994 is required, else part of the image creation process will fail.

    Check with: pnmcrop ‘ -version ’.

Several of the filters in those libraries are used during the PostScript to GIF conversion.

• If you want PNG images, you need pnmtopng (current version is 2.31). It is not part of netpbm and requires libpng-0.89c.tar.gz and libz (1.0.4) (or later versions). pnmtopng supports transparency and interlace mode. Netscape Navigator as well as MS IE do support inlined PNG images.

3. Segmentation of large documents If you wish to use this feature, you will have to upgrade your L^{A}TEX to L^{A}TEX 2ε. Some other hyperlinking features also require L^{A}TEX 2ε.

4. Transparent inlined images If you dislike the white background color of the generated inlined images then you should get either the netpbm library (instead of the older pbmplus) or install the

giftrans filter by Andreas Ley ley@rz.uni-karlsruhe.de. L^{A}TEX2HTML now sup-

ports the shareware program giftool (by Home Pages, Inc., version 1.0), too. It can also create interlaced GIFs.

If Ghostscript or the netpbm library are not available, it is still possible to use the translator with the ‘ -no images ’ option.

If you intend to use any of the special features of the translator (see page 32) then you have to include the html.sty file in any L^{A}TEX documents that use them.

If only a character-based browser, such as lynx, is available, or if you want the generated documents to be more portable, then the translator can be used with the ‘ -ascii mode ’ option (see Section 5.2.3).

2.2 Installation on Windows

For Windows 95, 98, and NT you will either need a newer 2html Release 99.1 or higher, or if you like to try an older release get the Windows 97.1 port by Fabrice Popineau from ftp://ftp.ese-metz.fr/pub/TeX/win32. Meanwhile, all the operating system dependent issues are integrated into the main release, thanks to the cool work of Marek Rouchal.

To install the tools required to run the translator, perform the steps below. Thanks to Jens Berger (jberger@mail.zedat.fu-berlin.de) for providing this list!

• install WinZip;

• install TEX/L^{A}TEX 2ε and dvips;

• install Perl;

E.g. ActivePerl 509 or higher from http://www.activestate.com. Windows 95

users will also need DCOM, it is listed on that download page, too.

• install GhostScript;
E.g. Aladdin GhostScript 5.50.

   • install the NetPBM tools library from ftp://ftp.ese-metz.fr/pub/TeX/win32/.

• unpack L^{A}TEX2HTML, e.g. under C:\TEXMF\LATEX2HTML;

• check that the path to GSWIN32C.EXE is added to the PATH variable in your

     AUTOEXEC.BAT;

• with L^{A}TEX2HTML 99.1 or higher, edit l2hconf.pin, then run CONFIG.BAT to install the translator; with older releases, edit LATEX2HTML.CONFIG and run

          cd c:\texmf\latex2html
          perl install-test

from within a DOS box.

• you might want to write a small .BAT file in your L^{A}TEX2HTML directory:

          perl c:\texmf\latex2html\latex2html %1 %2 %3 >> l2h.log

%1 is the name of the .TEX file, %2 and %3 are optional arguments to the translator

     such as ‘ -"-split 3" ’.

Note that if you want more than two arguments you will need to supply more parameters to the .BAT file.

• run it with a test document test.tex:

           l2h test

Maybe you will need to run L^{A}TEX before this, too!

2.3 Getting L^{A}TEX2HTML

L^{A}TEX2HTML is available through the Fedora, Debian, and Ubuntu package managers for Linux, and the macports package manager for MacOS.

Releases of L^{A}TEX2HTML may also be obtained at https://www.github.com/

latex2html/latex2html/releases ,

Finally there is the L^{A}TEX2HTML developers’ source repository, at https://www.github.

com/latex2html/latex2html/ .

The files to be found here are the most up-to-date with current developments, but they cannot be guaranteed to be fully reliable. New features may be still under development and not yet sufficiently tested for release. A daily updated compressed archive of the develop-

ers’ work may be downloaded from https://www.github.com/latex2html/latex2html/
archive/master.zip .

Having obtained a compressed tar version, save it into a file latex2html-2019.tar.gz say, then extract its contents with

    % gzip -d latex2html-2019.tar.gz
    % tar xvf latex2html-2019.tar

You should then have the following:

• README file;

• Changes index with latest changes;

• (no longer supplied);

• latex2html Perl script;

• texexpand Perl script^{5};

• latex2html.config configuration file;

• install-test Perl script, for installation and testing;

• dot.latex2html-init sample initialisation file;

• texinputs/ subdirectory, containing various L^{A}TEX style-files;

• versions/ subdirectory, containing code for specific HTML versions;

• makemap Perl script;

• example/ subdirectory, containing the segmentation example, described in detail in Section 4.10;

• .dvipsrc file;

• pstoimg Perl script for image conversion (replaces pstogif);

• configure-pstoimg Perl script for installation;

• local.pm Perl input file;

• icons.gif/ subdirectory, containing icons in GIF format;

• icons.png/ subdirectory, containing icons in PNG format;

• makeseg Perl script and examples to handle segmented documents via a generated Makefile, see makeseg.tex;

• docs/foilhtml/ contains L^{A}TEX package and Perl implementation by Boris Veytsman, to support FoilTEX to HTML translation;

• IndicTeX-HTML/ package that contains Perl and L^{A}TEX code for translating IndicTEX documents (see README file);

• docs/ subdirectory, containing the files needed to create a version of this manual;

• styles/ subdirectory, containing Perl code for handling some style-files;

• tests/ contains some test documents for L^{A}TEX2HTML.

5Initially written by Robert S. Thau, completely rewritten by Marek Rouchal and Jens Lippmann.

2.4 Installing L^{A}TEX2HTML

To install L^{A}TEX2HTML you MUST do the following:

1. Specify where Perl is on your system.

In each of the files latex2html, texexpand, pstoimg, install-test and makemap,

modify the first line saying where Perl is on your system. Some system administrators do not allow Perl programs to run as shell scripts. This means that you may not be able to run any of the above programs. In this case change the first line in each of these programs from #!/usr/local/bin/perl to:

    # *-*-perl-*-*
        eval 'exec perl -S  $0 "$@"'
        if $running_under_some_shell;

2. Copy the files to the destination directory.

Copy the contents of the texinputs/ directory to a place where they will be found by L^{A}TEX, or set up your TEXINPUTS variable to point to that directory.

3. Run install-test .

This Perl script will make some changes in the latex2html file and then check whether the path-names to any external utilities required by latex2html are correct. It will not actually install the external utilities. install-test asks you whether to configure for GIF or PNG image generation. Finally it creates the file local.pm which houses pathnames for the external utilities determined earlier.

You might need to make install-test executable before using it.

Use chmod +x install-test to do this.
You may also need to make the files pstogif, texexpand, configure-pstoimg and

latex2html executable if install-test fails to do it for you.

4. If you like so, copy or move the latex2html executable script to some location outside the $LATEX2HTMLDIR directory.

5. You might want to edit latex2html.config to reflect your needs. Read the instructions about $ICONSERVER carefully to make sure your HTML documents will be displayed right via the Web server. While you’re at it you may want to change some of the default options in this file. If you do a system installation for many users, only cope with general aspects; let the

user override these with $HOME/.latex2html-init.

Note that you must run install-test now; formerly you could manage without. If you want to reconfigure L^{A}TEX2HTML for GIF/PNG image generation, or because some of the external tools changed the location, simply rerun configure-pstoimg.

This is usually enough for the main installation, but you may also want to do some of the following, to ensure that advanced features of L^{A}TEX2HTML work correctly on your system:

• To use the new L^{A}TEX commands which are defined in html.sty: Make sure that L^{A}TEX knows where the html.sty file is, either by putting it in the same place as the other style-files on your system, or by changing your TEXINPUTS shell environment variable, or by copying html.sty into the same directory as your L^{A}TEX source file.

The environment variable TEXINPUTS is not to be confused with the L^{A}TEX2HTML installation variable $TEXINPUTS described next.

• There is an installation variable in latex2html.config called $TEXINPUTS, which tells L^{A}TEX2HTML where to look for L^{A}TEX style-files to process. It can also affect the input-path of L^{A}TEX when called by L^{A}TEX2HTML, unless the command latex is really a script which overwrites the $TEXINPUTS variable prior to calling the real latex. This variable is overridden by the environment variable of the same name if it is set.

• The installation variable $PK GENERATION specifies which fonts are used in the generation of mathematical equations. A value of “0” causes the same fonts to be used as those for the default printer. Because they were designed for a printer of much greater resolution than the screen, equations will generally appear to be of a lower quality than is otherwise possible. To cause L^{A}TEX2HTML to dynamically generate fonts that are designed specifically for the screen, you should specify a value of “1” for this variable. If you do, then check to see whether your version of dvips supports the command-line option ‘ -mode ’ . If it does, then also set the installation variable $DVIPS MODE to a low resolution entry from modes.mf, such as toshiba.

It may also be necessary to edit the MakeTeXPK script, to recognise this mode at the appropriate resolution.

If you have PostScript fonts available for use with L^{A}TEX and dvips then you can probably ignore the above complications and simply set $PK GENERATION to “0” and $DVIPS MODE to "" (the empty string). You must also make sure that gs has the locations of the fonts recorded in its gs fonts.ps file. This should already be the case where GS-Preview is installed as the viewer for .dvi-files, using the PostScript fonts.

If dvips does not support the ‘ -mode ’ switch, then leave $DVIPS MODE undefined, and verify that the .dvipsrc file points to the correct screen device and its resolution.

• The installation variable $AUTO PREFIX allows the filename-prefix to be automatically set to the base filename-prefix of the document being translated. This can be especially useful for multiple-segment documents.

• The makemap script also has a configuration variable $SERVER, which must be set to either CERN or NCSA, depending on the type of Web-server you are using.

• To set up different initialization files:

For a “per user” initialization file, copy the file dot.latex2html-init in the home directory of any user that wants it, modify it according to her preferences and re-

name it as .latex2html-init. At runtime, both the latex2html.config file and

$HOME/.latex2html-init file will be loaded, but the latter will take precedence.

You can also set up a “per directory” initialization file by copying a version of .latex2html-init in each directory you would like it to be effective. An initialization file /X/Y/Z/.latex2html-init will take precedence over all other initialization files if /X/Y/Z is the “current directory” when L^{A}TEX2HTML is invoked.

Warning: This initialization file is incompatible with any version of L^{A}TEX2HTML prior to v96.1 . Users must either update this file in their home directory, or delete it altogether.

• To make your own local copies of the L^{A}TEX2HTML icons: Please copy the icons/ subdirectory to a place under your WWW tree where they can be served by your server. Then modify the value of the $ICONSERVER variable in latex2html.config accordingly. Alternatively, a local copy of the icons can be included within the subdirectory containing your completed HTML documents. This is most easily done using the ‘ -local icons ’ command-line switch, or by setting $LOCAL ICONS to “1” in latex2html.config or within an initialization file, as described above.

• To make your own local copy of the L^{A}TEX2HTML documentation: This will also be a good test of your installation. Firstly, to obtain the .dvi version for printing, from within the docs/ directory it is sufficient to type:

 make manual.dvi

This initiates the following sequence of commands:

    latex manual.tex
    makeindex -s l2hidx.ist manual.idx
    makeindex -s l2hglo.ist -o manual.gls manual.glo
    latex manual.tex
    latex manual.tex

...in which the two configuration files l2hidx.ist and l2hglo.ist for the makeindex program, are used to create the index and glossary respectively. The 2nd run of latex is needed to assimilate references, etc. and include the index and glossary. (In case makeindex is not available, a copy of its outputs manual.ind and manual.gls are included in the docs/ subdirectory, along with manual.aux .) The 3rd run of latex is needed to adjust page-numbering for the Index and Glossary within the Table-of-Contents.

Next, the HTML version is obtained by typing:

make manual.html

This initiates a series of calls to L^{A}TEX2HTML on the separate segments of the manual; the full manual is thus created as a “segmented document” (see Section 4.10). The whole process may take quite some time, as each segment needs to be processed at least twice, to collect the cross-references from other segments.

The files necessary for correct typesetting of the manual to be found within the docs/ subdirectory. They are as follows:

  – style-files: l2hman.sty, html.sty, htmllist.sty, justify.sty,
    changebar.sty and url.sty
  – inputs: credits.tex, features.tex, hypextra.tex,
    licence.tex, manhtml.tex, manual.tex, overview.tex,
    problems.tex, support.tex and userman.tex

– sub-directory: psfiles/ containing PostScript graphics used in the printed version of this manual

– images of small curved arrows: up.gif, dn.gif

– filename data: l2hfiles.dat

  – auxiliaries: manual.aux, manual.ind, manual.gls

The last three can be derived from the others, but are included for convenience.

• To get a printed version of the ‘Changes’ section:

Due to the burgeoning size of the Changes file with successive revisions of L^{A}TEX2HTML, the ‘Changes’ section is no longer supported for the manual. Please refer to text file Changes instead which is part of the distribution.

• To join the community of L^{A}TEX2HTML users:

More information on a mailing list, discussion archives, bug reporting forms and more

     is available at https://www.latex2html.org/

• L^{A}TEX2HTML is actively supported by the international TEX Users Group (http:// www.tug.org). All users are encouraged to join http://www.tug.org, to keep up-to-date with the latest development in TEX, L^{A}TEX, L^{A}TEX2HTML and related programs. Consult the TUG Web pages at http://www.tug.org/.

2.5 Getting Support and More Information

A L^{A}TEX2HTML mailing list is managed by the international TEX User Group (http://www.

tug.org).

This mailing used was originally established at the Argonne National Labs, where it was based for several years. (Thanks to Ian Foster and Bob Olson, and others.) Since February 1999, it has been run by http://www.tug.org, thanks to Art Ogawa and Ross Moore.

To join send a message to: latex2html-request@tug.org

with the contents: subscribe . To be removed from the list send a message to: latex2html-request@tug.org with the contents: unsubscribe .

The mailing list also has a searchable online archive. It is recommended to start with this, to become familiar with the topics actually discussed, and to search for articles related to your own interests.

Enjoy!

3 Special Features

This section describes major features available for processing documents using L^{A}TEX2HTML. Firstly the means whereby L^{A}TEX2HTML can be configured to produce output for the different versions of HTML is discussed in Section 3.1. Following this is a description, in Section 3.2, of how to use languages other than English. The options available with the creation and reuse of images, are presented in Section 3.4, for those situations where a textual representation is inadequate or undesirable.

There are several strategies available for the presentation of mathematics according to the desired version of HTML. These are discussed in some detail, in Section 3.3. Environments such as figure, table, tabular and minipage have special features which are discussed in Section 3.5. Other supported packages are listed in Table 6.

3.1 Variation with HTML Versions

The Hypertext Mark-up Language (HTML) is an evolving standard, with different versions supporting different features. In order to make your documents viewable by the widest possible audience, you should use the most advanced HTML version with widely-accepted usage.

Sometimes it is known that the audience, for which a specific document is intended, has limited browser capabilities. Or perhaps special extended capabilities are known to be available. The L^{A}TEX2HTML translation may be customised to suit the available functionality.

Other HTML versions and extensions supported by L^{A}TEX2HTML are described below. See the description of the ‘ -html version ’ command-line option switch, on page 60.

Version 2.0 This provides only the functionality of the HTML 2.0 standard. There is little provision for aligning headings, paragraphs or images nor for super/subscripts to be generated. Images are created for tables and other environments that use <TABLE> tags with HTML 3.2; e.g. eqnarray and equation with equation numbering.

i18n (internationalised fonts) This extension (formerly known as HTML version 2.1) provides extensions for internationalisation. Most importantly, the default character set is no longer ISO–8859–1 but ISO–10646 (Unicode). This is a 16-bit character set and can thus display a much larger set of characters. There are also provisions for bidirectional languages (e.g. in Arabic the text is written from right to left, but numerals from left to right), and provisions in HTML to determine the character set and the language used.

Not all of the symbols are available in TEX, L^{A}TEX2HTML, or any browser yet available. However the ‘i18n’ extension to L^{A}TEX2HTML is in preparation for when such browsers do become available, and such characters will be required in Webaccessible documents.

math (HTML3 model) This extension (formerly referred to as HTML version 3.1) adds support for the HTML-Math model, originally part of the proposed HTML 3.0 standard, see above. The only available browser which can display this mark-up is Arena. Originally developed by the World Wide Web Consortium as a test-bed browser, it is no longer supported by them.

There has been a recent proposal for a Mathematical Markup Language (MathML) from the W3C Math Working Group. This would suggest that the HTML-Math model is unlikely ever to be adopted; better things being expected in the near future using MathML.

See also Section 3.3 for a discussion the the mechanisms available with L^{A}TEX2HTML for handling mathematical equations and expressions.

3.2 Internationalisation

A special variable $LANGUAGE TITLES in the initialisation or configuration files determines the language in which some section titles will appear. For example setting it to

 $LANGUAGE_TITLES = 'french';

will cause L^{A}TEX2HTML to produce “Table des matières” instead of “Table of Contents”. Furthermore, the value of the \today command is presented in a format customary in that language.

Languages currently supported are finnish, french, english, german and spanish. It is trivial to add support for another language by creating a file in the styles/ subdirectory, or by adding to the file latex2html.config. As a guide, here is the entry for French titles:

    sub french_titles {
        $toc_title = "Table des mati\\`eres";
        $lof_title = "Liste des figures";
        $lot_title = "Liste des tableaux";
        $idx_title = "Index";
        $ref_title = "R\\'ef\\'erences";
        $bib_title = "R\\'ef\\'erences";
        $abs_title = "R\\'esum\\'e";
        $app_title = "Annexe";
        $pre_title = "Pr\\'eface";
        $fig_name = "Figure";
        $tab_name = "Tableau";
        $part_name = "Partie";
        $prf_name = "Preuve";
        $child_name = "Sous-sections";
        $info_title = "\\`Apropos de ce document...";
        @Month = ('', 'janvier', "f\\'evrier", 'mars', 'avril', 'mai',
                  'juin', 'juillet', "ao\\^ut", 'septembre', 'octobre',
                  'novembre', "d\\'ecembre");
        $GENERIC_WORDS = "a|au|aux|mais|ou|et|donc|or|ni|car|l|la|le|les"
            . "|c|ce|ces|un|une|d|de|du|des";
    }

Notice how the backslash needs to be doubled, when a macro is needed (for accented characters, say). Also, the $GENERIC WORDS are a list of short words to be excluded when filenames are specially requested to be created from section-headings. In order to provide full support for another language you may also replace the navigation buttons which come with L^{A}TEX2HTML (by default in English) with your own. As long as the new buttons have the same file-names as the old ones, there should not be a problem.

3.2.1 Alternate Character Encodings

By default, L^{A}TEX2HTML assumes that input files are Unicode encoded with UTF8, and produces Unicode UTF8 output.

L^{A}TEX2HTML can handle input files in other encodings, indicated by including the inputenc package in the source:

     \usepackage[latin5]{inputenc}

In this case, L^{A}TEX2HTML will produce output in the same encoding, and will indicate the encoding in the HTML headers. The input encodings that are recognised are listed in Table 1.

3.2.2 Multi-lingual documents, using Images

Some multi-lingual documents can be constructed, when all the languages can be presented using characters from a single font-encoding, as discussed in the Section 3.2.1.

extension notes  encoding
unicode  (default) ISO–10646 (Unicode)
latin1  ISO–8859–1 (ISO-Latin-1)
latin2  ISO–8859–2 (ISO-Latin-2)
latin3  ISO–8859–3 (ISO-Latin-3)
latin4  ISO–8859–4 (ISO-Latin-4)
latin5  ISO–8859–9 (ISO-Latin-5)
latin6  ISO–8859–10 (ISO-Latin-6)
koi8-r  RFC 1489 (Russian)

Table 1: Supported Font-encodings

Another way to present multiple languages within a Web document is to create images of individual letters, words, sentences, paragraphs or even larger portions of text, which cannot be displayed within the chosen font-encoding. This is a technique that is used with IndicTEX/HTML, for presenting traditional Indic language scripts within Web pages. For these the L^{A}TEX source that is to be presented as an image needs special treatment using a “pre-processor”. For the special styles defined in IndicTEX/HTML, running the preprocessor is fully automated, so that it becomes just another step within the entire image-generation process.

The technique of using images, can be used with any font whose glyphs can be typeset using TEX or L^{A}TEX. Using TEX’s \font command, a macro is defined to declare the special font required; e.g. for Cyrillic characters, using the Univ. of Washington font:

     \font\wncyr = wncyr10

Now use this font switch immediately surrounded by braces:

     published by {\wncyr Rus\-ski\char26\ \char23zyk}.

to get:

published by ������� ����.

3.3 Mathematics

There are various different ways in which L^{A}TEX2HTML can handle mathematical expressions and formulas:

• give a textual representation (“simple” math);

• make an image of the complete formula or expression;

• combination of textual representation and images of sub-expressions;

• SGML-like representation built using abstract “entities”; e.g. for the HTML-Math model, or for MathML.

Which is the most appropriate normally depends on the context, or importance of the mathematics within the entire document. What L^{A}TEX2HTML will produce depends upon

1. the version of HTML requested; 2. whether or not the special ‘math’ extension has been loaded;

3. whether the ‘ -no math ’ command-line option has been specified, or (equivalently) the $NO SIMPLE MATH variable has been set in an initialisation file.

The strategies used to translate math expressions are summarised in Table 2 for HTML 3.0+ and Table 3 for HTML 2.0.

‘math’  switch  strategy adopted
not loaded  —  textual representation where possible,

else image of whole expressions

not loaded ‘ -no math ’ always generates an image of the whole

expression/environment

loaded  —  uses entities and <MATH> tags; e.g. for

                                 HTML-Math (or MathML in future)

loaded ‘ -no math ’ textual representation where possible, with images of sub-expressions

Table 2: Mathematics translation strategies, for HTML versions 3.0 and 3.2,

using <SUP> and <SUB> tags and <TABLE>s

Using the ‘ -no math ’ switch is best for having a consistent style used for all mathematical expressions, whether inline or in displays. The images are of especially good quality when “anti-aliasing” is being used (see page 65), provided the browser is set to have a light background colour. (When set against a gray or dark background, these images can become rather faint and hard to read.)

The final strategy in Table 2, using ‘ -no math ’ is the preferred method for good quality mathematics with HTML version 3.2 . It combines the browser’s built-in fonts with the best quality images, when needed. To obtain it use the command-line option switches:

      -no math -html version 3.2,math

This is what was used when creating the HTML version of this manual. Examples below show how to generate an image of a whole environment, even with these options in force.

Since the HTML 2.0 standard does not include superscripts and subscripts, via the <SUP> and <SUB> tags, the options are more limited. In this case creating images of sub-expressions is not so attractive, since virtually the whole expression would consist of images in all but the simplest of cases.

Here are some examples of mathematical expressions and environments processed by L^{A}TEX2HTML using different strategies. They are automatically numbered …

Φ_{l+1,m,n} = (Φ + h\frac{∂Φ}{∂x} + \frac{1}{2}h^{2} \frac{∂^{2}Φ}{∂x^{2}} + \frac{1}{6}h^{3} \frac{∂^{3}Φ}{∂x^{3}} + …)_{l,m,n}  (1)

… with some gratuitously ácçënted text in-between …

\frac{Φ_{l+1,m,n} − 2Φ_{l,m,n} + Φ_{l−1,m,n}}{h^{2}} + \frac{Φ_{l,m+1,n} − 2Φ_{l,m,n} + Φ_{l,m−1,n}}{h^{2}}

+\frac{Φ_{l,m,n+1} − 2Φ_{l,m,n} + Φ_{l,m,n−1}}{h^{2}} = −I_{l,m,n}(v) .  (2)

‘math’  switch  strategy adopted
not loaded  —  textual representation where possible,

else image of whole expressions

not loaded ‘ -no math ’ always generates an image of

the whole expression or environment

loaded — entities and <MATH> tags for HTML-Math loaded ‘ -no math ’ always generates an image of the whole

expression or environment

Table 3: Mathematics translation strategies, for HTML version 2.0

The latter example uses an eqnarray environment and the \nonumber command to suppress the equation number on the upper line.

In the on-screen version of these equations simple alphabetic characters that are not part of fractions appear in the (italiced) text-font selected using the browser’s controls. This may appear slightly different from the same symbol being used within a fraction, or other mathematical construction requiring an image to be generated. This is most apparent with the letter ‘h’ in the first equation and the subscripts at the end of the second equation.

By inserting an \htmlimage{} command into a math, equation or displaymath environment, a single image will be created for the whole environment. For an eqnarray environment, this will lead to having a single separate image for each of the aligned portions. The argument to \htmlimage need not be empty, but may contain information which is used to affect characteristics of the resulting image. An example of how this is used is given below, and a fuller discussion of the allowable options is given in Section 3.4.

Scale-factors for Mathematics. When an image is to be made of a mathematical formula or expression, it is generally made at a larger size than is normally required on a printed page. This is to compensate for the reduced resolution of a computer screen compared with laser-print. The amount of this scaling is given by the value of a configuration variable $MATH SCALE FACTOR, by default set to 1 in latex2html.config. A further variable $DISP SCALE FACTOR is used with ‘displayed math’ equations and formulas. This value multiplies the $MATH SCALE FACTOR to give the actual scaling to be used. The main purpose of this extra scaling is to allow some clarity in super/subscripts etc.

Anti-aliased Images. Figure 1 shows the same equations as previously, this time as images of the complete contents of the equation environment, and complete aligned parts of rows in an eqnarray. These are images, as they would appear if the HTML page were to be printed from the browser. A scaling of 60% has been applied to counteract the combined effects of the $MATH SCALE FACTOR of 1.4 and $DISP SCALE FACTOR of 1.2, used for the HTML version of this manual. For a comparison, the second group of images use antialiasing effects, whereas the first image does not; a 600 dpi printing is probably necessary to appreciate the difference in quality. Compare these images with those in Section 3.4.3. Note: To generate anti-aliased images using Ghostscript requires version 4.03 or later. These images of the whole environment were created using the \htmlimage command, to suppress the extended parsing that usually occurs when the ‘math’ extension is loaded; viz.

    \begin{equation}
    \htmlimage{no_antialias}

    \Phi_{l+1,m,n} = \Bigl(\Phi+h\frac{\partial\Phi}{\partial x} +
    ...
    \end{equation}
    %
    \begin{eqnarray}
    \htmlimage{}
    \frac{\Phi_{l+1,m,n}-2\Phi_{l,m,n}+\Phi_{l-1,m,n}}{h^{2}} +
    ...
    \end{eqnarray}

Further aspects of the options available when generating images are discussed in the next section, in particular with regard to the quality of printed images.

The \mbox command. Another way to force an image to be created of a mathematical expression, when global settings are not such as to do this anyway, is via the \mbox command having math delimiters within its argument.

Normally \mbox is used to set a piece of ordinary text within a mathematics environment. It is not usual to have math delimiters $...$ or \(...\) within the argument of an \mbox. Whereas earlier versions of L^{A}TEX2HTML simply ignored the \mbox command (treating its argument as normal text), the presence of such delimiters now results in an image being generated of the entire contents of the \mbox. It is not necessary for there to be any actual mathematics inside the \mbox’s contents; e.g. \mbox{...some text...${}$} will cause an image to be created of the given text.

The \parbox command. The \parbox[<align>]{<width>}{<text>} command also

generates an image of its contents, except when used within a tabular environment, or other similar table-making environment. Here the important aspect is the width specified for the given piece of text, and any special line-breaks or alignments that this may imply. Hence to get the best effect, L^{A}TEX is used to typeset the complete \parbox, with its specified width, alignment and contents, resulting in an image.

The heqn package. If you need HTML 2.0 compatible Web pages, and have a document with a great many displayed equations, then you might try using the heqn package. Inclusion of the heqn.sty file has absolutely no effect on the printed version of the article, but it does change the way in which L^{A}TEX2HTML translates displayed equations and equation arrays. It

(3)

(4)

Figure 1: Images of equation displays, at normal screen resolution causes the equation numbers of the equation environment to be moved outside of the images themselves, so that they become order-independent and hence recyclable. Images that result from the eqnarray environment are also recyclable, so long as their equation numbers remain unchanged from the previous run.

The \nonumber command is recognised in each line of the equation array, to suppress the equation number. A side-effect of this approach is that equation numbers will appear on the left side of the page. The heqn package requires the html package. Using HTML Version 3.2 the heqn package is quite redundant, since equation numbers are placed in a separate <TABLE> cell to the mathematical expressions themselves. It is not required and should not be requested, since this will override some of the improved functionality already available.

3.4 Figures and Image Conversion

L^{A}TEX2HTML converts equations, special accents, external PostScript files, and L^{A}TEX environments it cannot directly translate into inlined images. This section describes how it is possible to control the final appearance of such images. For purposes of discussion …

• “small images” refers to inline math expressions, special accents and any other L^{A}TEX command which causes an image to be generated; while …

• “figures” applies to image-generating L^{A}TEX environments (e.g. makeimage, figure, table (with HTML 2.0), and displayed math environments when required to generate images, etc.).

These parameters apply only to bitmapped image types, and have no effect with the default SVG image type. The size of all “small images” depends on a configuration variable $MATH SCALE FACTOR which specifies how much to enlarge or reduce them in relation to their original size in the PostScript version of the document. For example a scale-factor of 0.5 will make all images half as big, while a scale-factor of 2 will make them twice as big. Larger scale-factors result in longer processing times and larger intermediate image files. A scale-factor will only be effective if it is greater than 0. The configuration variable $FIGURE SCALE FACTOR performs a similar function for “figures”. Both of these variables are initially set to have value 1.

A further variable $DISP SCALE FACTOR is used with ‘displayed math’ equations and formulas; this value multiplies the $MATH SCALE FACTOR to give the actual scaling used. Values greater than 1 can be used to counteract readability problems with bitmapped images. Accordingly this manual actually uses values of 1.4 and 1.2 respectively, for $MATH SCALE FACTOR and $DISP SCALE FACTOR. These go well with the browser’s text-font set at 14 pt. The next larger size of 17 pt is then used for the <LARGE> tags in displayed equations.

A further variable $EXTRA IMAGE SCALE allows images to be created at a larger size than intended for display. The browser itself scales them down to the intended size, but has the extra information available for a better quality print. This feature is also available with single images. It is discussed, with examples, in Section 3.4.3.

\htmlimage{<options>} For finer control, several parameters affecting the conversion of a single image can be controlled with the command \htmlimage, which is defined in html.sty. With version v97.1 use of this command has been extended to allow it to control whether an image is generated or not for some environments, as well as specifying effects to be used when creating this image.

If an \htmlimage command appears within any environment for which creating an image is a possible strategy (though not usual, due to loading of extensions, say), then an image will indeed be created. Any effects requested in the <options> argument will be used. Having empty <options> still causes the image to be generated.

This ability has been used within this manual, for example with the mathematics images in Figure 1.

The <options> argument is a string separated by commas. Allowable options are:

• scale=<scale-factor>

allows control over the size of the final image.

• external

will cause the image not to be inlined; instead it will be accessible via a hyperlink.

• thumbnail=<scale-factor>

will cause a small inlined image to be placed in the caption. The size of the thumbnail depends on the <scale-factor>, as a factor of the ‘natural size’ of the image, ignoring

any $FIGURE SCALE FACTOR or $MATH SCALE FACTOR, etc. which may be applicable

to the full-sized version of the image. Use of the ‘thumbnail=’ option implies the ‘external’ option.

• map=<server-side image-map URL>

specifies that the image is to be made into an active image-map. (See Section 4.9 for more information.)

• usemap=<client-side image-map URL> same as previous item, but with the imagemap processed by the client. (See Section 4.9 for more information.)

• flip=<flip option>

specifies a change of orientation of the electronic image relative to the printed version. The <flip option> is any single command recognised by the pnmflip graphics utility. The most useful of these include:

– ‘rotate90’ or ‘r90’ This will rotate the image clockwise by 90^{∘}.

– ‘rotate270’ or ‘r270’ This will rotate the image counterclockwise by 90^{∘}.

– ‘leftright’ This will flip the image around a vertical axis of rotation.

– ‘topbottom’ This will flip the image around a horizontal axis of rotation.

• align=<alignment>

specifies how the figure will be aligned. The choices are: ‘top’, ‘bottom’, ‘middle’,

‘left’, ‘right’ and ‘center’.

The ‘middle’ option specifies that the image is to be left-justified in the line, but centered vertically. The ‘center’ option specifies that it should also be centered horizontally. This option is valid only if the HTML version is 3.0 or higher. The default alignment is ‘bottom’.

• transparent, no transparent or notransparent

specify that a transparent background should (not) be used with this image, regardless of the normal behaviour for similar images.

• antialias, no antialias or noantialias

specify that anti-aliasing should (not) be used with this image, regardless of the normal behaviour for similar images.

• extrascale=<scale-factor>

is used mainly used with a <scale-factor> of 1.5 or 2, when it is important to get printed versions of the completed HTML pages. The image is created scaled by the amount specified, but it is embedded in the HTML page with attributes to the <IMG> of HEIGHT=... and WIDTH=..., indicating the unscaled size. A browser is supposed to display the image at the requested size by scaling the actual image to fit, effectively imposing its own anti-aliasing. Some examples of this effect are show later, in Section 3.4.3. This effect can be applied to all images in a document by setting the $EXTRA IMAGE SCALE variable. However it may be desirable to also turn off “anti-aliasing”, as these effects serve similar purposes but need not work well together. Furthermore different browsers may give results of different quality. It may be necessary to experiment a little, in order to find the combination that works best at your site.

• height=<dimen> and width=<dimen>

are used to specify exactly the size to be occupied by the image on the HTML page. The value(s) given this way overrides the natural size of the image and forces the browser to shrink or stretch the image to fit the specified size. The <dimen> can be given as either (i) a number (of points); or (ii) with any of the units of cm, mm, in, pt; or (iii) fraction of \hsize or \textwidth, to become a percentage of the browser window’s width, or of \vsize or \textheight for a percentage height.

Note: images whose sizes are modified in this way may not be acceptable for imagerecycling, (see page 3.4.2). Instead they may need to be generated afresh on each run of L^{A}TEX2HTML through the same source document.

In order to be effective the \htmlimage command and its options must be placed inside the environment on which it will operate. Environments for alignment and changing the font size do not generate images of their contents. Any \htmlimage command may affect the surrounding environment instead; e.g. within a table or figure environment, but does not apply to a minipage.

When the \htmlimage command occurs in an inappropriate place, the following message is printed among the warnings at the end of processing. The actual command is shown, with its argument; also the environment name and identifying number, if there is one.

    The command "\htmlimage" is only effective inside an environment
    which may generate an image (e.g. "{figure}", "{equation}")
     center92: \htmlimage{ ... }

3.4.1 An Embedded Image Example

The effect of the L^{A}TEX commands below can be seen in the thumbnail sketch of Figure 2. A 5 pt border has also been added around the thumbnail, using \htmlborder command; this gives a pseudo-3D effect in some browsers.

\begin{figure}
   \htmlimage{thumbnail=0.5}
   \htmlborder{5}
   \centering \includegraphics[width=5in]{psfiles/figure}

        \latex{\addtocounter{footnote}{-1}}
        \caption{A sample figure showing part of a page generated by
           \latextohtml{} containing a customised navigation panel
           (from the
            CSEP project).}\label{fig:example}
    \end{figure}

Figure 2: A sample figure showing part of a page generated by L^{A}TEX2HTML containing a customised navigation panel (from the CSEP project).

The \htmlimage command is also often useful to cancel-out the effect of the configuration variable $FIGURE SCALE FACTOR. For example to avoid resizing a color screen snap despite

the value of $FIGURE SCALE FACTOR it is possible to use \htmlimage{scale=0} .

3.4.2 Image Sharing and Recycling

It is not hard too see how reasonably sized papers, especially scientific articles, can require the use of many hundreds of external images. For this reason, image sharing and recycling is of critical importance. In this context, “sharing” refers to the use of one image in more than one place in an article. “Recycling” refers to the use of an image left over from a previous run of L^{A}TEX2HTML. Without this ability, every instance of an image would have to be regenerated each time even the slightest change were made to the document.

All types of images can be shared. These include “small images” and figures with or without thumbnails and image-maps. Furthermore, most images can also be reused. The only exception are those which are order-sensitive, meaning that their content depends upon their location. Examples of order-sensitive images are equation and eqnarray environments, when ‘ -html version 2.0 ’ has been specified; this is because their figure numbers are part of the image.

Figures and tables with captions, on the other hand, are order-insensitive because the figure numbers are not part of the image itself.Similarly when HTML 3.2 code is being produced, equation numbers are no longer part of the image. Instead they are placed in a separate cell of a <TABLE>. So most images of mathematical formulas can be reused also.

3.4.3 Quality of Printed Images

(5)

(6)

Figure 3: Displayed math environments with extra-scale of 1.5

Since it is often desirable to get a good quality print on paper directly from the browser, Figure 3 shows the same equations as on page 17. This time the ‘extrascale=1.5’ option has been used. This value of 1.5 means that more than twice the number of pixels are available, for a cost of approximately 1.7 times the disk-space^{5}. On-screen these images appear slightly blurred or indistinct. However there can be marked improvement in the print quality, when printed from some browsers; others may show no improvement at all. The “anti-aliasing” helps on-screen. In the printed version jagged edges are indeed softened, but leave an overall fuzziness.

Figure 4 shows the same equations yet again; this time with ‘extrascale=2.0’. Now there are 4 times the pixels at a cost of roughly 2.45 times the disk space. Compared with the previous images (having 1.5 times extra-scaling), there is little difference in the on-screen images. Printing at 300 dpi shows only a marginal improvement; but at 600 dpi the results are most satisfying, especially when scaled to be comparable with normal 10 pt type, as here.

3.5 Figures, Tables and Arbitrary Images

This section is to explain how the translator handles figures, tables and other environments. Compare the paper with the online version.

When the common version of HTML was only 2.0, then almost all complicated environments were represented using images. However with HTML 3.2, there is scope for sensible layout of tables, and proper facilities for associating a caption with a figure or table. To take advantage of this, the figure environment now has its contents placed within <TABLE> tags; any caption is placed as its <CAPTION>.

For consistency with former practice, the contents of the figure environment are usually represented by generating an image. This is frequently exactly what is required; but not always. On page 43 it is described how to use the makeimage environment, defined in the

5This figure varies with the graphics format used, and the complexity of the actual image.

(8)

Figure 4: Displayed math environments with extra-scale of 2.0

html.sty package, to determine just which parts (if any) of a figure environment’s contents should be made into images, the remainder being treated as ordinary text, etc.

table and tabular environments. Similarly the makeimage environment can be used within a table, though usually this is used with a tabular or other table-making environment, such as tabbing or longtable or supertabular. Here is a simple example, from the L^{A}TEX ‘blue book’ .

gnats  gram $13.65
each  .01
gnu  stuffed 92.50
emur  33.33
armadillo frozen  8.99

Table 4: A sample table taken from [1]

Table 5 is a screen-shot of how the resulting table appears on-screen, using a typical browser supporting HTML 3.2. Here it is scaled down by 70% to compensate for the 14 pt fonts being used when the screen-shot was taken.

Table 5: Alternate view of the table from [1] minipage environments. The special feature of minipage environments is in the way \footnote and \footnotemark commands are handled. These are numbered separately from the rest of the footnotes throughout the document, and the notes themselves are collected together to be displayed at the end of the minipage’s contents.

Variable Meaning
none  none
Jacobi  m-step Jacobi iteration^{a}
SSOR  m-step SSOR iteration^{1}
IC  Incomplete Cholesky factorization^{b}
ILU  Incomplete LU factorization^{2}

aone footnote banother footnote

The code used for this example was as follows^{6}

    \begin{minipage}{.9\textwidth}
    \renewcommand{\thempfootnote}{\alph{mpfootnote}}
    \begin{tabular}{|l|l|} \hline
    \textbf{Variable} & \textbf{Meaning} \\ \hline
    none      & none                   \\
    Jacobi    & $m$-step Jacobi iteration\footnote[1]{one footnote} \\
    SSOR      & $m$-step SSOR iteration\footnotemark[1] \\
    IC        & Incomplete Cholesky factorization\footnote[2]{another footnote} \\
    ILU       & Incomplete LU factorization\footnotemark[2] \\ \hline
    \end{tabular}
    \end{minipage}

Warning: With some figures, especially when containing graphics imported using \includegraphics or other special macros, the background color may come out as a shade of grey, rather than white or transparent. This is due to a setting designed to enhance anti-aliasing of text within images; e.g. for mathematics. To alleviate this possible problem, the ‘ -white ’ command-line option can be used, to ensure a white background for images of figure environments. Alternatively, set the $WHITE BACKGROUND variable (see section 5.2.3).

3.6 Document Classes and Options

In general the standard L^{A}TEX document-classes: article, report, book, letter, slides are translated by L^{A}TEX2HTML in the same way. Currently the only real difference is with the display of section-numbering, when the ‘ -show section numbers ’ switch is used, and when numbering of theorem-like environments is linked to section-numbering.

These differences are achieved using a mechanism that automatically loads a file:

article.perl, report.perl, book.perl, letter.perl, slides.perl according to the re-

quested document-class. These files contain Perl code and are located in the styles/ directory. If a file of the same name exists in the working directory, this will be loaded instead.

Typically such files <class>.perl contain code to define subroutines or sets values for variables that will affect how certain translations are performed. There can be code that is executed only when specific class-options are specified along with the chosen document-class. For example, the foils.perl implementation of FoilTEX’s foils class defines code create a new sub-section for each ‘foil’. It also has code which allows L^{A}TEX2HTML to ignore those of FoilTEX’s special formatting commands that have no relevance when constructing an HTML page.

6Thanks to John Turner turner@lanl.gov for this example, which was used in developing code to handle minipage environments correctly.

Any options given on the \documentclass or \documentstyle line may also cause a file containing Perl code to be loaded. Such a file is named <option>.perl for the appropriate <option>. When such a file exists, in the local directory or in the styles/ directory, it typically contains Perl code to define subroutines or set values for variables that will affect how certain translations are performed. There can be code that is executed only for specific document-classes.

Since the files for class-options are loaded after those for the document-class, it is possible for the <option>.perl file to contain code that overrides settings made within the document-class file.

If a file named <class> <option>.perl happens to exist for a given combination of document-class <class> and class-option <option>, then this will be loaded. When such a file exists, reading and executing its contents is done, rather than executing any <class> <option> specific information that may be contained in <class>.perl or

<option>.perl .

Currently there are no special option or class-option files provided with the L^{A}TEX2HTML distribution. It is hoped that users will identify ways that specific features can be improved or adapted to specific classes of documents, and will write such files themselves, perhaps submitting them for general distribution.

Note: This mechanism for handling code specific to different document classes and class-options is more general than that employed by L^{A}TEX 2ε. New options can be defined for document-classes generally, or for specific classes, without the need to have corresponding .sty or .clo files. L^{A}TEX simply notes the existence of unusupported options—processing is not interrupted.

3.7 Packages and Style-Files

Similar to the document-class mechanism described in Section3.6, L^{A}TEX2HTML provides a mechanism whereby the code to translate specific packages and style-files is automatically loaded, if such code is available. For example, when use of a style such as german.sty is detected in a L^{A}TEX source document, either by

• a \usepackage command of L^{A}TEX 2ε;

• an option to the \documentstyle command of L^{A}TEX 2.09 ;

• an explicit \input or \include command;

the translator looks for a corresponding .perl file having the same file-name prefix; e.g. the file $LATEX2HTMLDIR/styles/german.perl. If such a .perl file is found, then its code will be incorporated with the main script, to be used as required.

This mechanism helps to keep the core script smaller, as well as making it easier for others to contribute and share solutions on how to translate specific style-files. The current distribution includes the files to support the styles listed in Table 6. These provide good examples of how you can create further extensions to L^{A}TEX2HTML.

Table 6: Supported L^{A}TEX2HTML packages and style-files.

.perl file  Description
alltt  Supports the L^{A}TEX 2ε’s alltt package.
amsfonts  provides recognition of the special AMS font symbols.
amsmath  same as amstex.perl.
amssymb  same as amsfonts.perl.
amstex  Supports much of the AMS-L^{A}TEX package (not yet complete).
babel  Interface to german.perl via the babel package.
changebar Provides rudimentary change-bar support.
chemsym  defines the standard atomic symbols.
color  Causes colored text to be processed as ordinary text by L^{A}TEX2HTML.
colordvi  supports the Crayola colors.
enumerate supports structured labels for enumerate environments.
epsbox  Processes embedded figures not enclosed in a figure environment.
epsfig  Processes embedded figures not enclosed in a figure environment.
finnish  Support for the Finnish language.
floatfig  Processes floating figures.
floatflt  Processes floating figures and tables.
foils  Supports FoilTEX system.
frames  Provides separate frames for navigation and footnotes.
francais  Support for the French language, same as french.perl.
french  Support for the French language.
german  Support for the German language.
germanb  Support for the German language, same as german.perl.
graphics  Supports commands in the graphics package.
graphicx  Supports the alternate syntax of graphics commands.
harvard  Supports the harvard style of citation (same as fnnharvard.perl).
heqn  Alters the way displayed equations are processed.
hthtml  gives an alternative syntax for specifying hyperlinks, etc.
htmllist  Provides support for fancy lists.
justify  supports paragraph alignment—no longer needed.
latexsym  supports the L^{A}TEX symbol font.
lgrind  macros for nice layout of computer program code.
longtable  supports use of long tables, as a single table.
makeidx  provides more sophisticated indexing.
multicol  suppresses requests for multi-columns.
natbib  Supports many different styles for citations and bibliographies.
nharvard  Supports harvard-style citations, using natbib.
seminar  for creation of overhead-presentation slides.
spanish  Support for the Spanish language.
supertabular supports use super-tables, as an ordinary table.
texdefs  Supports some raw TEX commands.
verbatim  Supports verbatim input of files.
verbatimfiles Supports verbatim input of files, also with line-numbering.
wrapfig  Supports wrapped figures.
xspace  Supports use of the xspace package and \xspace command.
xy  Supports use of the XY-pic graphics package.

The problem however, is that writing such extensions requires an understanding of Perl programming and of the way the processing in L^{A}TEX2HTML is organised. Interfaces that are more “user-friendly” are being investigated. Some of the techniques currently used are explained in Section 5.3.

3.7.1 Fancy List-Markers

An optional style-file htmllist.sty has been provided which produces fancier lists in the electronic version of the documentsuch as this. This file defines a new L^{A}TEX environment htmllist, which causes a user-defined item-mark to be placed at each new item of the list, and which causes the optional description to be displayed in bold letters. The filename prefix for the item-mark image can be given as an optional parameter; see example below. The images distributed with L^{A}TEX2HTML for this purpose are listed with the description of the \htmlitemmark command, which provides an alternative means of choosing the item-mark, and allows the image to be changed for different items in the list.

The mark is determined by the \htmlitemmark{<item-mark>} command. This command accepts either a mnemonic name for the <item-mark>, from a list of icons established at installation, or the URL of a mark not in the installation list. The command \htmlitemmark must be used inside the htmllist environment in order to be effective, and it may be used more than once to change the mark within the list. The item-marks supplied

with L^{A}TEX2HTML are BlueBall, RedBall, OrangeBall, GreenBall, PinkBall, PurpleBall,

WhiteBall and YellowBall. The htmllist environment is identical to the description environment in the printed version. An example of its usage is:

    \begin{htmllist}[WhiteBall]
    \item[Item 1:] This will have a white ball.
    \item[Item 2:] This will also have a white ball.
    \htmlitemmark{RedBall}%
    \item[Item 3:] This will have a red ball.
    \end{htmllist}

This will produce:

Item 1: This will have a white ball.

Item 2: This will also have a white ball.

Item 3: This will have a red ball.

One can also obtain L^{A}TEX 2ε style-files floatfig.sty and wrapfig.sty, which provide support for the floatingfigure and wrapfigure environments, respectively. These environments allow text to wrap around a figure in the printed version, but are treated exactly as an ordinary figures in the electronic version. They are described in The L^{A}TEX Companion[2].

3.7.2 Support for FoilTEX

The FoilTEX system presents some additional problems for L^{A}TEX2HTML:

• It has additional commands like \foilhead and \rotatefoilhead, that roughly correspond to sectioning commands,

• The images are produced at the sizes suitable for large screen presentation, but not for the HTML.

The package foils.perl deals with these problems. It treats foils as starred subsections and ignores FoilTEX-specific commands that have no meaning for HTML, like \LogoOn. The header \documentclass[+options]{foils} in the images.tex file is substituted by the

header \documentclass[$FOILOPTIONS]{$FOILCLASS}, where the variables $FOILOPTIONS

and $FOILCLASS can be set in the configuration file (by default they are '10pt' and 'article' correspondingly). A further variable $FOILHEADLEVEL holds the level of sectioning at which a ‘foil’ is to correspond; the default level is 4 (sub-section).

The L^{A}TEX style file foilhtml.sty in the texinputs/ directory provides some additional features for FoilTEX. It implements structural markup commands like \section, \tableofcontents for foils. See the directory docs/foilhtml/ for the details.

3.7.3 Indicating Differences between Document Versions

L^{A}TEX2HTML supports the L^{A}TEX 2ε changebar.sty package, written by Johannes Braams JLBraams@cistron.nl, for inserting change-bars in a document in order to indicate differences from previous versions. This is a very primitive form of version control and there is much scope for improvement.

Within the L^{A}TEX version of this manual two thicknesses of change-bar have been used. Thicker bars indicate changes introduced with version v97.1 , while thinner bars indicate earlier additions since v96.1 . Within the HTML version the change-bars clearly indicate the different revisions with explicit numbering.Within the HTML version, the graphic icons representing the changebars can be followed by some text indicating the new version. This is used repeatedly throughout the online version of this manual. It is achieved using the

command \cbversion{<version>}, immediately following the \begin{changebar}. This

sets a variable $cb version to be used both at the beginning and end of the environment. The value of this variable is retained, to be used with other changebar environments, unless changed explicitly by another occurrence of $cb version. Warning: L^{A}TEX2HTML will not correctly process changebar environments that contain sectioning commands, even when the (sub)sections or (sub)paragraphs are to occur on the same HTML page. If this is required, use a separate changebar environment within each (sub)section or (sub)paragraph.

3.8 Indexing

L^{A}TEX2HTML automatically produces an Index consisting of the arguments to all \index commands encountered, if there are any. A hyperlink is created to that point in the text where the \index command occurred.

More sophisticated indexing is available by loading the makeidx package. Most of the features described in [1, Appendix A] become available. This includes:

styled entries, using ‘@’ : Entries of the form \index{<sort-key>@<styled-text>} produce <styled-text> as the entry, but sorted according to <sort-key>.

hierarchical entries, using ‘!’ : Entries of the form \index{<item>!<sub-item>} set the <sub-item> indented below the <item>. Unlimited levels of hierarchy are possible, even though L^{A}TEX is limited to only 3 levels. The <sort-key>@<styled-text> can be used at each level.

explicit ranges, using ‘|(’ and ‘|)’ : This is perhaps more useful in the L^{A}TEX version. In the HTML version these simply insert words “from” and “to”, respectively, prior to the hyperlink to where the index-entry occurs.

|see{<index-entry>} : provides a textual reference to another indexed word or phrase, by inserting the word “see”. This can be used in conjunction with \htmlref to create a hyperlink to the <index-entry>; viz.

        \index{latexe@\LaTeXe |see{\htmlref{\LaTeX}{IIIlatex}}}

where a \label has been specified in some other index-entry, as follows:

        \index{latex@\LaTeX\label{IIIlatex}}

|emph : is handled correctly, by applying \emph to the text of the generated hyperlink.

|<style> : where <style> is the name of L^{A}TEX style-changing command, without the initial ‘\’; e.g. ‘emph’, ‘textbf’, ‘textit’, etc. The corresponding L^{A}TEX command is applied to the text of the generated hyperlink.

blank lines and alphabetization: Having precisely a single space-character after the | (e.g. \index{A| }) places a blank line before the index entry and omits the hyperlink. This is used mainly for visual formatting; it allows a break before the entries starting with each letter, say. Using a printable-key, as in \index{Q@Q, R| }, is appropriate when there are no indexed words starting with ‘Q’, say.

quoted delimiters: The three special delimiters can be used within the printable portion, if preceded by the double-quote character: "@, "|, "! and also "" for the quote character itself. Also \" produces an umlaut accent on the following character, when appropriate, else is ignored.

Furthermore, the printable part of an index entry can contain HTML anchors; that is, hyperlinks and/or \label{...}s. This allows index entries to contain cross-links to other entries, for example, as well as allowing index-entries to be the target of hyperlinks from elsewhere within the document.

The next section describes how this feature is used within this manual to create a Glossary, containing a short description of all file-names, configuration-variables and application software mentioned within the manual, integrated with the Index. All occurrences of the technical names can be easily found, starting from any other.

When a single item is indexed many times, it is sufficient to have a \label command appearing within the printable portion of the first instance of an \index{...} command for that item, within a single document segment.

If the index-entries are in different segments of a segmented document, it is sufficient to have the \index{...@...\label{...}} appearing within that segment, in which the item is indexed, whose indexing information is loaded earliest via a \internal[index]{...} command. When in doubt, include one \index{...@...\label{...}} per segment in which the item is indexed.

For cross-links to work effectively within segmented documents, the indexing command \index{...@...\label{...}} must occur earlier in the same segment than any use of

\index{...@...\htmlref{...}{...}} intended to create a link to that label. If the \label

occurs in a different segment, then a \internal[index]{...} command for that segment, may be needed at the beginning of the segment with the \htmlref . When this is done incorrectly, the resulting link will be to the segment where the indexed item occurred, rather than staying within the Index.

Since use of section-names, as the text for hyperlinks, can lead to a very long and cumbersome Index, especially when single items have been indexed many times, a further feature is provided to obtain a more compact Index.

Use of the command-line option ‘ -short index ’ causes a codified representation of the sectioning to be used, rather than the full section-name. The differences are as follows.

• For example, ‘2.1’ means sub-node #1 of node #2, viewing the entire document as a tree-like structure.

• The top-most node is simply denoted ‘^’.

• With a segmented document, each segment is codified separately using the <prefix> supplied for that segment. The Index includes a legend of these prefixes, each giving the title of the leading page from the segment, as a hyperlink to the place on that page where its child-links are displayed.

• Hyperlinks start on the same line as the index-key, rather than the next line, separated by ‘|’. This gives further compactification for easier browsing.

• If ‘ -prefix <prefix> ’ has been specified, then the <prefix> is prepended to the codified form. This is most useful for segmented documents. Now the top-most node is indicated by the bare <prefix>.

These features can also be obtained by setting the variable $SHORT INDEX to have value ‘1’, in a configuration or initialisation file; provided, of course, that the document loads the makeidx package.

3.8.1 Integrated Glossary and Index

A large number of different pieces of software are required to make L^{A}TEX2HTML work effectively, as well as many files containing data or code to work with parts of this software. For this reason, a Glossary is included with this manual. It contains the names of all files, configuration variables, application software and related technical terms, with a short description of what it is, or does, and perhaps a URL for further reference.

In the printed version each item in the Glossary is accompanied by the page-numbers on which the item is mentioned, somewhat like in the Index. For the HTML version, each glossary-item contains a hyperlink to an index-entry, which then has links to each occurrence. These extra index-entries do not appear in the printed version; indeed they also contain a hyperlink back to the corresponding glossary-entry.

This feature is currently available only when using the makeidx package, and needs also the html and htmllist packages. It was developed for version 96.1f by Ross Moore, incorporating an extensive revision of makeidx.perl, as well as additions to L^{A}TEX2HTML so that all aspects of indexing work correctly with segmented documents.

Since L^{A}TEX provides no guidelines for how a Glossary should be constructed, the technique used here will be explained in detail, for both the printed and HTML versions.

• Firstly the \makeglossary command, which is similar to \makeindex, must appear in the document preamble, so that L^{A}TEX will record uses of the \glossary{...} command within a file manual.glo.

This command is redundant in the HTML version, so is given a trivial definition which is ignored by L^{A}TEX.

• Next, the words, phrases or technical terms to be included in the Glossary are marked in the main text using the \glossary command, used indirectly via other macros. For example, file-names are inserted via \fn{html.sty}, \fn{dvips}, \appl{dvips} etc. which both insert the text and create the glossary-entry; viz.

    \newcommand{\fn}[1]{\htmlref{\texttt{#1}}{GGG#1}\glossary{#1}}
    \newcommand{\appl}[1]{\htmlref{\textsl{#1}}{GGG#1}%
      \Glossary{#1}{\textsl{#1}}}

• The expansions of \glossary, and the slightly more general \Glossary, are different for the printed and HTML versions. For the HTML version the following definitions occur within an htmlonly environment:

    \def\glossary#1{\index{#1@\texttt{#1} \label{III#1}%
      \htmlref{(G)}{GGG#1}}}
    \def\Glossary#1#2{\index{#1@{#2} \label{III#1}\htmlref{(G)}{GGG#1}}}
    \def\makeglossary{}
. . . while in L^{A}TEX we need only: \newcommand\Glossary[2]{\glossary{#1@#2}} .

Notice how the feature of makeidx, allowing the printable portion to be separate from the sorting-key, is used to allow text-styles to be included within both index-entries and glossary-entries. Indeed the purpose of \Glossary is to allow deviations from a fixed style, e.g.

    \newcommand{\MF}{\htmlref{\textsl{Metafont}}{GGGmetafont}%
      \Glossary{metafont}{\textsl{Metafont}}}%

Also notice that in the HTML version an index-entry is created that includes, within its printable portion, both a \label and a hyperlink. The former, having name III..., will ultimately reside on the Index page, while the latter will point to an anchor named GGG... on the Glossary page. These names must be distinct from any other names used with \labels elsewhere in the document, hence the use of prefixes III and GGG. A short string ‘(G)’ is used for the text of the hyperlink in the Index.

• The text descriptions of the glossary-items are stored in a file called l2hfiles.dat, with one description per line. For the HTML version this file is actually read as input:

    \section*{Glossary of variables and file-names\label{Glossary}}
    \begin{htmllist}\htmlitemmark{OrangeBall}
      \input l2hfiles.dat
    \end{htmllist}

For this reason alone it is desirable to have l2hfiles.dat sorted alphabetically.

• The mechanism used for the L^{A}TEX version also requires the file to be sorted strictly alphabetically, according to the sort-keys associated to each glossary entry. (This requirement could be relaxed, but only with a loss in efficiency; see below.) L^{A}TEX constructs its Glossary by running the makeindex utility on the file manual.glo, using the following command:

    makeindex -o manual.gls -s l2hglo.ist manual.glo

Its output, which includes page numbering for an index, is stored in manual.gls and subsequently read by L^{A}TEX using:

    \InputIfFileExists{manual.gls}{\clearpage\typeout{^^Jcreating Glossary...}}
    {\typeout{^^JNo Glossary, since  manual.gls  could not be found.^^J}}

The configuration file l2hglo.ist is included along with this manual. It contains a portion that inserts tricky TEX code at the beginning of manual.gls. This code extracts from l2hfiles.dat that line corresponding to each glossary entry, then typesets it itemized within an environment called theglossary.

    \newenvironment{theglossary}{\begin{list}{}{%
      \setlength{\labelwidth}{20pt}%
      \setlength{\leftmargin}{\labelwidth}%
      \setlength\itemindent{-\labelwidth}%
      \setlength\itemsep{0pt}\setlength\parsep{0pt}%
      \rmfamily}}{\end{list}}

Currently searching within l2hfiles.dat is only done sequentially, stopping at the end of the file. If an entry is not found then it is skipped and a message printed to the log; the next entry will search from the top of the file. If all entries are included and maintained in strict order, there will be no skipping and each line of l2hfiles.dat is read exactly once.

• Within l2hfiles.dat the data lines look like:

    \item[\gn{french.perl}] adds \Perl{} code to be compatible with the ...
    \item[\gn{\textsl  {ftp}}] `File Transfer Protocols', network ...
    \item[\gn{german.perl}] adds \Perl{} code to be compatible with the ...
    ...

For the L^{A}TEX version the \item[\gn{...}] is only used for pattern-matching, to find the correct data entry. All typesetting is controlled from within manual.gls. However the HTML version requires the following definition:

    \newcommand{\gn}[1]{\texttt{#1}\label{GGG#1}\htmlref{\^}{III#1}}%

which establishes the hyperlink to the Index, marked by ‘^’, and provides the \label to create the target in the Glossary for any \glossary{...} command having the corresponding argument.

4 Hypertext Extensions to L^{A}TEX

This section describes how you can define hypertext entries in your HTML documents from within your L^{A}TEX source, as well as other effects available in HTML for which there need be no direct L^{A}TEX analog for a printed document. These are implemented as new L^{A}TEX commands which have special meaning during the translation by L^{A}TEX2HTML into HTML, but are mostly ignored when processed by L^{A}TEX.

The new commands described in the sections below are defined mainly in the html package, with L^{A}TEX definitions in the file html.sty, which is part of the L^{A}TEX2HTML distribution. It must be included in any L^{A}TEX document using these features, by one of the following methods:

• including html as an optional argument to \documentstyle in L^{A}TEX 2.09 ;

• including html in a L^{A}TEX 2ε \usepackage command.

It is not sufficient to load the style file via an \input or \include command, such as \input html.sty . This will load the required definitions for L^{A}TEX, but will not load the html.perl package file for L^{A}TEX2HTML. Warning: Some of these features, but not all, are also available with L^{A}TEX 2.09. Users of L^{A}TEX2HTML are strongly advised to upgrade their L^{A}TEX installations to L^{A}TEX 2ε.

Several new environments are defined, in particular for specifying large (or small) sections of the text which are appropriate to only one version of the document—either the HTML or the L^{A}TEX typeset version. Their use is discussed in Sections 4.2 and 4.4.

\begin{rawhtml} for including raw HTML tags and SGML-like markup.

\begin{htmlonly} for material intended for the HTML pages only.

\begin{latexonly} for material intended for the L^{A}TEX version only. Note that any macrodefinitions or changes to counter-values are local to within this environment.

%begin{latexonly} for material intended for the L^{A}TEX version only. Macro-definitions and changes to counter-values are retained outside of this (pseudo-)environment.

\begin{imagesonly} for material intended to be used in the images.tex file only.

\begin{comment} for user-comments only, currently ignored in both the HTML and L^{A}TEX versions. (To put HTML comments into the HTML files, use the rawhtml environment.)

\begin{makeimage} creates an image of its contents, as typeset by L^{A}TEX. This is also used to prevent an image being made of the complete contents of a figure environment, allowing more natural processing.

\begin{htmllist} defined in htmllist.sty and htmllist.perl, this produces coloured

balls tagging the items in a descriptive list, as used throughout the HTML version of this manual.

Warning: When using these environments it is important that the closing delimiter, \end{htmlonly} say, occurs on a line by itself with no preceding spaces, <tab>s or any other characters. (Otherwise L^{A}TEX will not recognise the intended end of the environment when processing for the .dvi version.) Similarly there should be nothing on the same line after the opening environment delimiter, \begin{htmlonly} say.

The following commands are defined for L^{A}TEX in html.sty. Corresponding Perl implementations are either in html.perl or in the latex2html script itself.

\latextohtml expands to the name L^{A}TEX2HTML, of this translator;

\htmladdnormallink creates a (perhaps named) textual hyperlink to a specified <URL>;

\htmladdnormallinkfoot same as \htmladdnormallink, but L^{A}TEX also prints the <URL>

in a footnote;

\htmladdimg places an image (perhaps aligned) on the HTML page; ignored by L^{A}TEX.

\hyperref creates a textual hyperlink to where a \label command occurred within the same document. This is the recommended substitute for L^{A}TEX’s \ref command.

\htmlref creates a textual hyperlink to the place where a \label command occurred; no reference is printed in the L^{A}TEX version.

\hypercite creates a textual hyperlink to the bibliography page where citation details are shown. This is the recommended substitute for L^{A}TEX’s \cite command.

\htmlcite creates a textual hyperlink to the bibliography page where citation details are shown; no citation marker is printed in the L^{A}TEX version.

\externalref creates a textual hyperlink to where a \label command occurred within a different document that has also been processed by L^{A}TEX2HTML; ignored in L^{A}TEX.

\externalcite creates a textual hyperlink to where a reference occurs in a bibliography page from a different document that has also been processed by L^{A}TEX2HTML; ignored in L^{A}TEX.

\externallabels allows hypertext links to a different document; ignored in L^{A}TEX.

The following commands, also defined for L^{A}TEX in html.sty, are normally used only when creating segmented documents, see Section 4.10.

\segment directs that an \input file <file> should be regarded as a separate “segment” of a larger L^{A}TEX2HTML document. In L^{A}TEX the file is input as usual, after counter values have first been written to a file, named <file>.ptr .

\startdocument tells L^{A}TEX2HTML where the end of the preamble occurs for a document segment; ignored in L^{A}TEX. (A segment cannot have a \begin{document} command, unless it is shielded from L^{A}TEX within an htmlonly environment.)

\internal reads internal information from another document, so that symbolic references can be treated as if part of the current document; ignored in L^{A}TEX.

\htmlhead places a sectional heading on a HTML page; used mainly with the document segmentation feature. It is ignored in L^{A}TEX.

\htmlnohead suppresses the section-heading for a document segment; ignored in L^{A}TEX.

\segmentcolor read from the .ptr file, this sets the text color for a document segment; ignored in L^{A}TEX.

\segmentpagecolor read from the .ptr file, this sets the background color for a document segment; ignored in L^{A}TEX.

The following commands are shorthand forms for some of the “conditional” environments listed above.

\html for putting small pieces of text into the HTML version only;

\latex for putting small pieces of text into the L^{A}TEX version only;

\latexhtml puts one piece of text into the L^{A}TEX version, another into the HTML version.

The following commands implement effects on the HTML pages for which there is no direct L^{A}TEX counterpart. Most of these commands are discussed in detail in Section 4.8.

\HTMLcode a general command for placing raw HTML tags, with attributes and contents; tags and attributes are ignored in L^{A}TEX, but not the contents. (See Section 38.)

\htmlrule places a (perhaps styled) horizontal line on the HTML page; ignored in L^{A}TEX.

\strikeout places text between <STRIKE>...</STRIKE> tags; ignored in L^{A}TEX.

\htmlimage used for fine control over the size of individual images, and other graphics effects (e.g. making a ‘thumbnail’ version); ignored in L^{A}TEX. (See page 18 for details.)

\htmlborder places a border around the contents of an environment, but placing the environment as a cell inside a <TABLE>; ignored in L^{A}TEX.

\tableofchildlinks determines where the table of childlinks should be placed on the HTML page; ignored in L^{A}TEX.

\htmlinfo determines where the “About this document...” information should be placed; ignored in L^{A}TEX.

\htmladdtonavigation appends a button to the navigation panels; ignored in L^{A}TEX.

\bodytext allows the contents of the <BODY...> tag to be set explicitly for the current and subsequent HTML pages; ignored in L^{A}TEX.

\htmlbody allows an attribute to be added or changed within the <BODY...> tag of HTML; ignored in L^{A}TEX.

\htmlbase Allows a URL to be specified within the <BASE...> tag for all the HTML pages produced; ignored in L^{A}TEX.

\htmltracing{<level>} specifies that extra tracing messages be generated, according to the <level>; ignored in L^{A}TEX. (See page 67 for levels of verbosity.)

\htmltracenv{<level>} same as \htmltracing except that this command is evaluated in sequence with environments; ignored in L^{A}TEX. (See also page 67.)

\HTMLset programmer’s device, allowing an arbitrary Perl variable to be set or changed dynamically during the L^{A}TEX2HTML processing; ignored in L^{A}TEX.

\HTMLsetenv Same as the preceding \HTMLset command, except that this one is processed in order, as if it were an environment; ignored in L^{A}TEX.

Most of the new environments listed above can also be used with delimiter macros \<env-name>...\end<env-name>. This alternative style, which is common with AMS-TEX, is discouraged for general L^{A}TEX usage (even by the AMS itself) in favour of the usual

\begin{<env-name>}...\end{<env-name>} markup notation. (Safety features that are

available with the usual \begin...\end mechanism may not always work in the best way with this alternative style of environment delimiter. These comments apply to both the L^{A}TEX and L^{A}TEX2HTML processing.)

\rawhtml...\endrawhtml old AMS-style variant of rawhtml environment.

\htmlonly...\endhtmlonly old AMS-style variant of htmlonly environment.

\latexonly...\endlatexonly old AMS-style variant of latexonly environment.

\imagesonly...\endimagesonly old AMS-style variant of imagesonly environment.

\comment...\endcomment old AMS-style variant of comment environment.

Warning: These ‘pseudo’-environments are not as reliable as their L^{A}TEX counterparts. In particular, the \begin<env-name> and \end<env-name> commands should appear on lines by themselves, preferably with no preceding spaces or <tab> characters. This requirement is analogous to the warning at the bottom of page 33 for conditional environments.

4.1 Hyper-links in L^{A}TEX

Arbitrary hypertext references are created using the \htmladdnormallink and \htmladdimg commands. These have syntax:

     \htmladdnormallink{<text>}{<URL>}
     \htmladdnormallink[<name>]{<text>}{<URL>}
     \htmladdimg{<URL>}
     \htmladdimg[<align>]<URL>}
     \htmladdnormallinkfoot{<text>}{<URL>}
     \htmladdnormallinkfoot[<name>]{<text>}{<URL>}

\htmladdnormallink The \htmladdnormallink command expects some text as the first

argument and a URL as the second argument. When processed by L^{A}TEX (i.e. in the .dvi or .ps output files), the URL will have no effect. But when processed by the translator, the URL will be used to provide an active hypertext link (to another file, picture, sound-file, movie, etc.) e.g.

     \htmladdnormallink{<URL>}
       {http://www.ncsa.uiuc.edu/demoweb/url-primer.html}

The optional argument to \htmladdnormallink allows a name to be specified for the place in the document where the hyperlink occurs. This is done via the NAME="<name>" attribute for the <A...> anchor tag in HTML . Such a name can be used as the target for a hyperlink using the \htmlref command, described in Section 4.5.

\htmladdimg In a similar way, the argument of the \htmladdimg command should be a URL pointing to an image. This URL is ignored in the L^{A}TEX hard copy output. The optional argument to \htmladdimg allows an alignment for the image to be given: center, right or left. In the latter cases, the image is bound to the specified side of the browser’s window. Subsequent text paragraphs ‘flow around’ the other side of the image.

In fact any valid set of “attributes” for the <IMG> tag in HTML can be specified as the optional <align> parameter. In particular the WIDTH, HEIGHT and BORDER attributes can be set, perhaps overriding the natural size of the image.

\htmladdnormallinkfoot The \htmladdnormallinkfoot command takes the same argu-

ments, and when generating HTML has the same effect, as \htmladdnormallink. However when processed by L^{A}TEX it places the URL as a footnote.

Warning: The tilde (~) character is commonly used within hyperlink URLs. It is a quirk of TEX and L^{A}TEX that it must be generated via \~{}, else the ~ will be interpreted as an accent on the following character.

4.2 Including Arbitrary HTML Mark-up and Comments

L^{A}TEX2HTML provides the ability to include raw HTML tags and text within the HTML version of a document, without requiring corresponding material for the L^{A}TEX typeset version. This ability can be used to

• include HTML markup for effects that have no corresponding concept within a L^{A}TEX typeset document (see the following example)

• take advantage of new HTML facilities as soon as they become available, and there are browsers capable of displaying them.

• include arbitrary SGML-like markup, for use with special browsers that know how to sensibly handle the resulting files.

\begin{rawhtml} The simplest way to include raw HTML tags and/or text is by using the rawhtml environment. (An alternative way is to use the \HTML command, described in Section 4.3, which allows macros to be expanded to give the required tags, attributes and contents.) Note the warning on page 33 concerning how the environment delimiters should be used in the L^{A}TEX source code.

A particularly good use of the rawhtml environment is in the creation of interactive electronic forms from within a L^{A}TEX document. When producing the paper (.dvi) version of a document the rawhtml environment is ignored. Here is an example:

    \begin{rawhtml}
    <HR>
    <FORM ACTION="http://cbl.leeds.ac.uk/nikos/doc/error.html">
    <OL>
    <LI> <INPUT TYPE="checkbox" NAME="wp" VALUE="word"> Word for
    Windows.
    <LI> <INPUT TYPE="checkbox" NAME="wp" VALUE="wp"> Word Perfect.
    <LI> <INPUT TYPE="checkbox" NAME="wp" VALUE="latex"> LaTeX.
    <LI> Plain Text Editors (Please Specify): <INPUT TYPE="text" NAME="other_ed">
    </OL>
    So, what do think (comments please): <BR>
    <INPUT TYPE="text" SIZE=45 NAME="other_wp">
    <INPUT TYPE="submit" VALUE="submit this form but don't expect much!">
    </FORM>
    <HR>
    \end{rawhtml}

The result is shown in Figure 5.

Figure 5: An electronic form. In the online version the form would be active.

\beginrawhtml...\endrawhtml This is an alternative way to specify a chunk of raw HTML code, using the old AMS-style of delimiting environments. Use of this style is discouraged; the rawhtml environment is preferred.

\begin{comment} This environment is simple for the convenience of “commenting-out” large sections of source code. The contents of this environment is completely ignored, both in the L^{A}TEX and HTML versions. Such an environment is already used in AMS-L^{A}TEX, and perhaps with other packages. It is defined here for its general utility. To insert SGML-style comments into the HTML files, use the rawhtml environment as follows.

    \begin{rawhtml}
    <!--  this text is treated as a comment
          perhaps extending over several lines
    -->
    \end{rawhtml}

Note the warning on page 33 concerning how the environment delimiters should be used in the L^{A}TEX source code.

\comment...\endcomment This is an alternative way to specify a chunk of material intended to be ignored in both the L^{A}TEX and HTML versions, using the old AMS-style of delimiting environments. Use of this style (though convenient for typing) is discouraged, since it is not as reliable as using the comment environment.

4.3 Arbitrary Tags and Attributes

For version 97.1 of L^{A}TEX2HTML there is a new command which provides an extremely flexible way to include HTML 3.2 tags, along with any values for the “attributes” of that tag, if desired.

     \HTMLcode[<attribs>]{<tag>}
     \HTMLcode[<attribs>]{<tag>}{<contents>}

When the <tag> also needs a closing tag (e.g <I>...</I>) the <contents> must be given, enclosed in braces. Both the opening and closing tags then will be placed correctly.

Warning: In version 97.1 this command was actually called \HTML. However style files may well define \HTML to mean something else, like a styled version of the HTML acronym. So in version 98.1 the name has been changed to \HTMLcode.

If no other definition of \HTML exists, then this command will be defined, to work the same as \HTMLcode.

An important aspect of this is that any of the <tag>, <attribs> and <contents> may be given wholly by expanding a L^{A}TEX macro, or may contain arbitrary macros, perhaps including other \HTMLcode commands. The contents of Figure 6 was constructed using this feature; its L^{A}TEX source follows.

Figure 6: Example use of macros for raw HTML code.

\newcommand{\myalign}{center}
\newcommand{\mylist}{UL}
\newcommand{\myitem}[2]{\HTMLcode[disc]{LI}{\simpletest{#1}{#2}}}
\newcommand{\simpletest}[2]{%
\HTMLcode{#1}{ a simple test of ``#2'',} using \HTMLcode{CODE}{<#1>} .}
\newcommand{\tableopts}{10,border=5}
\newcommand{\tablelist}[4][left]{\HTMLcode[#1]{DIV}{
\HTMLcode[\tableopts]{TABLE}{
\HTMLcode[bottom]{CAPTION}{
#3
}\HTMLcode{TR}{\HTMLcode{TD}{
\HTMLcode{#2}{
#4

}}}
}}\HTMLcode[all]{BR}}
\tablelist[\myalign]{\mylist}{%
\textbf{A listing of the different text styles available in HTML 3.2}}{%
\myitem{B}{bold-face}
\myitem{I}{italics}
\myitem{TT}{teletype-text}
\myitem{U}{underlining}
\HTMLcode[circle]{LI}{\simpletest{STRIKE}{strikeout}}
\myitem{EM}{emphasis style}
\myitem{STRONG}{strong style}
\myitem{CODE}{code style}
\myitem{CITE}{citation style}
\myitem{DFN}{definition style}
\HTMLcode[square]{LI}{\simpletest{SAMP}{sample style}}
\HTMLcode[square]{LI}{\simpletest{KBD}{keyboard style}}
\myitem{VAR}{variable style}}

The above code demonstrates many aspects of the way \HTML commands can be used.

nesting: \HTML commands can be nested to arbitrary depth.

macros: Macros can be used to specify all or part of each argument.

within macros: \HTMLcode commands work correctly within the expansions of other macros.

attribute values: Information within <attribs> can be specified in a very loose way, as a comma-separated list of key/value pairs or as single values. Not even the commas are necessary: space(s), <tab>s or newlines are equally effective. Indeed the horizontal rules preceding and following the table were specified by:

        \HTMLcode[50\% 3 noshade center]{HR}

attribute names: Usually it is not necessary to know the names of the attributes to the tags that are to be used. It is sufficient just to give the values; these will be matched to the appropriate attribute, according to the type of data required. (If names are given, these are case-insensitive.)

newlines: Although L^{A}TEX ignores linebreaks within the source code, this is not so with L^{A}TEX2HTML. The strange spreading-out of the definition of the \tablelist command above was done with the purpose solely of making the code in the resulting HTML files more easily readable, to a human. (As most browsers ignore those newlines anyway, more compact code would have rendered the same on-screen.)

Some further aspects of the use of this \HTML command are not apparent from the above example.

invalid <tag> : If a <tag> is specified that is not part of the HTML 3.2 specifications, then it and its attributes are not placed into the HTML document created by L^{A}TEX2HTML. Any <contents> is included as ordinary data; i.e. as text in paragraphs, etc. required attributes: Some tags have attributes which are required to have values, if that

tag is to be included in an HTML document. Using the \HTML command, if any such attribute is not given an appropriate value then the tag is ignored. Any <contents> are included in the document, as ordinary character data.

valid HTML : Currently there is no checking that the <contents> of a <tag> contains only data (perhaps including other tags) allowed by the DTD for HTML 3.2.

The requirement to produce valid HTML currently rests with the user.

This issue will be addressed in forthcoming revisions of L^{A}TEX2HTML.

extra attributes and values: The list of attributes for a <tag> can include key-value pairs whose keys do not match any valid attribute for the <tag>. Such key-value pairs are simply ignored. Similarly extra data values are ignored, as are values that do not match the requirements for any valid attribute.

attributes with similar data-types: Several attributes to a <tag> may use values having the same or similar data-types. First any key-value pairs are processed. Remaining values are allocated to those attributes which do not already have a value. An ordering of the attributes is used, based on a perceived likelihood of each attribute being required to be changed from its default setting.

4.4 Conditional Text

\begin{latexonly} and \begin{htmlonly} Conditional text can be specified using the environments latexonly and htmlonly. These allow writing parts of a document which are intended only for electronic delivery or only for paper-based delivery.

This would be useful for example in adding a long description of a multi-media resource in the paper version of a document. Such a description would be redundant in the electronic version, as the user can have direct access to this resource.

Here is an example of the use of the latexonly environment, used on page 37 of this manual:

   \begin{latexonly}
    \begin{figure}
        \begin{center}
        \fbox{\includegraphics[width=4in]{psfiles/eform}}
        \end{center}
        \caption{An electronic form. Of course in the online version of this
         document the form above would be active.}
    \end{figure}
   \end{latexonly}

Note the warning at the bottom of page 33 concerning how the environment delimiters should be used in the L^{A}TEX source code.

\htmlonly...\endhtmlonly This is an alternative way to specify a chunk of material intended for the HTML version only, using the old AMS-style of delimiting environments. Use of this style is discouraged; the htmlonly environment is preferred.

\latexonly...\endlatexonly This is an alternative way to specify a chunk of material intended for the L^{A}TEX typeset version only, using the old AMS-style of delimiting environments. Use of this style is discouraged; the latexonly environment or the unscoped %begin{latexonly} construction are preferred. Note the warning at the bottom of page 33 concerning how the environment delimiters should be used in the L^{A}TEX source code.

\latex, \html and \latexhtml There are also shorthand notations to accomplish the same thing as in the latexonly environment and htmlonly environment, but with less typing.

• The \latex{...} command causes everything within the braces to be processed by L^{A}TEX, but ignored by L^{A}TEX2HTML.

• Conversely, the \html{...} command causes everything within the braces to be ignored by L^{A}TEX and processed by L^{A}TEX2HTML.

• Finally the command \latexhtml{...}{...} causes everything within the first set of braces to be processed exclusively by L^{A}TEX, with the contents of the second set of braces processed solely by L^{A}TEX2HTML.

Warning: Only small pieces of text work reliably in this way. With whole paragraphs or contained sub-environments, the “conditional” environments should be used instead.

%begin{latexonly} Another variant of the latexonly environment is available, in which everything between %begin{latexonly} and %end{latexonly} is ignored by L^{A}TEX2HTML. The difference is that the latexonly environment puts the contents into a group, in which all definitions are local. There is no such scoping with the %begin...%end variant, since L^{A}TEX sees the initial %s simply as starting comments.

The following example should clarify what happens:

    \newcommand{\A}{The letter A.}
    \newcommand{\B}{The letter B.}
      \begin{latexonly}
      \renewcommand{\A}{Not the letter A.}
      \end{latexonly}
      %begin{latexonly}
      \renewcommand{\B}{Not the letter B.}
      %end{latexonly}
    \begin{document}
    \A \B
    \end{document}

If you process this with L^{A}TEX, the result is: The letter A. Not the letter B. Note the warning at the bottom of page 33 concerning how the environment delimiters should be used in the L^{A}TEX source code.

Warning: Be careful when using L^{A}TEX commands which alter the values of counters (e.g. numbered figures or equations) in conditional text, because this may cause the counter values in the electronic version to lose synchronisation with the values of the corresponding counters in the L^{A}TEX version.

\begin{imagesonly} This environment is used to put L^{A}TEX code into the images.tex file, to be used when generating images. Typically this is used to add commands to the preamble of images.tex, such as setting the text or background color. However code can be added at any other point as well; e.g. to change the background color of all images after a certain point in the document. Note the warning at the bottom of page 33 concerning how the environment delimiters should be used in the L^{A}TEX source code.

\begin{makeimage} This is a special environment which forces an image to be made of its contents. That is, one gets effectively a snapshot of a portion of a page that has been typeset using L^{A}TEX. Within the normal L^{A}TEX typeset version of the document, this environment is completely transparent, adding its contents to the page as usual.

One further important use of the makeimage environment is as follows. If a makeimage environment occurs as a sub-environment within a figure environment, then an image will not be made of the figure’s contents. Instead, the contents are treated as normal text, each part being handled as if there were no figure at all, except that everything is placed within a single cell of a <TABLE>...</TABLE> construction in HTML 3.2. The contents of any \caption commands are placed between <CAPTION>...</CAPTION> tags for the <TABLE>.

Normally an image of the entire contents of the figure would be placed within the single cell of the <TABLE>. Now images are made of any subparts of those figure’s contents that really need it, in particular the makeimage sub-environments. An empty makeimage subenvironment does not generate an image of itself, yet still it inhibits an image being made of the whole figure. These comments apply also to table environments.

4.5 Symbolic References shown as Hyperized Text

In printed documents cross-references are shown through a numeric or symbolic indirection e.g. “see Figure 1” (numeric indirection), or “see section ‘Changes’ ” (symbolic indirection). L^{A}TEX2HTML can mirror this mechanism using the same numeric or symbolic references, or when these are not appropriate by using iconic references.

In a hypertext document however, cross-references can be shown without any indirection, just by highlighting a relevant piece of text. This can make a document more readable as it removes unnecessary information.

\hyperref A single new L^{A}TEX command \hyperref can be used for specifying how a cross-reference should appear, both in the printed document and in the hypertext version. For example, assuming that the label {sec:cond} is defined somewhere within a document, the command \hyperref, taking 4 arguments, can be used in that document as follows:

    \emph{Is the concept of
    \hyperref
                   % This will be highlighted in the hypertext version
    {conditional text}                      % argument #1
                   % This will be shown in the printed version
                   % followed by a numeric reference ...
    {conditional text (see Section }        % argument #2
                   % ... followed by this text
    { for more information)}                % argument #3
                   % This is the common label
    {sec:cond}                              % argument #4
    a good idea? }

Here is how it will be shown:

Is the concept of conditional text (see Section 4.5 for more information) a good idea?

In the hypertext version what would appear is:

Is the concept of \underline{conditional text} a good idea? (Of course \underline{conditional text} would be an active hypertext link.) An extended syntax for \hyperref uses an optional argument, which determines what information is to be placed in the L^{A}TEX version of the document. The value of this optional argument can also affect the number of required arguments. These forms are recognised:

     \hyperref[ref]{<HTML-text>}{<LaTeX-text>}{<post-LaTeX>}{<label>}
     \hyperref{<HTML-text>}{<LaTeX-text>}{<post-LaTeX>}{<label>}
     \hyperref[pageref]{<HTML-text>}{<LaTeX-text>}{<post-LaTeX>}{<label>}
     \hyperref[page]{<HTML-text>}{<LaTeX-text>}{<post-LaTeX>}{<label>}
     \hyperref[noref]{<HTML-text>}{<LaTeX-text>}{<label>}
     \hyperref[no]{<HTML-text>}{<LaTeX-text>}{<label>}

The first two are the defaults, where L^{A}TEX uses \ref{<label>}. With the next two L^{A}TEX uses \pageref{<label>}, while with the final two L^{A}TEX completely ignores the <label>, setting just the <LaTeX-text>. For creating hyperlinks to other documents using symbolic reference <label>s, see also the \externalref command, described on page 47. The preceding paragraph is an example of the use of the \hyperref[page] option. Its source code is:

    For creating hyperlinks to other documents
    using symbolic reference \Meta{label}s,
    see also the \Lc{externalref}
    \hyperref[page]{command}{command, described on page~}{}{externref}.

which appears in the HTML version as:

For creating hyperlinks to other documents, using symbolic reference <label>s, see also the \externalref \underline{command}.

with the \underline{command} being an active hyperlink. In fact both \hyperref and the \htmlref command, to be described next, permit textual hyperlinks based on symbolic <label>s from external files.

\htmlref Another command also defined in html.sty is \htmlref which has the same effect as \hyperref during the conversion to HTML. It takes two arguments, some text and a label. In the HTML version the text will be “hyperized”, pointing to the label. In the paper version the text will be shown as it is and the label will be ignored; e.g.

    With \verb|\htmlref| \htmlref{it's easy to make links}{fig:example}.

which produces:

With \htmlref it’s easy to make links. In the HTML version it is shown as:

With \htmlref \underline{it’s easy to make links}.

4.6 Hypertext Links in Bibliographic References (Citations)

If a report or a book that is cited (using the \cite command) is available (or there is information about it) on the World-Wide Web, then it is possible to add the appropriate hypertext links in your bibliographic database (the .bib) file.

Here is an example of a bibliographic entry for the original L^{A}TEX [1] blue book:

    @string{tugURL="\htmladdnormallink
    {http://www.tug.org/}{http://www.tug.org}"}
    @string{danteURL="\htmladdnormallink
    {http://www.dante.de/}{http://www.dante.de}"}
    @book{lamp:latex,
    title = "LaTeX User's Guide \& Reference Manual, 2nd edition",
    year = 1994 ,
    author = "Leslie Lamport",
    Publisher = "Addison--Wesley Publishing Company, Inc.",
    note = "Online information on {\TeX} and {\LaTeX} is available at "
     # tugURL # " and " # danteURL }

See the bibliography for how this will appear. No other modifications are required; L^{A}TEX and BibTEX should work as normal. Note that it would be sensible to put the @string commands into a separate file, urls.bib say, loaded with the main file via \bibliography{urls,...}.

The natbib package, written for L^{A}TEX by Patrick Daly, provides even more flexibility in the way a reference may be cited. All the features of this package are implemented for L^{A}TEX2HTML via the natbib.perl file. (Indeed there is even a mode whereby natbib handles the Harvard style of citation. This requires loading also the nharvard package.)

Thanks... to Martin Wilck for the bulk of the work in producing this extension, and to Ross Moore for necessary adjustments to allow it to work correctly with the document segmentation strategy.

\hypercite Analogous to \hyperref is the \hypercite command, which allows a freeform textual hyperlink to the bibliography, whereas the L^{A}TEX typeset version contains the usual citation code. The allowed syntax is as follows.

     \hypercite[int]{<HTML-text>}{<LaTeX-text>}{<opt-LaTeX>}{<label>}
     \hypercite[cite]{<HTML-text>}{<LaTeX-text>}{<opt-LaTeX>}{<label>}
     \hypercite{<HTML-text>}{<LaTeX-text>}{<opt-LaTeX>}{<label>}
     \hypercite[nocite]{<HTML-text>}{<LaTeX-text>}{<label>}
     \hypercite[no]{<HTML-text>}{<LaTeX-text>}{<label>}
     \hypercite[ext]{<HTML-text>}{<LaTeX-text>}{<label>}

The first three forms are equivalent; L^{A}TEX uses \cite[<opt-LaTeX>]<label> , after placing the <LaTeX-text>. Note that {<opt-LaTeX>} must be specified, even if empty ‘{}’.

Similarly the latter three forms are equivalent, with L^{A}TEX using \nocite{<label>} , to force the particular reference to appear on the bibliography page, even though no explicit marker is placed at this point. (Thus there is no need for an optional <opt-LaTeX> argument.) Within the HTML version a hyperlink is produced when the <HTML-text> is not empty. External label files are also searched, in order to match the symbolic <label>, see

also \externalcite on page 47.

Earlier in this manual the following source code was used:

    commands described in the \LaTeX{} \htmlcite{blue book}{lamp:latex},
    ...
    as well as many other \LaTeX{} constructions, such as are described in
    the \LaTeX{} \hypercite{\emph{Companion}}{\emph{Companion}}{}{goossens:latex}
    and \LaTeX{} \hypercite{\emph{Graphics Companion} (e.g. \Xy-pic)}%
    {\emph{Graphics Companion}}{\Xy-pic}{goossens:latexGraphics};

which produces:

commands described in the L^{A}TEX blue book ,

...

as well as many other L^{A}TEX constructions, such as are described in the L^{A}TEX Companion[2] and L^{A}TEX Graphics Companion[3, XY-pic];

whereas in the HTML version one sees:

commands described in the L^{A}TEX \underline{blue book},

...

as well as many other L^{A}TEX constructions, such as are described in the L^{A}TEX \underline{Companion} and L^{A}TEX \underline{Graphics Companion (e.g. XY-pic)};

\htmlcite Analogous to \htmlref is the \htmlcite command, which creates a textual hyperlink to a place on the document’s bibliography page, but without displaying any reference marker in the L^{A}TEX typeset version. (See above for an example.)

The \externalcite command, described on page 47, provides a similar facility when the bibliography page is “external”; that is, not part of the current document.

4.7 Symbolic References between Living Documents

The method of the previous section to generated symbolic hyperized links can easily be extended to external documents processed by L^{A}TEX2HTML. When L^{A}TEX2HTML processes a document, it generates a Perl file named <prefix>labels.pl which contains a list of all the symbolic labels that were defined, along with their locations. The <prefix> is empty unless otherwise specified, to allow different document segments to share the same directory.

\externallabels Links to an external document are then possible once a connection is established to that document’s labels.pl file. This connection is established by the

\externallabels command:
     \externallabels{<URL to directory of external document>}
                    {<local copy of external document labels.pl file>}

The first argument to \externallabels should be a URL to the directory containing the external document. The second argument should be the full path-name to the labels.pl file belonging to the external document. Note that for remote external documents it is necessary to copy the labels.pl file locally so that it can be read when processing a local document that uses it. The command \externallabels can be used once for each external document in order to import the external labels into the current document. A warning is given if labels.pl cannot be found.

If a symbolic reference made in either of the commands described in Section 4.5 is not defined within the document itself, L^{A}TEX2HTML will look for that reference in one of the external files^{7}. After any modifications in an external document (sections added/deleted, segmentation into different physical parts, etc.) a new labels.pl will be generated. If the \externallabels command in another document contains the correct address to an updated copy of the labels.pl file, then the cross-references will be re-aligned after running the local document through the translator.

There is also a mechanism analogous to the label–ref pairs of L^{A}TEX, which can be used only within a single document. These labels are called internal labels, as opposed to the external labels defined above. They are used extensively with the document segmentation strategy described in Section 4.10.

Either type of label is defined with a L^{A}TEX \label command. Labels can be referenced within a document using a \ref command. When processed by L^{A}TEX, each \ref command is replaced by the section number in which the corresponding \label occurred. When processed by the translator, each \ref is replaced by a hypertext link to the place where the corresponding \label occurred.

\externalref This mechanism can be extended to external documents:

     \externalref{<symbolic label in remote document>}

The argument to \externalref may be any symbolic label defined in the labels.pl file of any of the external documents. Such references to external symbolic labels are then translated into hyper-links pointing to the external document.

\externalcite Analogous to \externalref, the \externalcite command is used to

create a citation link, where the bibliography page is not part of the current document. As with \externalref symbolic labels for the bibliography page must have been loaded using

\externallabels.

A particularly important use for this is in allowing multiple documents to access information in a common bibliographic listing. For example: all of an author’s publications; a comprehensive listing of publications in a particular field; the (perhaps yearly) output of publications from a particular organisation or institution.

Thanks... to Uffe Engberg for suggesting this feature.

4.7.1 Cross-Referencing Example

To understand this mechanism better consider how you would maintain a link to this section (of the hypertext version of this document) from one of your documents, without using labels. Sure enough you can get the name of the physical file that this section is in. This however is quite likely to change, and any links to it would become invalid. To update your link, the name of the new file must be found and your link changed by hand. Also there is no general updating mechanism, so the only way to find out if your document is pointing to the right place is by actually following the link, then doing a manual update^{8}.

Next consider how it could be done with symbolic labels. First you have to import the labels used in this document by copying the file labels.pl, saving it in /tmp/labels.pl say, then adding anywhere in your document:

    \externallabels{http://cbl.leeds.ac.uk/nikos/tex2html/doc/manual}%
                   {/tmp/labels.pl}

7Care must be taken to ensure that critical symbolic references are unique across related documents.

8Link validation can be done automatically but the updating must be done manually when filenames have changed (assuming no other symbolic label mechanism is available).

After that you can use the label ‘crossrefs’ defined at the beginning of this section^{9} as follows:

    \externalref{crossrefs}

This will be translated into the appropriate hyper-link to this page. If there are any changes in this document and you would like to bring your document up-to date, you have to copy labels.pl again and rerun the translator on your document. Of course if I move the directory containing the HTML files for this document somewhere else, then you would have to make a change in the argument of the \externallabels command to reflect this.

It is obvious that some level of collaboration is required between authors trying to maintain cross-references between different documents. Using symbolic labels makes this a lot easier (especially for documents written by the same author).

4.8 Miscellaneous commands for HTML effects

The html package, through the L^{A}TEX input file html.sty, and its Perl counterpart html.perl, implements several new commands that are intended entirely for effects within the produced HTML files. In L^{A}TEX these commands, their arguments, and any optional arguments are completely ignored.

\htmlrule and \htmlrule* One such device provided by html.sty, is the \htmlrule

command. This puts a horizontal rule into the HTML file only; being ignored in the .dvi version. It is useful to provide extra visual separation between paragraphs, without creating a new HTML page, such as might warrant extra vertical space within the printed version.

Much variation can be obtained in the horizontal rule that is produced, using extended forms of the \htmlrule command:

     \htmlrule
     \htmlrule*
     \htmlrule[<attribs>]
     \htmlrule*[<attribs>]

Whereas a “break” tag <BR> normally precedes the <HR> generated by the \htmlrule command, this break is omitted when using the \htmlrule* variant. Furthermore, the optional argument <attribs> can be used to specify attributes for both the <HR> and <BR> tags. More specifically, <attribs> should be a list of attribute-names and/or key-value pairs <key>=<value> separated by spaces or commas. This list is parsed to extract those attributes applicable to the <HR> tag, and those applicable to the <BR> (with the unstarred variant).

Using HTML 3.2, this allows variations to be specified for:

• the (vertical) thickness of the horizontal line in pixels: SIZE=<num>;

• the (horizontal) width of the line in pixels or points: WIDTH=<width>;

   • alignment: WIDTH="..." taking left, right or center;

• removal of the shadowed effect NOSHADE;

• positioning of the rule with respect to text-flows: CLEAR="..." taking left, all,

     right or none.

Some examples of these effects appear on the HTML version of this page.

9You either have to guess the role of each label by looking at the labels.pl file or by asking the author!

\strikeout{<text>} With this command the <text> is processed as normal in the HTML version, then placed between <STRIKE>...</STRIKE> tags. Thus a horizontal line should be drawn through the middle of the <text>. Currently the command and the <text> are ignored in the L^{A}TEX version.

\tableofchildlinks As an extra aid to navigation within a long page, containing several (sub)subsections or deeper levels of sectioning, there is the \tableofchildlinks command. This does not generate anything new, for a table of the child links on or from a page is generated automatically by L^{A}TEX2HTML.

However if this command, or its variant \tableofchildlinks*, occurs within the source code to appear on a particular HTML page, then the child-links table will be placed at that point where the command occurs. Normally a break tag <BR> is inserted to separate the table of child-links from the surrounding text. The \tableofchildlinks* omits this extra break when it would result in too much space above the table.

For example throughout this section of the HTML version of the manual, all subsections in which several explicit commands have been discussed have their child-links table placed at the top of the page, using \tableofchildlinks*. This helps to quickly find the description of how the commands are used.

\htmlinfo Normally an “About this document...” page is created at the end of the HTML document, containing technical information about how the document was created, by whom, or any other information contained in the $INFO variable. This information can be made to appear at any other place within the document by specifying \htmlinfo at the desired place in the source. For example, the information may be best suited for the title-page.

The variant \htmlinfo* places the information, but leaves out the standard “About this document...” header. Instead the \htmlhead command can be used to place an alternative heading, prior to the \htmlinfo* command. Neither this heading nor the $INFO contents appears in the L^{A}TEX typeset version.

\bodytext{<options>} The text and background colors, and colors for the text of hypertext links can be set on an HTML page by giving appropriate attributes with the <BODY...> tag. This is particularly easy to do using the \bodytext command, which simply inserts the <code> as the desired list of attributes.

Warning: Any previous settings for the <BODY...> tag are discarded. Furthermore no checking is done to verify whether the given <options> indeed contains a list of attributes and values valid for the <BODY...> tag. When using \bodytext you are assumed to know precisely what you are doing!

Other packages contain commands which alter the contents of the <BODY...> tag; notably the color.perl implementation of L^{A}TEX’s color package, and the (prototype) frames package, by Martin Wilck and Ross Moore. In both these packages the requested information is checked for validity as an attribute within the <BODY...> tag.

\htmlbody{<options>} This is similar to the \bodytext command, except that it adds the value of an attribute, or allows an existing value to be changed. Thus it can be used to alter just a single one of the text and background colors, colors for the text of hypertext links or add a background pattern. The <options> are given as key-value pairs; some checking is done to ensure the validity of the attributes whose values are being set.

\htmlbase{<URL>} This specifies that the given <URL> be included in the <HEAD> section of each HTML page via a tag: <BASE HREF="<URL>". Such a feature is particularly useful…

• when preparing a document whose final location may be different from where it was created; By making all internal references be relative (to the the provided <URL>), a whole directory tree containing the document and all its subparts can be moved to elsewhere. A single edit in each HTML file produces the complete document intact at the new location.

• by allowing just single page to be copied to another location, but act as if it were part of the original document (provided this is accessible across the Web). Relative URLs within the copied page are relative to the base <URL>, rather than relative to the new location.

• Other uses for this feature are likely to become apparent.

\HTMLset{<which>}{<value>} and \HTMLsetenv{<which>}{<value>} The \HTMLset

command provides a mechanism whereby an arbitrary Perl variable can be assigned a value dynamically, during the L^{A}TEX2HTML processing. A variable having name ‘$<which>’ is assigned the specified <value>, overwriting any value that may exist already. The \HTMLsetenv is for the same purpose, but it is expanded in order as if it were an environment, rather than a command.

Warning: This is intended for Perl programmers only. Use this command at your own risk!

\latextohtml expands to the name L^{A}TEX2HTML, of this translator. Commands for parts of names of important L^{A}TEX packages are also included with L^{A}TEX2HTML: e.g. TEX, L^{A}TEX, AMS, XY. (This is to make it easy to refer to these products, in a consistent way within the HTML pages; you may still need L^{A}TEX definitions for the typeset version.)

4.9 Active Image Maps

Image maps are images with active regions in which a Web-surfer can click, to send him off to another sector of cyberspace. L^{A}TEX2HTML can design either inline “figures” or external ones (with or without a thumbnail version) to be image-maps. However HTML requires a URL of a HTML map-file, which associates the coordinates of each active region in the map with a destination URL. Usually this map file is kept on the server machine, however HTML 3.2 also allows it to reside on the client side for faster response. Both configurations are supported by L^{A}TEX2HTML through the \htmlimage options ‘map=’ and ‘usemap=’ respectively.

Keeping such a map file up to date manually can be tedious, especially with dynamic documents under revision. An experimental program makemap helps automate this process. This program (which is really a Perl script) takes one mandatory argument and an optional argument. The mandatory argument is the name of a user-map file, defined below. The optional argument is the name of the directory where the HTML map file(s) are to be placed.

The best way of describing how this works is by example. Suppose a document has two figures designated to become active image-maps. The first figure includes a statement like:

    \begin{figure}
    \htmlimage{map=/cgi-bin/imagemap/BlockDiagram.map,...}
    . . .
    \end{figure}

The second figure has a line like:

    \begin{figure}
    \htmlimage{map=/cgi-bin/imagemap/FlowChart.map,...}
    . . .
    \end{figure}

A typical user-map file, named report.map, might contain the following information^{10}:

    #
    #  Define the location(s) of the labels.pl file(s):
    #
    +report/ <URL>
    #
    #  Define map #1:
    #
    BlockDiagram.map:
    label1  rect    288,145 397,189
    label2  rect    307,225 377,252
    label2  default
    #
    #  Define map #2
    #
    FlowChart.map:
    label3  circle  150,100 200,100
    label4  default

In this file, comments are denoted by a #-sign in column 1. The line beginning with +report states that the symbolic labels are to be found in the labels.pl contained in the directory report/, and that its associated URL is as stated. Any number of external labels.pl files may be so specified. The block diagram image has two active regions. The first is a rectangle bounded by corners (288, 145) and (397, 189), while the second is a rectangle bounded by corners (307, 225) and (377, 252). These coordinates can be obtained with the aid of a program such as xv. If the user clicks in the first rectangle, it will cause a branch to the URL associated with symbolic label label1 defined in the labels.pl file found in directory report/. The single active region in the flow chart figure is a circle centred at (150, 100) and passing through point (200, 100). Clicking in this region will cause a branch to symbolic label label3. Labels label2 and label4 will be visited if the user clicks anywhere outside of the explicit regions. If any labels are not defined in any of the labels.pl files mentioned, they will be interpreted as URLs without translation.

The HTML image-maps are generated and placed in directory report/ by invoking the

command: makemap report.map report .

4.10 Document Segmentation^{11}

One of the greatest appeals of the World-Wide Web is its high connectivity through hyperlinks. As we have seen, the L^{A}TEX author can provide these links either manually or symbolically. Manual links are more tedious because a URL must be provided by the author for every link, and updated every time the target documents change. Symbolic links are more convenient, because the translator keeps track of the URLs. Earlier releases of L^{A}TEX2HTML required the entire document to be processed together if it was to be linked symbolically. However it was easy for large documents to overwhelm the memory capacities of moderatesized computers. Furthermore, processing time could become prohibitively high, if even a small change required the entire document to be reprocessed.

10This file is designed for an NCSA server. CERN servers use “rect” instead of “rectangle,” specify a radius instead of an outer point in the circle, and enclose point coordinates by parentheses.

11This feature is supported only for users of L^{A}TEX 2ε.

For these reasons, program segmentation was developed. This feature enables the author to subdivide his document into multiple segments. Each segment can be processed independently by L^{A}TEX2HTML. Hypertext links between segments can be made symbolically, with references shared through auxiliary files. If a single segment changes, only that segment needs to be reprocessed (unless a label is changed that another segment requires). Furthermore, the entire document can be processed without modification by L^{A}TEX to obtain the printed version. The top level segment that L^{A}TEX reads is called the parent segment. The others are called child segments. Document segmentation does require a little more work on the part of the author, who will now have to undertake some of the book-keeping formerly performed by L^{A}TEX2HTML. The following four L^{A}TEX extensions carry out segmentation:

 \segment{<file>}{<sec-type>}{<heading>} This command indicates the start of a

new program segment. The segment resides in <file>.tex, represents the start of a new L^{A}TEX sectional unit of type <sec-type> (e.g., \section, \chapter, etc.) and has a heading of <heading>. (A variation \segment* of this command, is provided for segments that are not to appear in the table of contents.) These commands perform the following operations in L^{A}TEX:

1. The specified sectioning command is executed.

2. L^{A}TEX will write its section and equation counters into an auxiliary file, named <file>.ptr. It will also write an \htmlhead command to this file. This information will tell L^{A}TEX2HTML how to initialise itself for the new document segment.

3. L^{A}TEX will then proceed to input and process the file <file>.tex.

The \segment and \segment* commands are ignored by L^{A}TEX2HTML.

\internal[<type>]{<prefix>} This command directs L^{A}TEX2HTML to load inter-segment information of type <type> from the file <prefix><type>.pl . Each program segment must be associated with a unique filename-prefix, specified either through a command-line option, or through the installation variable $AUTO PREFIX . The information <type> must be one of the following:

internals This is the default type, which need not be given. It specifies that the internal labels from the designated segment are to be input and made available to the current segment.

contents The table of contents information from designated segment are to be made available to the current segment.

sections Sectioning information is to be read in. Note that the segment containing the table of contents requires both contents and sections information from all other program segments.

figure Lists of figures from other segments are to be read.

table Lists of tables from other segments are to be read. index Index information from other segments is to be read.

images Allows images generated in other segments to be reused with the current segment.

Note: If extensive indexing is to be used, then it is advisable to keep each <prefix> quite short. This is because the hyper-links in the index have text strings constructed from this <prefix>, when using the makeidx package. Having long names with multiply-indexed items results in an extremely inelegant, cumbersome index. See Section 3.8 for more details.

\startdocument The \begin{document} and \end{document} statements are contained

in the parent segment only. It follows that the child segments cannot be processed separately by L^{A}TEX without modification. However they can be processed separately by L^{A}TEX2HTML, provided it is told where the end of the L^{A}TEX preamble is; this is the function of the \startdocument directive. It substitutes for \begin{document} in child segments, but is otherwise ignored by both L^{A}TEX and L^{A}TEX2HTML.

\htmlhead{<sec-type>}{<heading>} This command is generated automatically by a \segment command. It is not normally placed in the document at all; instead it facilitates information being passed from parent to child via the <file>.ptr file. It identifies to L^{A}TEX2HTML that the current segment is a L^{A}TEX sectional unit of type <sec-type>, with the specified heading. This command is ignored by L^{A}TEX. From version v97.1 , it is possible to use this command to insert extra section-headings, for use in the HTML version only.

\htmlnohead When placed at the top of the preamble of a document segment, the \htmlnohead command discards everything from the current page that has been placed already. Usually this will be just the section-head, from the \htmlhead command in the .ptr file. Numbering and color information is unaffected. This allows an alternative heading to be specified, or no heading at all in special circumstances; e.g. the page contains a single large table with a caption.

\segmentcolor{<model>}{<color>} This command is generated automatically by a \segment command. It is not normally placed in the document at all; instead it facilitates information being passed from parent to child via the <file>.ptr file. It specifies to L^{A}TEX2HTML that text in the document should have the color <color> .

\segmentpagecolor{<model>}{<color>} This command is generated automatically by a \segment command. It is not normally placed in the document at all; instead it facilitates information being passed from parent to child via the <file>.ptr file. It specifies to L^{A}TEX2HTML that the background of in the document should have the color

<color> .

The use of the segmenting commands is best illustrated by the example below. You might want to check your segmented document for consistency using the -unsegment command line option.

4.10.1 A Segmentation Example

The best way to illustrate document segmentation is through a simple example. Suppose that a document is to be segmented into one parent and two child segments. Let the parent segment be report.tex, and the the two child segments be sec1.tex and sec2.tex. The latter are translated with filename prefixes of s1 and s2, respectively. This example is included with recent distributions of L^{A}TEX2HTML, having more prolific comments than are shown here.

The text of report.tex is as follows:

\documentclass{article}          % Must use LaTeX 2e
\usepackage{html,makeidx,color}
\internal[figure]{s1}            % Include internal information
\internal[figure]{s2}            % from children
\internal[sections]{s1}
\internal[sections]{s2}
\internal[contents]{s1}
\internal[contents]{s2}
\internal[index]{s1}
\internal[index]{s2}
\begin{document}                 % The start of the document
\title{A Segmentation Example}
\date{\today}
\maketitle
\tableofcontents
\listoffigures
% Process the child segments:
\segment{sec1}{section}{Section 1 title}
\segment{sec2}{section}{Section 2 title}
\printindex
\end{document}

This file obtains the information necessary to build an index, a table of contents and a list of figures from the child segments. It then proceeds to typeset these.

The first child segment sec1.tex is as follows:

\begin{htmlonly}
\documentclass{article}
\usepackage{html,color,makeidx}
\input{sec1.ptr}
\end{htmlonly}
\internal{s2}
\startdocument
Here is some text.
\subsection{First subsection}
Here is subsection 1\label{first}.
\begin{figure}
\colorbox{red}{Some red text\index{Color text}}
\caption[List of figure caption]{Figure 1 caption}
\end{figure}
Reference\index{Reference} to \ref{second}.

The first thing this child segment does is establish the L^{A}TEX packages it requires, then loads the counter information that was written by the \segment command that invoked it. Since this segment contains a symbolic reference (second) to the second segment, it must load the internal labels from that segment.

The final segment sec2.tex is as follows:

    \begin{htmlonly}
    \documentclass{article}
    \usepackage{html,makeidx}
    \input{sec2.ptr}
    \end{htmlonly}
    \internal{s1}
    \startdocument
    Here is another section\label{second}.
    Plus another\index{Reference, another} reference\ref{first}.
    \begin{figure}
    \fbox{The figure}
    \caption{The caption}
    \end{figure}

This segment needs to load internal labels from the first one, because of the reference to ‘first’. These circular dependencies (two segments referencing each other) are either not allowed or handled incorrectly by the Unix utility make, without resorting to time stamps and some trickery. A time-stamp is a zero-length file whose only purpose is to record its creation time. Besides evaluating segment interdependence, another function of make is to provide inter-segment navigation information.

A sample Makefile is included in the distribution. This correctly generates the fully-linked document. The first time it is invoked, it runs:

   • latex on report.tex twice;
   • dvips to generate report.ps ;
   • latex2html on sec1.tex ;

• latex2html on sec2.tex . At this point sec2.html is completely linked, since the labels from the sec1 were available;

• latex2html on sec1.tex to pick up the labels from sec2 ;

   • latex2html on report.tex .

Proper operation of make depends on the fact that L^{A}TEX2HTML updates its own internal label file only if something in its current program segment causes the labels to change from the previous run. This ensures that L^{A}TEX2HTML is not run unnecessarily. It is also usual for the information page to be suppressed by specifying ‘ -info 0 ’ for all but the top-level document.

In the above example, all segments are built within the same sub-directory report/ of the directory containing the L^{A}TEX source files. This is achieved simply by using the option ‘ -dir report ’ with each. All the images and <prefix><type>.pl files are created and stored within this directory.

Sometimes it is desirable to build one or more segments within separate sub-directories. This is especially so when a segment has a large number of images, or if it is required to be part of more than one combined document. In this case the ‘ -dir <dir> ’ options can be different, or omitted entirely. For inter-segment referencing to work, a “relative path” must be included as part of the <prefix> with each \internal command; e.g.

   \internal[figure]{../sect1/s1}

5 User Manual

L^{A}TEX2HTML is a program for creating hyperlinked sets of HTML pages from a document marked-up using L^{A}TEX commands. Previous sections have discussed the results of specific L^{A}TEX commands. In this section we discuss instead the extensive range of command-line switches and options, and other aspects of Perl code, that affect the way the translation is performed.

To use L^{A}TEX2HTML to translate a file <file>.tex containing L^{A}TEX commands, type:

     latex2html <file>.tex

This will create a new directory called <file> which will contain the generated HTML files, some log files and possibly some images. To view the result use an HTML browser, such as NCSA Mosaic or Netscape Navigator, on the main HTML file which is <file>/<file>.html . The file will contain navigation links to the other parts of the generated document. The .tex suffix is optional and will be supplied by the program if it is omitted by the user. Other suffixes are acceptable also, such as .doc .

It is possible to customise the output from L^{A}TEX2HTML using a number of command-line options (see Section 5.2) with which you can specify:

• how to break up the document;

• where to put the generated files, and what are their names;

• the title for the whole document;

• the signature at the end of each page is;

• how many navigation panels to provide, what links to put in them;

• what other documents this one links to;

• extra information to include about the document;

• whether to retain the original L^{A}TEX section-numbering scheme;

• and many other things that affect how the information is obtained, processed or displayed in the resulting .html files and images.

The L^{A}TEX2HTML script includes a short manual which can be viewed with the command:

     perldoc latex2html

5.1 Developing Documents using L^{A}TEX2HTML

Although any document containing L^{A}TEX commands can be translated by the L^{A}TEX2HTML translator, the best results are obtained when that document is itself a valid L^{A}TEX document. Indeed it is generally a good idea to develop documents so that they produce good readable results in both the L^{A}TEX typeset version as well as a set of HTML pages. This is not just a nicety; there are several good practical reasons for doing this.

L^{A}TEX macros: The macro commands that L^{A}TEX2HTML recognises are based upon corresponding commands for L^{A}TEX. If one tries to use syntax that is incorrect for L^{A}TEX then there is no reason why L^{A}TEX2HTML should be able to “get it right”, by somehow recognising the true intent.

error checking: Processing the document first using L^{A}TEX is the easiest, and quickest, way to check for valid syntax. Whereas L^{A}TEX stops at each error (when run in interactive mode), allowing a fix to be made “on the spot” or a “stop-fix-restart”, L^{A}TEX2HTML does not stop when it detects an error in L^{A}TEX syntax. Useful messages are given concerning missing or unmatched braces, but other apparent anomalies generate only warning messages, which are saved to the end. (Some warnings are also shown immediately when the $VERBOSITY variable is set to at least 3.) In practice it can be much quicker to test for invalid syntax using L^{A}TEX before attempting to use the L^{A}TEX2HTML translator.

Furthermore, L^{A}TEX warns of cross-reference labels that have not been defined. This is useful to help avoid having hyperlinks which point to nowhere.

The case of missing braces, or an unmatched opening brace, is an error that L^{A}TEX2HTML actually handles better than L^{A}TEX (or rather, the underlying TEX processor). Whereas TEX only detects an error when something else goes wrong later in the processing, L^{A}TEX2HTML shows where the unmatched brace itself occurs.

auxiliary file: Some information that L^{A}TEX2HTML might need is normally read from the .aux file for the document being processed; or perhaps from the .aux file of a different document, of which the current document is just a portion. Clearly a valid L^{A}TEX run is required to produce the .aux file.

Of course if no information in the .aux file is actually used, then this L^{A}TEX run can be neglected. Also, if the .aux file has already been created, and edits are made on the source which do not alter the information stored within the .aux file, then there is no need for a fresh L^{A}TEX run (except for the purposes of error-checking).

bibliography: Suppose the document requires a bibliography, or list of references, which is to be prepared using BibTEX. L^{A}TEX2HTML reads citation information from the .aux file, and can import the bibliography itself from the .bbl file. However these must first be created using L^{A}TEX.

document segmentation: With the document segmentation technique, discussed fully in Section 4.10, it is vitally important that the full document processes correctly in L^{A}TEX. The desired effect is that of a single large document, whereas the pieces will actually be processed separately. To achieve this, L^{A}TEX writes vital information into special .ptr files. Like the .aux file, these files are read by L^{A}TEX2HTML to get sectionnumbering and other information to be used while processing each segment.

print quality: When a document contains automatically-generate images, these images are usually bitmaps designed for viewing on-screen. In general the resolution of these is too poor to give a good result when printed on a high-resolution laser-printer. Thus if it is likely that the reader will want to obtain a printed version of your document, then it is best to include a hyperlink to the typeset .dvi version, or to a PostScript conversion of the .dvi file. (In either case, a link to a compressed version is even better.)

5.2 Command-Line Options

The command-line options described here can be used to change the default behaviour of L^{A}TEX2HTML. Alternatively, often there is a corresponding environment variable whose value may be set or changed within a .latex2html-init initialisation file, in order to achieve the same result. There are so many options that they are listed here in groups, according to the nature of the effects they control. When a large number of such options are required for the processing of a document, it is usual to store the switches and their desired settings within a Makefile, for use with the Unix make utility. Now a simple command such as:

 make mydocument

can initiate a call to latex2html that would otherwise take many lines of typing. Indeed it could instigate several such calls to L^{A}TEX2HTML, or to other programs such as L^{A}TEX and BibTEX, dvips and others. The document segmentation feature, discussed in Section 4.10, uses this technique extensively.

5.2.1 Options controlling Titles, File-Names and Sectioning

-t <top-page-title> (Same as setting: $TITLE = "<top-page-title>";)

Name the document using this title.

-short extn (Same as setting: $SHORTEXTN = 1;)

Use a filename prefix of .htm for the produced HTML files. This is particularly useful for creating pages to be stored on CD-ROM or other media, to be used with operating systems that require a 3-character extension.

-long titles <num> (Same as setting: $LONG TITLES = <num>;)

Instead of the standard names: node1.html, node2.html,... the filenames for each HTML page are constructed from the first <num> words of the section heading for that page, separated by the ‘ ’ character. Commas and common short words (a an to by of and for the) are omitted from both title and word-count.

Warning: Use this switch with great caution. Currently there are no checks for uniqueness of names or overall length. Very long names can easily result from using this feature.

-custom titles (Same as setting: $CUSTOM TITLES = 1;)

Instead of the standard names: node1.html, node2.html, ... the filenames for each HTML page are constructed using a Perl subroutine named custom title hook . The user may define his/her own version of this subroutine, within a .latex2html-init file say, to override the default (which uses the standard names). This subroutine takes the section-heading as a parameter and must return the required name, or the empty string (default).

-dir <output-directory> (Same as setting: $DESTDIR = "<output-directory>";)

Redirect the output to the specified directory. The default behaviour is to create (or reuse) a directory having the same name as the prefix of the document being processed.

-no subdir (Same as setting: $NO SUBDIR = 1;)

Place the generated HTML files into the current directory. This overrides any $DESTDIR setting.

-prefix <filename-prefix> (Same as setting: $PREFIX = "<filename-prefix>";)

The <filename-prefix> will be prepended to all .gif, .pl and .html files produced, except for the top-level .html file; it may include a (relative) directory path. This will enable multiple products of L^{A}TEX2HTML to peacefully coexist in the same directory. However, do not attempt to simultaneously run multiple instances of L^{A}TEX2HTML using the same output directory, else various temporary files will overwrite each other.

-auto prefix (Same as setting: $AUTO PREFIX = 1;)

Constructs the prefix as ‘<title>–’ to be prepended to all the files produced, where <title> is the name of the L^{A}TEX file being processed. (Note the ‘–’ in this prefix.) This overrides any $PREFIX setting.

-no auto link (Same as setting: $NO AUTO LINK = 1;)

If $NO AUTO LINK is empty and variables $LINKPOINT and $LINKNAME are defined appropriately (as is the default in the latex2html.config file), then a hard link to the main HTML page is produced, using the name supplied in $LINKNAME. Typically this is index.html; on many systems a file of this name will be used, if it exists, when a browser tries to view a URL which points to a directory. On other systems a different value for $LINKNAME may be appropriate. Typically $LINKPOINT has value $FILE.html, but this may also be changed to match whichever HTML page is to become the target of the automatic link. Use of the ‘ -no auto link ’ switch cancels this automatic linking facility, when not required for a particular document.

-split <num> (Same as setting: $MAX SPLIT DEPTH = <num>;) (default is 8) Stop splitting sections into separate files at this depth. Specifying ‘ -split 0 ’ will put the entire document into a single HTML file. See below for the different levels of sectioning. Also see the next item for how to set a “relative” depth for splitting.

-split +<num> (Same as setting: $MAX SPLIT DEPTH = -<num>;) (default is 8) The level at which to stop splitting sections is calculated “relative to” the shallowest level of sectioning that occurs within the document. For example, if the document contains \section commands, but no \part or \chapter commands, then ‘ -split +1 ’ will cause splitting at each \section but not at any deeper level; whereas ‘ -split

+2 ’ or ‘ -split +3 ’ also split down to \subsection and \subsubsection commands

respectively. Specifying ‘ -split +0 ’ puts the entire document into a single HTML file.

-link <num> (Same as setting: $MAX LINK DEPTH = <num>;) (default is 4) For each node, create links to child nodes down to this much deeper than the node’s sectioning-level. Specifying ‘ -link 0 ’ will show no links to child nodes from that page, ‘ -link 1 ’ will show only the immediate descendents, etc. A value at least as big as that of the ‘ -split <num> ’ depth will produce a mini table-of-contents (when not empty) on each page, for the tree structure rooted at that node.

When the page has a sectioning-level less than the ‘ -split ’ depth, so that the a mini table-of-contents has links to other HTML pages, this table is located at the bottom of the page, unless placed elsewhere using the \tableofchildlinks command.

On pages having a sectioning-level just less than the ‘ -split ’ depth the mini tableof-contents contains links to subsections etc. occurring on the same HTML page. Now the table is located at the top of this page, unless placed elsewhere using the

\tableofchildlinks command.

-toc depth <num> (Same as setting: $TOC DEPTH = <num>;) (default is 4) Sectioning levels down to <num> are to be included within the Table-of-Contents tree.

-toc stars (Same as setting: $TOC STARS = 1;)

Sections created using the starred-form of sectioning commands are included within the Table-of-Contents. As with L^{A}TEX, normally such sections are not listed.

-show section numbers (Same as setting: $SHOW SECTION NUMBERS = 1;) Show section numbers. By default section numbers are not shown, so as to encourage the use of particular sections as stand-alone documents.

-unsegment (Same as setting: $UNSEGMENT = 1;)

Treat a segmented document (see the section about document segmentation) like it were not segmented. This will cause the translator to concatenate all segments and process them as a whole. You might find this useful to check a segmented document for consistency.

For all documents the sectioning levels referred to above are:

0 document
1 part
2 chapter
3 section
4 subsection
5 subsubsection
6 paragraph
7 subparagraph
8 subsubparagraph

These levels apply even when the document contains no sectioning for the shallower levels; e.g. no \part or \chapter commands is most common, especially when using L^{A}TEX’s article document-class.

5.2.2 Options controlling Extensions and Special Features

The switches described here govern the type of HTML code that can be generated, and how to choose between the available options when there are alternative strategies for implementing portions of L^{A}TEX code.

-html version (2.0|3.2|4.0|5.0)[,(math|i18n)]*
(Same as setting: $HTML VERSION = "...  ";)

This specifies both the HTML version to generate, and any extra (non-standard) HTML features that may be required. The version number corresponds to a published DTD for an HTML standard. A corresponding Perl file in the versions/ subdirectory is loaded; these files are named ‘html<num>.pl’.

Following the version number, a comma-separated list of extensions can be given. Each corresponds to a file ‘<name>.pl’ also located in the versions/ subdirectory. When such a file is loaded the resulting HTML code can no longer be expected to validate with the specified DTD. An exception is ‘ math ’ when the ‘ -no math ’ switch is also used, which should still validate.

Currently, versions 2.0, 3.2. 4.0 and 5.0 are available. The default version is usually set to be ‘5.0’, within latex2html.config .

-no tex defs (Same as setting: $TEXDEFS = 0;) (default is 1)

When $TEXDEFS is set (default) the file texdefs.perl will be read. This provides code to allow common TEX commands like \def, \newbox, \newdimen and others, to be recognised, especially within the document preamble. In the case of \def, the definition may even be fully interpreted, but this requires the pattern-matching to be not too complicated.

If $TEXDEFS is ‘0’ or empty, then texdefs.perl will not be loaded; the translator will make no attempt to interpret any raw TEX commands. This feature is intended to enable sophisticated authors the ability to insert arbitrary TEX commands in environments that are destined to be processed by L^{A}TEX anyway; e.g. figures, , pictures, etc. However this should rarely be needed, as now there is better support for these types of environment. There are now other methods to specify which chunks of code are to be passed to L^{A}TEX for explicit image-generation; see the discussion of the makeimage environment on page 43.

-external file <filename> (Same as setting: $EXTERNAL FILE = "<filename>";)

Specifies the prefix of the .aux file that this document should read. The .aux extension will be appended to this prefix to get the complete filename, with directory path if needed. This file could contain necessary information regarding citations, figure, table and section numbers from L^{A}TEX and perhaps other information also. Use of this switch is vital for document segments, processed separately and linked to appear as if generated from a single L^{A}TEX document.

-font size <size> (Same as setting: $FONT SIZE = "<size>";)

This option provides better control over the font size of environments made into images using L^{A}TEX. <size> must be one of the font sizes that L^{A}TEX recognizes; i.e. ‘10pt’, ‘11pt’, ‘12pt’, etc. Default is ‘10pt’, or whatever option may have been specified on the \documentclass or \documentstyle line. Whatever size is selected, it will be magnified by the installation variables $MATH SCALE FACTOR, $FIGURE SCALE FACTOR and $DISP SCALE FACTOR as appropriate.

Note: This switch provides no control over the size of text on the HTML pages. Such control is subject entirely to the user’s choices of settings for the browser windows.

-scalable fonts (Same as setting: $SCALABLE FONTS = 1;)

This is used when scalable fonts, such as PostScript versions of the TEX fonts, are available for image-generation. It has the effect of setting $PK GENERATION to ‘1’, and $DVIPS MODE to be empty, overriding any previous settings for these variables.

-no math (Same as setting: $NO SIMPLE MATH = 1;)

Ordinarily simple mathematical expressions are set using the ordinary text font, but italiced. When part of the expression can not be represented this way, an image is made of the whole formula. This is called “simple math”. When $NO SIMPLE MATH is set, then all mathematics is made into images, whether simple or not.

However, if the ‘ math ’ extension is loaded, using the ‘ -html version ’ switch described earlier, then specifying ‘ -no math ’ produces a quite different effect. Now it is the special <MATH> tags and entities which are cancelled. In their place a sophisticated scheme for parsing mathematical expressions is used. Images are made of those sub-parts of a formula which cannot be adequately expressed using (italiced) text characters and <SUB> and <SUP> tags. See Section 3.3 for more details.

-local icons (Same as setting: $LOCAL ICONS = 1;) A copy of each of the icons actually used within the document is placed in the directory along with the HTML files and generated images. This allows the whole document to be fully self-contained, within this directory; otherwise the icons must be retrieved from a (perhaps remote) server. It is also the default behavior if $ICONSERVER is not set. The icons are normally copied from a subdirectory of the $LATEX2HTMLDIR, set within latex2html.config. An alternative set of icons can be used by specifying a (relative) directory path in $ALTERNATIVE ICONS to where the customised images can be found.

-init file <file> Load the specified initialisation file. This Perl file will be loaded after

loading $HOME/.latex2html-init , or .latex2html-init in the local directory, if

either file exists. It is read at the time the switch is processed, so the contents of the file may change any of the values of any of the variables that were previously established, as well as any default options. More than one initialisation file can be read in this way.

-no fork (Same as setting: $NOFORK = 1;)

When set this disables a feature in the early part of the processing whereby some memory-intensive operations are performed by ‘forked’ child processes. Some singletask operating systems, such as DOS, do not support this feature. Having $NOFORK set then ensures that unnecessary file-handles that are needed with the forked processes, are not consumed unnecessarily, perhaps resulting in a fatal Perl error.

-iso language <type> This enables you to specify a different language type than 'EN' to be used in the DTD entries of the HTML document, e.g. 'EN.US'.

-short index (Same as setting: $SHORT INDEX = 1;)

Creates shorter Index listings, using codified links; this is fully compatible with the makeidx package.

-no footnode (Same as setting: $NO FOOTNODE = 1;)

Suppresses use of a separate file for footnotes; instead these are placed at the bottom of the HTML pages where the references occur. When this option is used, it is frequently desirable to change the style of the marker used to indicate the presence of a footnote. This is done as in L^{A}TEX, using code such as follows.

    \renewcommand{\thefootnote}{\arabic{footnote}}

All the styles \arabic, \alph, \roman, \Alph and \Roman are available.

-numbered footnotes (Same as setting: $NUMBERED FOOTNOTES = 1;)

If this is set you will get every footnote applied with a subsequent number, to ease readability.

-address <author-address> (Same as setting: $ADDRESS = "<author-address>";)

Sign each page with this address. See latex2html.config for an example using Perl code to automatically include the date. A user-defined Perl subroutine called &custom address can be used instead, if defined; it takes the value of $ADDRESS as a parameter, which may be used or ignored as desired. At the time when this subroutine will be called, variables named $depth, $title, $file hold the sectioning-level, title and filename of the HTML page being produced; $FILE holds the name of the filename for the title-page of the whole document.

-info <string> (Same as setting: $INFO = "<string>";)

Generate a new section “About this document” containing information about the document being translated. The default is to generate such a section with information on the original document, the date, the user and the translator. An empty string (or the value ‘0’) disables the creation of this extra section. If a non-empty string is given, it will be placed as the contents of the “About this document” page instead of the default information.

5.2.3 Switches controlling Image Generation

These switches affect whether images are created at all, whether old images are reused on subsequent runs or new ones created afresh, and whether anti-aliasing effects are used within the images themselves.

-image type Specify the type of images to be generated. Depending on your setup, LaTeX2HTML can generate svg, png or gif images. Vector formats such as svg look better at high resolution, while bitmap formats such as png or gif are generally faster to download and to render.

-use dvipng Use the dvipng program to generate png images, rather than dvips followed by gs. This method produces better alignment of math formulas which extend significantly above or below the line of text in which they are contained. An example of this behavior can be seen in the file tests/eq line spacing.tex. The dvipng method also eliminates the ugly crop marks that appear with 12pt documents. It does not respect the $MATH SCALE FACTOR option.

-(no)use pdftex By default, pdflatex is used to process input files. Specify -nouse pdftex for documents that rely on standard, dvi-producing latex.

The pdflatex method uses the pdflatex program followed by pdfcrop and gs to generate images, rather than latex followed by dvips. This method can be useful for pdfLaTeX documents which cannot be translated by latex. The pdflatex method generally produces better alignment of math formulas and eliminates ugly crop marks. It does not respect the $MATH SCALE FACTOR option.

The pdflatex method uses the pdfwrite GhostScript driver by default. If called together with the -use dvipng option, it will use the png16m driver and produce slightly different math alignment.

-use luatex Use the lualatex program followed by pdfcrop and gs to generate images, rather than latex followed by dvips. This method can be useful for LuaLaTeX documents which cannot be translated by latex or pdflatex. This method can sometimes produce slightly better alignment of math formulas and eliminate ugly crop marks. It does not respect the $MATH SCALE FACTOR option.

This method uses the pdfwrite GhostScript driver by default. If called together with the -use dvipng option, it will use the png16m driver and produce slightly different math alignment.

An example of using the -use luatex option can be seen on the file

tests/polyglossia.tex.

-use luadvi Use the dvilualatex program instead of latex to generate images. This method can be useful for dvilualatex documents which cannot be translated by latex. It can be combined with the -use dvipng option as usually.

-ascii mode (Same as setting: $ASCII MODE = $EXTERNAL IMAGES = 1;) Use only ASCII characters and do not include any images in the final output. With ‘ -ascii mode ’ the output of the translator can be used on character-based browsers, such as lynx, which do not support inlined images (via the <IMG> tag).

-nolatex (Same as setting: $NOLATEX = 1;)

Disable the mechanism for passing unknown environments to L^{A}TEX for processing. This can be thought of as “draft mode” which allows faster translation of the basic document structure and text, without fancy figures, equations or tables.

(This option has been superseded by the ‘ -no images ’ option, see below.)

-external images (Same as setting: $EXTERNAL IMAGES = 1;) Instead of including any generated images inside the document, leave them outside the document and provide hypertext links to them.

-ps images (Same as setting: $PS IMAGES = $EXTERNAL IMAGES = 1;) Use links to external PostScript files rather than inlined images in the chosen graphics format.

-discard (Same as setting: $DISCARD PS = 1;)

The temporary PostScript files are discarded immediately after they have been used to create the image in the desired graphics format.

-no images (Same as setting: $NO IMAGES = 1;)

Do not attempt to produce any inlined images. The missing images can be generated “off-line” by restarting L^{A}TEX2HTML with the option ‘ -images only ’.

-images only (Same as setting: $IMAGES ONLY = 1;)

Try to convert any inlined images that were left over from previous runs of L^{A}TEX2HTML.

-reuse <reuse option> (Same as setting: $REUSE = <reuse option>;)

This switch specifies the extent to which image files are to be shared or recycled. There are three valid options:

0 Do not ever share or recycle image files.

This choice also invokes an interactive session prompting the user about what to do about a pre-existing HTML directory, if it exists.

1 Recycle image files from a previous run if they are available, but do not share identical images that must be created in this run.

2 Recycle image files from a previous run and share identical images from this run. This is the default.

Section 3.4.2 provides additional information about image-reuse.

-no reuse (Same as setting: $REUSE = 0;)

Do not share or recycle images generated during previous translations. This is equivalent to ‘ -reuse 0 ’. (This will enable the initial interactive session during which the user is asked whether to reuse the old directory, delete its contents or quit.)

-antialias (Same as setting: $ANTI ALIAS = 1;) (Default is 0.) Generated images of figure environments and external PostScript files should use antialiasing. By default anti-aliasing is not used with these images, since this may interfere with the contents of the images themselves.

-antialias text (Same as setting: $ANTI ALIAS TEXT = 1;) (Default is 1.) Generated images of typeset material such as text, mathematical formulas, tables and the content of makeimage environments, should use anti-aliasing effects. The default is normally to use anti-aliasing for text, since the resulting images are much clearer on-screen. However the default may have been changed locally.

-no antialias (Same as setting: $ANTI ALIAS = 0;) (Default is 0.) Generated images of figure environments and external PostScript files should not use anti-aliasing with images, though the local default may have been changed to use it.

-no antialias text (Same as setting: $ANTI ALIAS TEXT = 0;) (Default is 1.) Generated images of typeset material should not use anti-aliasing effects. Although on-screen images of text are definitely improved using anti-aliasing, printed images can be badly blurred, even at 300dpi. Higher resolution printers do a much better job with the resulting grey-scale images.

-white (Same as setting: $WHITE BACKGROUND = 1;) (Default is 1.) Ensures that images of figure environments have a white background. Otherwise transparency effects may not work correctly.

-no white (Same as setting: $WHITE BACKGROUND = '';) (Default is 1.) Cancels the requirement that figure environments have a white background.

-ldump (Same as setting: $LATEX DUMP = 1;) (Default is 0.)

Use this if you want to speed up image processing during the 2nd and subsequent runs of L^{A}TEX2HTML on the same document. The translator now produces a L^{A}TEX format-dump of the preamble to images.tex which is used on subsequent runs. This significantly reduces the startup time when L^{A}TEX reads the images.tex file for imagegeneration.

This process actually consumes additional time on the first run, since L^{A}TEX is called twice — once to create the format-dump, then again to load and use it. The pay-off comes with the faster loading on subsequent runs. Approximately 1 Meg of disk space is consumed by the dump file.

5.2.4 Switches controlling Navigation Panels

The following switches govern whether to include one or more navigation panels on each HTML page, also which buttons to include within such a panel.

-no navigation (Same as setting: $NO NAVIGATION = 1;)

Disable the mechanism for putting navigation links in each page. This overrides any settings of the $TOP NAVIGATION, $BOTTOM NAVIGATION and

$AUTO NAVIGATION variables.

-top navigation (Same as setting: $TOP NAVIGATION = 1;)

Put navigation links at the top of each page.

-bottom navigation (Same as setting: $BOTTOM NAVIGATION = 1;) Put navigation links at the bottom of each page as well as the top.

-auto navigation (Same as setting: $AUTO NAVIGATION = 1;)

Put navigation links at the top of each page. Also put one at the bottom of the page, if the page exceeds $WORDS IN PAGE number of words (default = 450).

-next page in navigation (Same as setting: $NEXT PAGE IN NAVIGATION = 1;) Put a link to the next logical page in the navigation panel.

-previous page in navigation

(Same as setting: $PREVIOUS PAGE IN NAVIGATION = 1;)

Put a link to the previous logical page in the navigation panel.

-contents in navigation (Same as setting: $CONTENTS IN NAVIGATION = 1;) Put a link to the table-of-contents in the navigation panel if there is one.

-index in navigation (Same as setting: $INDEX IN NAVIGATION = 1;) Put a link to the index-page in the navigation panel if there is an index.

5.2.5 Switches for Linking to other documents

When processing a single stand-alone document, the switches described in this section should not be needed at all, since the automatically generated navigation panels, described in Section 5.2.4, should generate all the required navigation links. However if a document is to be regarded as part of a much larger document, then links from its first and final pages, to locations in other parts of the larger (virtual) document, need to be provided explicitly for some of the buttons in the navigation panel.

The following switches allow for such links to other documents, by providing the title and URL for navigation panel hyperlinks. In particular, the “Document Segmentation” feature of Section 4.10 necessarily makes great use of these switches. It is usual for the text and targets of these navigation hyperlinks to be recorded in a Makefile, to avoid tedious typing of long command-lines having many switches.

-up url <URL> (Same as setting: $EXTERNAL UP LINK = "<URL>";)

Specifies a universal resource locator (URL) to associate with the “UP” button in the navigation panel(s).

-up title <string> (Same as setting: $EXTERNAL UP TITLE = "<string>";)

Specifies a title associated with this URL.

-prev url <URL> (Same as setting: $EXTERNAL PREV LINK = "<URL>";)

Specifies a URL to associate with the “PREVIOUS” button in the navigation panel(s).

-prev title <string> (Same as setting: $EXTERNAL PREV TITLE = "<string>";)

Specifies a title associated with this URL.

-down url <URL> (Same as setting: $EXTERNAL DOWN LINK = "<URL>";)

Specifies a URL for the “NEXT” button in the navigation panel(s).

-down title <string> (Same as setting: $EXTERNAL DOWN TITLE = "<string>";)

Specifies a title associated with this URL.

-contents <URL> (Same as setting: $EXTERNAL CONTENTS = "<URL>";)

Specifies a URL for the “CONTENTS” button, for document segments that would not otherwise have one.

-index <URL> (Same as setting: $EXTERNAL INDEX = "<URL>";)

Specifies a URL for the “INDEX” button, for document segments that otherwise would not have an index.

-biblio <URL> (Same as setting: $EXTERNAL BIBLIO = "<URL>";)

Specifies the URL for the bibliography page to be used, when not explicitly part of the document itself.

Warning: On some systems it is difficult to give text-strings <string> containing space characters, on the command-line or via a Makefile. One way to overcome this is to use the corresponding variable. Another way is to replace the spaces with underscores (_).

5.2.6 Switches for Help and Tracing

The first two of the following switches are self-explanatory. When problems arise in processing a document, the switches ‘ -debug ’ and ‘ -verbosity ’ will each cause L^{A}TEX2HTML to generate more output to the screen. These extra messages should help to locate the cause of the problem.

-tmp <path> Define a temporary directory to use for image generation. If <path> is 0, the standard temporary directory /tmp is used.

-h(elp) Print out the list of all command-line options.

-v Print the current version of L^{A}TEX2HTML.

-debug (Same as setting: $DEBUG = 1;)

Run in debug-mode, displaying messages and/or diagnostic information about files read, and utilities called by L^{A}TEX2HTML. Shows any messages produced by these calls.

More extensive diagnostics, from the Perl debugger, can be obtained by appending the string ‘-w-’ to the 1st line of the latex2html (and other) Perl script(s).

-verbosity <num> (Same as setting: $VERBOSITY = <num>;)

Display messages revealing certain aspects of the processing performed by L^{A}TEX2HTML on the provided input file(s). The <num> parameter can be an integer in the range 0 to 8. Each higher value adds to the messages produced.

0. No special tracing; as for versions of L^{A}TEX2HTML prior to v97.1 .

1. (This is the default.) Show section-headings and the corresponding HTML file names, and indicators that major stages in the processing have been completed.

2. Print environment names and identifier numbers, and new theorem-types. Show warnings as they occur, and indicators for more stages of processing. Print names of files for storing auxiliary data arrays.

3. Print command names as they are encountered and processed; also any unknown commands encountered while pre-processing. Show names of new commands, environments, theorems, counters and counter-dependencies, for each document partition.

4. Indicate command-substitution the pre-process of math-environments. Print the contents of unknown environments for processing in L^{A}TEX, both before and after reverting to L^{A}TEX source. Show all operations affecting the values of counters. Also show links, labels and sectioning keys, at the stages of processing.

5. Detail the processing in the document preamble. Show substitutions of new environments. Show the contents of all recognised environments, both before and after processing. Show the cached/encoded information for the image keys, allowing two images to be tested for equality.

6. Show replacements of new commands, accents and wrapped commands.

7. Trace the processing of commands in math mode; both before and after.

8. Trace the processing of all commands, both before and after.

The command-line option sets an initial value only. During processing the value of $VERBOSITY can be set dynamically using the \htmltracing{...} command, whose argument is the desired value, or by using the more general \HTMLset command as

     follows: \HTMLset{VERBOSITY}{<num>} .

5.2.7 Other Configuration Variables, without switches

The configuration variables described here do not warrant having a command-line switch to assign values. Either they represent aspects of L^{A}TEX2HTML that are specific to the local site, or they govern properties that should apply to all documents, rather than something that typically would change for the different documents within a particular sub-directory.

Normally these variables have their value set within the latex2html.config file. In the following listing the defaults are shown, as the lines of Perl code used to establish these values. If a different value is required, then these can be assigned from a local .latex2html-init initialisation file, without affecting the defaults for other users, or documents processed from other directories.

$dd holds the string to be used in file-names to delimit directories; it is set internally to ‘/’, unless the variable has already been given a value within latex2html.config .

Note: This value cannot be set within a .latex2html-init initialisation file, since its value needs to be known in order to find such a file.

$LATEX2HTMLDIR Read by the install-test script from latex2html.config, its value is

inserted into the latex2html Perl script as part of the installation process.

$LATEX2HTMLSTYLES = "$LATEX2HTMLDIR/styles"; Read from the latex2html.config

file by install-test , its value is checked to locate the styles/ directory.

$LATEX2HTMLVERSIONS = "$LATEX2HTMLDIR/versions"; The value of this variable should

be set within latex2html.config to specify the directory path where the version and extension files can be found.

$ALTERNATIVE ICONS = ''; This may contain the (relative) directory path to a set of customised icons to be used in conjunction with the ‘ -local icons ’ switch.

$TEXEXPAND = "$LATEX2HTMLDIR/texexpand"; Read by the install-test Perl script

from latex2html.config, its value is used to locate the texexpand Perl script.

$PSTOIMG = "$LATEX2HTMLDIR/pstoimg"; Read by the install-test Perl script from

latex2html.config, its value is used to locate the pstoimg Perl script.

$IMAGE TYPE = '<image-type>'; Set in latex2html.config, the currently supported
     <image-type>s are: svg, png and gif.
$DVIPS = 'dvips'; Read from latex2html.config by install-test, its value is checked

to locate the dvips program or script.

There could be several reasons to change the value here:

• add a switch -P<printer> to load a specific configuration-file; e.g. to use a specific set of PostScript fonts, for improved image-generation.

• to prepend a path to a different version of dvips than normally available as the system default (e.g. the printing requirements are different).

• to append debugging switches, in case of poor quality images; one can see which paths are being searched for fonts and other resources.

• to prepend commands for setting path variables that dvips may need in order to locate fonts or other resources.

If automatic generation of fonts is required, using Metafont, the following configuration variables are important.

$PK GENERATION = 1; This variable must be set, to initiate font-generation; otherwise fonts will be scaled from existing resources on the local system. In particular this variable must not be set, if one wishes to use PostScript fonts or other scalable font resources (see the ‘ -scalable fonts ’ switch).

$DVIPS MODE = 'toshiba'; The mode given here must be available in the modes.mf file, located with the Metafont resource files, perhaps in the misc/ subdirectory.

$METAFONT DPI = 180; The required resolution, in dots-per-inch, should be listed specifically within the MakeTeXPK script, called by dvips to invoke Metafont with the correct parameters for the required fonts.

$LATEX = 'latex'; Read from latex2html.config by install-test, its value is checked

to locate the latex program or script.

If L^{A}TEX is having trouble finding style-files and/or packages, then the default command can be prepended with other commands to set environment variables intended to resolve these difficulties;

e.g. $LATEX = 'setenv TEXINPUTS <path to search> ; latex' .

There are several variables to help control exactly which files are read by L^{A}TEX2HTML and by L^{A}TEX when processing images:

$TEXINPUTS This is normally set from the environment variable of the same name. If difficulties occur so that styles and packages are not being found, then extra paths can be specified here, to resolve these difficulties.

$DONT INCLUDE This provides a list of filenames and extensions to not include, even if requested to do so by an \input or \include command. (Consult latex2html.config for the default list.)

$DO INCLUDE = ''; List of exceptions within the $DONT INCLUDE list. These files are to be read if requested by an \input or \include command.

$ICONSERVER = '<URL>'; This is used to specify a URL to find the standard icons, as used for the navigation buttons. Names for the specific images size, as well as size information, can be found in latex2html.config. The icons themselves can be replaced by customised versions, provided this information is correctly updated and the location of the customised images specified as the value of $ICONSERVER. When the ‘ -local icons ’ switch is used, so that a copy of the icons is placed with the HTML files and other generated images, the value of $ICONSERVER is not needed within the HTML files themselves.

$NAV BORDER = <num>; The value given here results in a border, measured in points, around each icon. A value of ‘0’ is common, to maintain strict alignment of inactive and active buttons in the control panels.

$LINKNAME = '"index.$EXTN"'; This is used when the $NO AUTO LINK variable is empty, to allow a URL to the working directory to be sufficient to reach the main page of the completed document. It specifies the name of the HTML file which will be automatically linked to the directory name. The value of $EXTN is .html unless $SHORTEXTN is set, in which case it is .htm .

$LINKPOINT = '"$FILE$EXTN"'; This specifies the name of the HTML file to be duplicated, or symbolically linked, with the name specified in $LINKNAME. At the appropriate time the value of $FILE is the document name, which usually coincides with the name of the working directory.

$CHARSET = 'iso 8859 1'; This specifies the character set used within the HTML pages produced by L^{A}TEX2HTML. If no value is set in a configuration or initialisation file, the default value will be assumed. The lowercase form $charset is also recognised, but this is overridden by the uppercase form.

$ACCENT IMAGES = 'large'; Accented characters that are not part of the ISO-Latin fonts can be generated by making an image using L^{A}TEX. This variable contains a (commaseparated) list of L^{A}TEX commands for setting the style to be used when these images are made. If the value of this variable is empty then the accent is simply ignored, using an un-accented font character (not an image) instead.

Within the color.perl package, the following variables are used to identify the names of files containing specifications for named colors. Files having these names are provided, in the $LATEX2HTMLSTYLES directory, but they could be moved elsewhere, or replaced by alternative files having different names. In such a case the values of these variables should be altered accordingly.

$RGBCOLORFILE = 'rgb.txt';
$CRAYOLAFILE = 'crayola.txt';

The following variables may well be altered from the system defaults, but this is best done using a local .latex2html-init initialisation file, for overall consistency of style within documents located at the same site, or sites in close proximity.

$default language = 'english'; This establishes which language code is to be placed within the <!DOCTYPE...> tag that may appear at the beginning of the HTML pages produced. Loading a package for an alternative language can be expected to change the value of this variable. See also the $TITLES LANGUAGE variable, described next.

$TITLES LANGUAGE = 'english'; This variable is used to specify the actual strings used for standard document sections, such as “Contents”, “References”, “Table of Contents”, etc. Support for French and German titles is available in corresponding packages. Loading such a package will normally alter the value of this variable, as well as the $default language variable described above.

$WORDS IN NAVIGATION PANEL TITLES = 4; Specifies how many words to use from section titles, within the textual hyperlinks which accompany the navigation buttons.

$WORDS IN PAGE = 450; Specifies the minimum page length required before a navigation panel is placed at the bottom of a page, when the $AUTO NAVIGATION variable is set.

$CHILDLINE = "<BR><HR>\n"; This gives the HTML code to be placed between the childlinks table and the ordinary contents of the page on which it occurs.

$NETSCAPE HTML = 0; When set, this variable specifies that HTML code may be present which does not conform to any official standard. This restricts the contents of any <!DOCTYPE...> tag which may be placed at the beginning of the HTML pages produced.

$BODYTEXT = ''; The value of this variable is used within the <BODY...> tag; e.g. to set text and/or background colors. It’s value is overridden by the \bodytext command, and can be added-to or parts changed using the \htmlbody command or \color and \pagecolor from the color package.

$INTERLACE = 1; When set, interlaced images should be produced.

This requires graphics utilities to be available to perform the interlacing operation.

$TRANSPARENT FIGURES = 1; When set, the background of images should be made transparent; otherwise it is white. This requires graphics utilities to be available which can specify the color to be made transparent.

$FIGURE SCALE FACTOR = 1; Scale factor applied to all images of figure and other environments, when being made into a bitmapped image.

Note that this does not apply to recognised mathematics environments, which instead use the contents of $MATH SCALE FACTOR and $DISP SCALE FACTOR to specify scaling.

$MATH SCALE FACTOR = 1; Scale factor applied to all images of mathematics, both inline and displayed. A value of 1.4 is a good alternative, with anti-aliased images.

$DISP SCALE FACTOR = 1; Extra scale factor applied to images of displayed math environments. When set, this value multiplies $MATH SCALE FACTOR to give the total scaling. A value of ‘1.2’ is a good choice to accompany $MATH_SCALE_FACTOR = 1.4; .

$EXTRA IMAGE SCALE This may hold an extra scale factor that can be applied to all generated images. When set, it specifies that a scaling of $EXTRA IMAGE SCALE be applied when images are created, but to have their height and width recorded as the un-scaled size. This is to coax browsers into scaling the (usually larger) images to fit the desired size; when printed a better quality can be obtained. Values of ‘1.5’ and ‘2’ give good print quality at 600dpi.

$PAPERSIZE = 'a5'; Specifies the size of a page for typesetting figures or displayed math, when an image is to be generated. This affects the lengths of lines of text within images. Since images of text or mathematics should use larger sizes than when printed, else clarity is lost at screen resolutions, then a smaller paper-size is generally advisable. This is especially so if both the $MATH SCALE FACTOR and $DISP SCALE FACTOR scaling factors are being used, else some images may become excessively large, including a lot of blank space.

$LINE WIDTH = 500; Formerly specified the width of an image, when the contents were to be right- or center-justified. (No longer used.)

The following variables are used to access the utilities required during image-generation. File and program locations on the local system are established by the configure-pstoimg Perl script and stored within $LATEX2HTMLDIR/local.pm as Perl code, to be read by pstoimg when required.

After running the configure-pstoimg Perl script it should not be necessary to alter the values obtained. Those shown below are what happens on the author’s system; they are for illustration only and do not represent default values.

   • $GS LIB = '/usr/local/share/ghostscript/4.02';
   • $PNMCAT = '/usr/local/bin/pnmcat';
   • $PPMQUANT = '/usr/local/bin/ppmquant';
   • $PNMFLIP = '/usr/local/bin/pnmflip';
   • $PPMTOGIF = '/usr/local/bin/ppmtogif';
   • $HOWTO TRANSPARENT GIF = 'netpbm';
   • $GS DEVICE = 'pnmraw';
   • $GS = '/usr/local/bin/gs';
   • $PNMFILE = '/usr/local/bin/pnmfile';
   • $HOWTO INTERLACE GIF = 'netpbm';
   • $PBMMAKE = '/usr/local/bin/pbmmake';
   • $PNMCROP = '/usr/local/bin/pnmcrop';
   • $TMP = '/usr/var/tmp';

The following variables are no longer needed, having been replaced by the more specific information obtained using the Perl script configure-pstoimg.

     $USENETPBM = 1;
     $PBMPLUSDIR = '/usr/local/bin';

5.3 Extending the Translator

In Section 3.7 is was seen how the capabilities of L^{A}TEX2HTML can be extended to cope with L^{A}TEX commands defined in packages and style-files. This is in addition to defining simple macros and environments, using \newcommand and \newenvironment. The problem however, is that writing such extensions for packages requires an understanding of Perl programming and of the way the processing in L^{A}TEX2HTML is organised.

The default behaviour for unrecognised commands is for their arguments to remain in the HTML text, whereas the commands themselves are passed on to L^{A}TEX, in an attempt to generate an image. This is far from ideal, for it is quite likely to lead to an error in L^{A}TEX due to not having appropriate arguments for the command.

Currently there are several related mechanisms whereby a user can ask for particular commands and their arguments to be either

   • ignored (via: &ignore commands );

• passed on to L^{A}TEX for processing as an image

     (via: &process commands in tex );

• passed to L^{A}TEX for processing as an image, as if an environment

     (via: &process commands inline in tex );

• passed on to L^{A}TEX, setting parameters to be used in subsequent images

     (via: &process commands nowrap in tex );

• processed as “order-sensitive”, as if an environment rather than a command

     (via: &process commands wrap deferred );

The string beginning &.... is the name of the Perl subroutine that controls how the L^{A}TEX commands are to be subsequently treated during processing by L^{A}TEX2HTML. Towards the end of the latex2html script, one finds a list of L^{A}TEX commands to be handled by each of these subroutines. These lists even include some common TEX commands.

Analogous lists occur in most of the package extension files. In many cases the commands are for fine-tuning the layout on a printed page. They should simply be ignored; but any parameters must not be allowed to cause unwanted characters to appear on the HTML pages. Customised extensions using these mechanisms may be included in the

$LATEX2HTMLDIR/latex2html.config file, or in a personal $HOME/.latex2html-init ini-

tialisation file, as described next.

5.3.1 Asking the Translator to Ignore Commands

Commands that should be ignored may be specified in the .latex2html-init file as input to the &ignore_commands subroutine. Each command which is to be ignored should be on a separate line followed by compulsory or optional argument markers separated by #’s e.g.^{12}:

     <cmd_name1> # {} # [] # {} # [] ...
     <cmd_name2> # <<pattern>> # [] ...

{}’s mark compulsory arguments and []’s optional ones, while <<pattern>> denotes matching everything up to the indicated string-pattern, given literally (e.g. \\endarray); spaces are ignored. Special characters such as $ , & , \ itself and perhaps some others, need to be “escaped” with a preceding \ .

Some commands may have arguments which should be left as text even though the command should be ignored (e.g. \hbox, \center, etc.). In these cases arguments should be left unspecified. Here is an example of how this mechanism may be used:

    &ignore_commands( <<_IGNORED_CMDS_);
    documentstyle # [] # {}
    linebreak # []
    mbox
    <add your commands here>
    _IGNORED_CMDS_

12It is possible to add arbitrary(?) Perl code between any of the argument markers which will be executed when the command is processed. For this however a basic understanding of how the translator works and of course Perl is required.

5.3.2 Asking the Translator to Pass Commands to L^{A}TEX

Commands that should be passed on to L^{A}TEX for processing because there is no direct translation to HTML may be specified in the .latex2html-init file as input to the process_commands_in_tex subroutine. The format is the same as that for specifying commands to be ignored. Here is an example:

    &process_commands_in_tex (<<_RAW_ARG_CMDS_);
    fbox # {}
    framebox # [] # [] # {}
    <add your commands here>
    _RAW_ARG_CMDS_

5.3.3 Handling “order-sensitive” Commands

Some commands need to be passed to L^{A}TEX, but using the &process_commands_in_tex subroutine gives incorrect results. This may occur when the command is “order-sensitive”, using information such as the value of a counter or a boolean expression (or perhaps requiring a box to have been constructed and saved). Try using the &process_commands_inline_in_tex subroutine instead. Commands declared this way are first “wrapped” within a dummy environment, which ensures that they are later processed in correct order with other environments and order-sensitive commands.

Other commands may need to be passed to L^{A}TEX, not to create an image themselves, but to affect the way subsequent images are created. For example a color command such as \color{red} should set the text-colour to ‘red’ for all subsequent text and images. This must be sent to L^{A}TEX so that it is processed at exactly the right time; i.e. before the first image required to be ‘red’ but following any images that are not intended to be affected by this colour-change. The subroutine process_commands_nowrap_in_tex is designed specifically to meet such requirements.

Commands can be order-sensitive without having to be passed to L^{A}TEX. Indeed even if a Perl subroutine has been carefully written to process the command, it may still give wrong results if it is order-sensitive, depending on the values of counters, say. To handle such cases there is the &process_commands_wrap_deferred subroutine. This also “wraps” the command within a dummy environment, but when that environment is processed the contents are not sent to L^{A}TEX, as in the previous case. All of the standard L^{A}TEX commands to change, set or read the values of counters are handled in this way.

5.4 Customising the Navigation Panels

The navigation panels are the strips containing “buttons” and text that appears at the top and perhaps at the bottom of each generated page and provides hypertext links to other sections of a document. Some of the options and variables that control whether and where it should appear are presented in Section 5.2.4.

A simple mechanism for appending customised buttons to the navigation panel is provided by the command \htmladdtonavigation. This takes one argument which L^{A}TEX2HTML appends to the navigation panel. For example,

    \htmladdtonavigation{\htmladdnormallink
       {\htmladdimg{http://server/mybutton.gif}}{http://server/link}}

will add an active button mybutton.gif pointing to the specified location. Apart from these facilities it is also possible to specify completely what appears in the navigation panels and in what order. As each section is processed, L^{A}TEX2HTML assigns relevant information to a number of global variables. These variables are used by the subroutines

top navigation panel and bottom navigation panel, where the navigation panel is con-

structed as a string consisting of these variables and some formatting information. These subroutines can be redefined in a system- or user-configuration file (respectively,

$LATEX2HTMLDIR/latex2html.config and $HOME/.latex2html-init). Any combination

of text, HTML tags, and the variables mentioned below is acceptable.

The control-panel variables are:

Iconic links (buttons):

$UP Points up to the “parent” section;

$NEXT Points to the next section;

$NEXT GROUP Points to the next “group” section;

$PREVIOUS Points to the previous section;

$PREVIOUS GROUP Points to the previous “group” section;

$CONTENTS Points to the contents page if there is one;

$INDEX Points to the index page if there is one.

Textual links (section titles):

$UP TITLE Points up to the “parent” section;

$NEXT TITLE Points to the next section;

$NEXT GROUP TITLE Points to the next “group” section;

$PREVIOUS TITLE Points to the previous section;

$PREVIOUS GROUP TITLE Points to the previous “group” section.

If the corresponding section exists, each iconic button will contain an active link to that section. If the corresponding section does not exist, the button will be inactive. If the section corresponding to a textual link does not exist then the link will be empty.

The “next group” and “previous group” are rarely used, since it is usually possible to determine which are the next/previous logical pages in a document. However these may be needed occasionally with segmented documents, when the segments have been created with different values for the $MAX SPLIT DEPTH variable. This is quite distinct from the segmented document effect in which the first page of one segment may have its ‘PREVIOUS’ button artificially linked to the first page of the previous segment, rather than the last page. The number of words that appears in each textual link is controlled by the variable $WORDS IN NAVIGATION PANEL TITLES which may also be changed in the configuration files. Below is an example of a navigation panel for the bottom of HTML pages. (Note that the “.” is Perl’s string-concatenation operator and “#” signifies a comment.)

    sub bot_navigation_panel {
        #  Start with a horizontal rule and descriptive comment
        "<HR>\n" . "<!--Navigation Panel-->".
        # Now add a few buttons, with a space between them
           "$NEXT $UP $PREVIOUS $CONTENTS $INDEX $CUSTOM_BUTTONS" .
        # Line break

           "\n<BR>" .
        # If ``next'' section exists, add its title to the navigation panel
           ($NEXT_TITLE ? "\n<B> Next:</B> $NEXT_TITLE" : undef) .
        # Similarly with the ``up'' title ...
           ($UP_TITLE ? "\n<B>Up:</B> $UP_TITLE\n" : undef) .
        # ... and the ``previous'' title
           ($PREVIOUS_TITLE ? "\n<B> Previous:</B> $PREVIOUS_TITLE\n" : undef) .
    }

Note that extra buttons may be included by defining suitable code for the container $CUSTOM BUTTONS . The use of explicit ‘newline’ (\n) characters is not necessary for the on-screen appearance of the navigation panel within a browser window. However it maintains an orderly organisation within the .html files themselves, which is helpful if any hand-editing is later required, or simply to read their contents. The corresponding subroutine for a navigation-panel at the top of a page need not use the rule <HR>, and would require a break (<BR>) or two at the end, to give some visual separation from the following material.

6 Known Problems

Here are some of the problems that were known to exist with previous versions of L^{A}TEX2HTML. Most of those that were real errors are either fixed completely in more recent versions (v98.1 and later), or are much less likely to occur within correct L^{A}TEX source. (Some are not really errors but indications of poor style in the user’s choices among various ways to organise their source code.) Several are indeed limitations inherent in the way L^{A}TEX2HTML currently performs its processing.

Unrecognised Commands and Environments: Unrecognised commands are ignored and any arguments are left in the text. Unrecognised environments are passed to L^{A}TEX and the result is included in the document as one or more inlined images. There are very few standard L^{A}TEX commands that are not recognised. Many common TEX commands are recognised also, even though not explicitly mentioned in the L^{A}TEX blue book[1]. Any aberrant commands should be reported to the L^{A}TEX2HTML mailing list, see Section 2.5.

Index: The translator generates its own index by saving the arguments of the \index command. The contents of the theindex environment are ignored. When using the makeidx package, very sophisticated Indexes can be built automatically. The Index for this manual is a good example.

New Definitions: New definitions (\newcommand, \newenvironment, \newtheorem and

\def), will not work as expected if they are defined more than once. Only the last definition will be used throughout the document. Stylistically it is bad to declare new environments or theorems outside of the document preamble, so this should not cause any real problems. Changes to commands using \def or \renewcommand should usually be made only locally, within special environments, to set a needed parameter; e.g. a basic length in a picture environment. But when such environments force an image to be generated, then L^{A}TEX will make the correct redefinition.

Scope of declarations and environments: If the scope of a declaration or environment crosses section boundaries, then the output may not be as expected, because each section is processed independently.

This is inherent to the way L^{A}TEX2HTML does its processing. It will not be different unless later versions change this strategy; (e.g. if L^{A}TEX2HTML-NG ever becomes fully integrated.)

Math-mode font-size changes: Math-mode font changes made outside the math-mode are not honoured. Thus the two equations in $a_b$ and {\LARGE $a_b$} would come out looking the same. The trick is to write $a_b and $\mbox{\LARGE $a_b$}$.

6.1 Troubleshooting

Here are some curable symptoms:

Cannot run any of the Perl programs: If your Perl installation is such that Perl programs are not allowed to run as shell scripts you may be unable to run latex2html, texexpand, pstoimg and install-test. In this case change the first line in each of these programs from

    #!/usr/local/bin/perl

to

    : # *-*-perl-*-*
        eval 'exec perl -S  $0 "$@"'
        if $running_under_some_shell;

The install-test script gives uninformative error messages: If, for any reason, you have trouble running install-test, do not despair. Most of what it does is to do with checking your installation rather than actually installing anything. To do a manual installation just change the variable $LATEX2HTMLDIR in the beginning of the file latex2html to point to the directory where the L^{A}TEX2HTML files can be found.

Also, make sure that the files pstoimg, texexpand and latex2html are executable; if necessary use the Unix chmod command to make them executable.

It just stops. Check the style files that you are using. It is possible that you are using a style file which contains raw TEX commands. In such a case start L^{A}TEX2HTML with the option ‘ -dont include <style-file name> ’. Alternatively, add the name of the style to the variable $DONT INCLUDE in your $HOME/.latex2html-init file. If you don’t have such a file then create one and add the lines:

     $DONT_INCLUDE = "$DONT_INCLUDE" . ":<style file name>";
     1;        # This must be the last line

Another reason why L^{A}TEX2HTML might stop is that the L^{A}TEX source file itself contains raw TEX commands. In this case you may put such commands inside a latexonly environment (see Section 4.4).

The $VERBOSITY variable can be used to create tracing messages, which may help to locate which command or environment was being processed when everything stopped.

It appears to be doing nothing. Perhaps the processor has fallen into an unending loop. Usually there will be a bad definition, or other faulty source code, which has caused this. The $VERBOSITY variable (see page 67)can be set to generate tracing messages, which may help to locate which command or environment is being processed repeatedly. Try setting a value of ‘3’; e.g. using the commandline switch ‘ -verbosity 3 ’. This will print command and environment names, as thaey are processed. It should soon become apparent where any such looping occurs.

It just fills the endlessly with dots. No ‘perhaps’ here; the processor has definitely fallen into an unending loop. See the previous item for how to detect which command or environment is causing the problem.

Perl cannot parse the latex2html script: If Perl refuses to start L^{A}TEX2HTML and issues errors, your Perl version is not up to date. Update your Perl to 5.003 or later. You can check which version of Perl you are using by invoking Perl with the ‘ -v ’ option.

If Perl issues errors during runtime, this is most probably related to bugs within L^{A}TEX2HTML or one of its modules. In this case you will need help from the developers or experienced users; this can be obtained via the discussion list.

It does not show any of your images: You can’t run L^{A}TEX2HTML in a subdirectory that contains a dot within the directory name, such as latex2html-98.1, or in name of any higher directory. This is because dvips’s -o option will change 98.1 into 98.001 and use that as the resulting output file, instead of image001 . The PostScript files will be placed higher up in the directory tree.

For instance, if pwd returns something like:

   /usr/people/nelson/files/latex2html-98.1/work/tests

and you run L^{A}TEX2HTML, then dvips will generate image output files here:

   /usr/people/nelson/files
called latex2html-98.001, latex2html-98.002, . . . instead of image001, image002,

image003, … in the subdirectory where your .html files were created. As a result the images will not show in your documents.

If you are getting ‘File Not Found’ errors, then reprocess your job using the ‘ -debug ’ switch, to see what options are passed to dvips. If there is a ‘.’ in some parts of any directory name, then look above that directory to see if files are being generated there.

One obvious fix is to rename the offending directory to remove the ‘.’ from its name.

If that is not possible, then define an alternative location for image generation to take place; set $TMP to contain the name for this location. Typically $TMP = '/usr/tmp'; . (This use of $TMP is a good thing to do anyway, especially if your Unix account is subject to quota limitations.)

It stops after having run L^{A}TEX, displaying a message about dvips: See the previous item for the reason and ‘fix’. This version of L^{A}TEX2HTML detects the problem and will not allow images to be generated in the wrong place.

dvips complains about incorrect arguments: Please use a version which supports the command-line options ‘ -M ’, ‘ -S ’, ‘ -o ’ and ‘ -i ’. “Recent” versions, at least after 5.516, do support them.

It gives an “Out of memory” message and dies: Try splitting your source file into more than one file, using the L^{A}TEX commands \input or \include. Also, try using the ‘ -no images ’ option.

Perhaps the processor has fallen into an infinite loop. Usually there will be a bad definition, or other faulty source code, which has caused this. See an earlier problem for how to set the $VERBOSITY variable to help locate the bad code leading to this memory exhaustion.

As a last resort you may consider increasing the virtual memory (swap space) of your machine.

install-test issues “dbm” related error messages:

L^{A}TEX2HTML requires a DataBase Management system (NDBM, GDBM, or SDBM) in order to run. This is usually part of each Unix-like operating system and SDBM is part of Perl 5, but sometimes this is either missing on your operating system or left out in a binary Perl distribution.

latex2html issues “dbm” related error messages: If you get warnings like

   ndbm store returned -1, errno 28, key "xyz" at latex2html line 123

this is related to an overflow of L^{A}TEX2HTML internals. You will need help from the list, here.

If you get real error messages which cause L^{A}TEX2HTML to abort, run install-test to check if your DataBase management works. You will probably need to re-install Perl 5 (see above topic).

This can happen when an image is being created from a large piece of L^{A}TEX source code. The image-reuse mechanism uses the code itself to construct a database key. If too long, the key is invalid and may crash DBM or NDBM. (In fact this error should no longer occur in v97.1 , so please advise the L^{A}TEX2HTML developers if you get error messages of this kind.) The message should contain the name of environment which caused the problem, along with an identifying number; e.g. eqnarray268. To find which exact piece of code this represents, run L^{A}TEX2HTML again, using the ‘ -debug ’ switch. Then look at the files in the TMP subdirectory of the working directory

named TMP/part001, TMP/part002, etc. Use the unix grep command: grep 268

<dir>/TMP/part* to find that number in these files. This should enable you to locate exactly where the problem occurs.

One solution may be to wrap the whole environment within \begin{makeimage} and \end{makeimage}. This will still cause the image to be created, but uses just the environment name and number as the database key.

The \verb"ABC" command doesn’t work: This is an unfortunate bug which can be avoided by using any character other than quotes; e.g. \verb+ABC+.

Cannot get the “tilde” (˜ ) to show: The  trick  here  is  to
use  \~{}.  Alternatively  try  using  something  like:

{\htmladdnormallink{mylink}{http://host/\%7Eme/path/file.html}}.

Warning: Some browsers may not be able to interpret the %7E as a “tilde” character.

Try using \char126 . In any case tildes within \htmladdnormallink and similar commands should be handled correctly.

Macro definitions don’t work correctly: As mentioned in other places, not all plain TEX \def-initions can be converted. But you may also have problems even when using L^{A}TEX definitions (with \newcommand and \newenvironment) if such definitions make use of sectioning or \verbatim commands. These are handled in a special way by L^{A}TEX2HTML and cannot be used in macro definitions.

\input commands: There is a bug in the expansion of \input commands which may cause a problem when more than one \input command appears on the same line. There is no quick fix other than suggesting that you put each \input command on a line by itself, in the L^{A}TEX source files.

\input commands in verbatim environments: These should no longer cause problems (actually since 97.1). \input commands are also handled correctly within comment environments declared using \excludecomment and when loading multiple .aux files, due to use of the \include command. Alternatively you might want to use either the verbatim or the verbatimfiles package.

Optional arguments in description environments: If you have optional arguments for the \item command in a description environment containing nested “]” characters then these may not show up correctly. To avoid the problem enclose them in {}s;

e.g. \item[{[nested [angle [brackets] are ok]]}]

L^{A}TEX2HTML behaves differently even when you run it on the same file: If you notice any strange side-effects from previous runs of L^{A}TEX2HTML, try using the option ‘ -no reuse ’ and choose (d) when prompted. This will clear any intermediate files generated during previous runs. Note that this option will disable the image-reuse mechanism.

Cannot convert PostScript images which are included in the L^{A}TEX file: It is likely that the macros you are using for including PostScript files (e.g. \epsffile) are not understood by L^{A}TEX2HTML. To avoid this problem enclose them in an environment which will be passed to L^{A}TEX anyway; e.g.

     \begin{figure}
     \epsffile{<PostScript file name>}
     \end{figure}

Another reason why this might happen is that your shell environment variable TEXINPUTS may be undefined. This is not always fatal but if you have problems you can use full path-names for included PostScript files (even when the PostScript files are in the same directory as the L^{A}TEX source file). Alternatively try setting TEXINPUTS to ‘.::’. With some TEX and L^{A}TEX installations setting TEXINPUTS to ‘.::’ may cause problems in the normal operation of L^{A}TEX. If you get errors such as L^{A}TEX complaining that it can no longer find any style files then you must set

TEXINPUTS to "<path to your LaTeX installation>:." if you want to use both

L^{A}TEX and L^{A}TEX2HTML.

Some of the inlined images are in the wrong places: There are several known ways that this may occur.

• Perhaps one of the inlined images is more than a page (paper page) long. This is sometimes the case with very large tables or large PostScript images. In this case you can try specifying a larger paper size (e.g. ‘a4’, ‘a3’ or even ‘a0’) instead of the default (‘a5’) using the L^{A}TEX2HTML variable $PAPERSIZE in the

file latex2html.config.

This error should no longer occur, with v97.1 . Please report it on the mailinglist, if it does.

• More likely is that some inappropriate L^{A}TEX code has caused an error, resulting in an extra page (or pages) being generated. Examine the images.log file, to see if it reports any L^{A}TEX errors.

• A much rarer reason is that by default the dvips program reverses the PostScript pages it generates. If your dvips program behaves in this way try changing

the line: $DVIPS = "dvips"; to: $DVIPS = "dvips -r0"; within the file
latex2html.config.

• Yet another reason for images appearing out of place, especially while developing a document, is that the browser’s image cache is providing out-of-date versions rather than getting the latest version afresh. When this occurs there will often be images stretched or shrunk to fit the wrong sized imaging area; this symptom is browser-dependent. Flushing the cache, then reloading the HTML document, should clear up the problem.

Unacceptable quality of converted images: Try changing the size of the image (see Section 3.4).

The bibliographic references are missing: Run latex and then bibtex on the original source file in order to generate a .bbl file. L^{A}TEX2HTML may need the .bbl file in order to generate the references.

The labels of figures, tables or equations are wrong: This can happen if you have used any figures, tables, equations or any counters inside conditional text; i.e. in a latexonly or a htmlonly environment.

Problems after changing the configuration files: Please make sure that the last line

in the configuration files (i.e. .latex2html-init and latex2html.config) is:
1;      # This is the last line

This is a Perl quirk.

Problems when producing the .dvi version: If you are using any of the new L^{A}TEX commands which are defined in the html.sty file make sure that html.sty is included; e.g. as one of the optional arguments to the \documentclass command. Of course you also have to make sure that L^{A}TEX knows where the html.sty file is, either by putting it in the same place as the other style-files on your system, or by changing your TEXINPUTS shell environment variable ^{13}.

Some of the fonts are translated incorrectly: There is a fault in way the L^{A}TEX scoping rules have been interpreted in L^{A}TEX2HTML. Consider this:

\ttfamily fixed-width font.
\begin{something}
nothing here
\end{something}
default font.

13If you don’t know how to do either of these things, copy (or link) html.sty to the directory of your L^{A}TEX document.

When processed by L^{A}TEX, the effect of the \tt command is delimited by the beginning of the environment “something”, so that “default font” will appear in the default font. But L^{A}TEX2HTML will not recognise “something” as a delimiter and “default font” will appear in the wrong font. To avoid this problem (until it is fixed) you may delimit the scope of some commands explicitly using {}’s; i.e.

    \texttt{fixed-width font}.
    \begin{something}
    nothing here
    \end{something}
    default font.

Nesting of font changing commands is now handled correctly. Such problems should no longer occur, though it always helps to use explicitly delimited ‘pseudo’-environments; e.g. {\bf ... }, or L^{A}TEX’s commands requiring an explicit argument, such as

\textbf.

Cannot get it to generate inlined images: Run L^{A}TEX2HTML on your document, using the ‘ -debug ’ switch. Look in the directory where the HTML files are generated for two files named images.tex and images.log. Examine their contents. Do you notice anything unusual in them?

Copy images.tex into the directory of your original L^{A}TEX file and run latex on images.tex. Can you see any errors in images.log? If so, can you fix images.tex to get rid of the errors? After fixing images.tex you can put it back in the directory of HTML, then run L^{A}TEX2HTML on the original document using the ‘ -images only ’ switch.

However if you make changes or additions to the original source then the same problems may occur again. Thus it is better to understand why the changes were required and alter your document’s source code appropriately.

If you get into a mess delete all the image files and run L^{A}TEX2HTML again. Often it is sufficient to just delete the file images.pl. If you still get into a mess, try running L^{A}TEX2HTML with the options ‘ -no reuse ’ and

‘ -no images ’; e.g.
    cblipca% latex2html -no_reuse -no_images test.tex
    This is LaTeX2HTML Version 95 (Tue Nov 29 1994) by Nikos Drakos,
    Computer Based Learning Unit, University of Leeds.
    OPENING /tmp_mnt/home/cblelca/nikos/tmp/test.tex
    Cannot create directory /usr/cblelca/nikos/tmp/test: File exists
    (r) Reuse the images in the old directory OR
    (d) *** DELETE *** /usr/cblelca/nikos/tmp/test AND ITS CONTENTS OR
    (q) Quit ?
    :d
    Reading ...
    Processing macros ....+.
    Reading test.aux ......................
    Translating ...0/1........1/1.....

    Writing image file ...
    Doing section links .....
    *********** WARNINGS ***********
    If you are having problems displaying the correct images with Mosaic,
    try selecting "Flush Image Cache" from "Options" in the menu-bar
    and then reload the HTML file.
    Done.

Then try to have a look in the file images.tex (as described earlier) and perhaps fix it. Once you are happy that images.tex is OK, run L^{A}TEX2HTML again with the option

‘ -images only ’.

Some problems in displaying the correct inlined images, may be due to the image caching mechanisms of your browser. With some browsers a simple “Reload Current Document” will be enough to refresh the images but with others (e.g. Mosaic) you may need to request for the cache to be refreshed. With Mosaic try selecting “Flush Image Cache” from “Options” in the menu-bar and then reload the HTML file.

It cannot do slides, memos, etc. If you use SliTEX you can go a long way just by replacing the {slides} argument of the \documentclass command with something like {article} just before using L^{A}TEX2HTML. One problem may be that all your slides will end up in the same HTML file. If you use lslide.sty you may get much better results.

7 Credits

Beginnings, 1993–1994

L^{A}TEX2HTML was written by Nikos Drakos at the Computer Based Learning Unit, University of Leeds. Several people have contributed suggestions, ideas, solutions, support and encouragement. Some of these are Roderick Williams, Ana Maria Paiva, Jamil Sawar and Andrew Cole at the Computer Based Learning Unit.

CERN The idea of splitting L^{A}TEX files into more than one component, connected via hyperlinks, was first implemented in Perl by Toni Lantunen at CERN. Thanks to Robert Cailliau of the World-Wide Web Project, also at CERN, for providing access to the source code and documentation (although no part of the original design or the actual code has been used).

Robert S. Thau has contributed the new version of texexpand. Also, in order to translate the “document from hell” (!!!) he has extended the translator to handle \def commands, nested math-mode commands, and has fixed several bugs.

Phillip Conrad and L. Peter Deutsch. The pstogif Perl script uses the pstoppm.ps PostScript program, originally written by Phillip Conrad (Perfect Byte, Inc.) and modified by L. Peter Deutsch (Aladdin Enterprises).

Roderick Williams The idea of using existing symbolic labels to provide cross-references between documents was first conceived during discussions with Roderick.

Eric Carroll who first suggested providing a command like \hyperref .

Franz Vojik provided the basic mechanism for handling foreign accents.

Todd Little The ‘ -auto navigation ’ option was based on an idea by Todd.

Axel Belinfante provided the Perl code in the makeidx.perl file, as well as numerous suggestions and bug-reports.

Verena Umar (from the Computational Science Education Project (ORNL)) has been a very patient tester of some early versions of L^{A}TEX2HTML and many of the current features are a result of her suggestions.

Ian Foster and Bob Olson. Thanks to Ian Foster and Bob Olson at the Argonne National Labs, for setting up the mailing list.

Later Developments, 1995–1996

Since 1995 the power and usefulness of L^{A}TEX2HTML has been enhanced significantly. The revisions later than v95.1 have been largely due to the combined efforts of many people, other than the original author. Interested users have supplied patches to fix a fault, or implement a feature that previously was not supported. Often a question or complaint to the discussion-group (see Section 2.5) has spurred someone else to provide the necessary “patch”.

Arising from this work, special credit is due to:

Marcus Hennecke for his many extensive revisions; Mark Noworolski for coordinating v95.3;

Sidik Isani for his improvement in GIF quality;

Michel Goossens was the driving force behind the upgrade to L^{A}TEX 2ε compatibility, and other features developed at CERN;

Herb Swan for coordinating v96.1 of L^{A}TEX2HTML, including much of the Perl code for the new features that were introduced, and for providing a series of bug-fix revisions prior to v96.1 rev-f;

Ross Moore who has revised and extended this manual, helped design and test the segmentation strategy, and later revisions of v96.1 . Ross organised the release of v96.1 rev-g and provided many of the improvements incorporated into v96.1 rev-h.

Martin Wilck for the initial work on implementation of frames. Also Martin did most of the work implementing the extensive citation and bibliographic features of the natbib package, written by Patrick Daly. He also provided the makeseg Perl script to create Makefiles for segmented documents.

Jens Lippmann for organising the releases v96.1 rev-h to v98.1. Jens made significant contributions to the internal workings of L^{A}TEX2HTML, as well as cleaning up much of its source code.

Many others, too many to mention, contributed bug-reports, fixes and other suggestions.

Thanks also to Donald Arseneau for allowing his url.sty to be distributed with this manual. Similarly, thanks to Johannes Braams for changebar.sty. Both of these are useful utilities which enhance the appearance of the printed manual. In particular, changes introduced with v98.1 and its revisions are denoted by thin change-bars, while thicker bars denote changes introduced with v98.2 and later releases.

Developments: late 1996 to mid 1997

During the latter part of 1996 there was much work on improving the capabilities of L^{A}TEX2HTML. Some of this was due to the World Wide Web Consortium^{14}’s proposals for HTML 3.2, becoming a formal recommendation in November 1996, and their subsequent acceptance in January 1997. Existing L^{A}TEX markup for effects such as centering, left- or right-justification of paragraphs, flow of text around images, table-layout with formal captions, etc. could now be given a safe translation into HTML 3.2, compliant with a standard that would guarantee that browsers would be available to view such effects.

At the same time developers were exploring ways to enhance the overall performance of L^{A}TEX2HTML. As a result the current v97.1 release has significant improvements in the following areas:

image-generation is much faster, requires less memory and inline images are aligned more accurately;

image quality is greatly improved by the use of anti-aliasing effects for on-screen clarity, in particular with mathematics, text and line-drawings; memory-requirements are much reduced, particularly with image-generation;

14http://www.w3c.org/

mathematics can now be handled using a separate parsing procedure; images of sub-parts of expressions can be created, rather than using a single image for the whole formula;

macro definitions having a more complicated structure than previously allowed, can now be successfully expanded;

counters and numbering are no longer entirely dependent on the .aux file generated by L^{A}TEX;

decisions about which environments to include or exclude can now be made;

HTML effects for which there is no direct L^{A}TEX counterpart can be requested in a variety of new ways;

HTML code produced by the translator is much neater and more easily readable, containing more comments and fewer redundant breaks and <P> tags.

error-detection of simple L^{A}TEX errors, such as missing or unmatched braces, is now performed — a warning message shows a line or two of the source code where the error has apparently occurred;

For these developments, thanks goes especially to:

Jens Lippmann for creating and maintaining the CVS repository at https://www. github.com/latex2html/latex2html/ . This has made it much easier for the contributions from different developers to be collected and maintained as a “development version” which is kept up-to-date and available at all times. Together with Marek Rouchal he produced an extensive rewrite of the texexpand utility.

Ross Moore for extensive work on almost all aspects of the L^{A}TEX2HTML source and documentation, combining code for L^{A}TEX, Perl, HTML and other utilities. Most of the coding for the new features based on HTML 3.2, many of the new packages, faster image-generation and the improved support for mathematics and other environments, is his work.

Marek Rouchal for extending the former pstogif utility, transforming it into pstoimg which now allows for alternative image formats, such as PNG. Also he produced the neat configure-pstoimg script, which eases L^{A}TEX2HTML installation, and a rewrite

of texexpand.

Marcus Hennecke who has always been there, up-to-date with developments in HTML and related matters concerning Web publishing, and tackling the issues involved with portability of L^{A}TEX2HTML to Unix systems on various platforms.

Furthermore Marcus has produced L^{A}TEX2HTML-NG, a version of L^{A}TEX2HTML which handles expansion of macros in a more “TEX-like” fashion. This should lead to further improvements in speed and efficiency, while allowing complicated macro definitions to work as would be expected from their expansions under L^{A}TEX. (This requires Perl 5 , using some programming features not available with Perl 4 .)

Fabrice Popineau has produced an adaptation for the Windows NT platform, of L^{A}TEX2HTML v97.1 .

Uli Wortmann showed how to configure Ghostscript to produce anti-aliasing effects within images.

Axel Ramge for various suggestions and examples of enhancements, and the code to avoid a problem with Ghostscript.

Thanks also to all those who have made bug-reports, supplied fixes or offered suggestions as to features that might allow L^{A}TEX2HTML to be used more efficiently in particular circumstances. Most of these have been incorporated into this new version v97.1 , though perhaps not in the form originally envisaged. Such feedback has contributed enormously to helping make L^{A}TEX2HTML the easy to use, versatile program that it has now become.

Keep the ideas coming!

1st L^{A}TEX2HTML Workshop Darmstadt, 15 February 1997

Thanks again to Jens Lippmann and members of the LiPS Design Team for organising this meeting; also to the Fachbereich Informatik at Darmstadt for use of their facilities. This was an opportunity for many of the current L^{A}TEX2HTML developers to actually meet for the first time; rather than communication by exchange of electronic mail messages.

• Nikos Drakos talked about the early development of L^{A}TEX2HTML, while…

• … Ross Moore, Jens Lippmann and Marek Rouchal described recent improvements.

• Michel Goossens presented a list of difficulties encountered with earlier versions of L^{A}TEX2HTML, and aspects requiring improvement. Almost all of these now have been addressed in the v97.1 release, so far as is possible within the bounds inherent in the HTML 3.2 standard.

• Kristoffer Rose showed how it is possible to create GIF89 animations from pictures generated by TEX or L^{A}TEX, using the XY-pic graphics package and extensions.

Also present were representatives from the DANTE e.V. Praesidium and members of the L^{A}TEX3 development team. In all it was a very pleasant and constructive meeting.

TUG’97 — Workshop on L^{A}TEX2HTML University of San Francisco, 28 July 1997

On the Sunday afternoon (2.00pm–5.00pm) immediately prior to the TUG meeting, there will be a workshop on L^{A}TEX2HTML, conducted by Ross Moore^{15}.

Admission: $50, includes a printed copy of the latest L^{A}TEX2HTML manual.

15Mathematics Department, Macquarie University, Sydney, Australia

TEXNortheast TUG Conference, TEX/L^{A}TEX Now March 22–24, 1998, New York City

Includes a workshop/presentation by Ross Moore^{16}.

Euro-TEX’98, 10th European TEX Conference St. Malo, France — 29–31 March, 1998

Several of the L^{A}TEX2HTML developers will be present. All European (and other) L^{A}TEX2HTML users are encouraged to attend.

16Mathematics Department, Macquarie University, Sydney, Australia

Developments: late 1997 to early 1998

Much of the work contributed to L^{A}TEX2HTML during this time was related to bug fixing and maintaining the 97.1 release, in order to reach a more stable and reliable version which produces HTML code conforming to the W3C standards/drafts. To keep in context with this view, support for HTML 4 has been incorporated into the translator. There have been improvements to the way math code is handled, as well as font-changing and numbering commands. These now are expected to work much closer to the way that L^{A}TEX handles them. Furthermore, missing L^{A}TEX style translations for basic L^{A}TEX and AMS-TEX document classes were added to the distribution: book.perl, report.perl, article.perl, letter.perl, amsbook.perl and amsart.perl. New styles implementing L^{A}TEX packages

include seminar.perl, inputenc.perl and chemsym.perl naming but a few.

The aim is ultimately to support all L^{A}TEX, AMS-TEX etc. packages in the standard L^{A}TEX distribution, or for which there is published documentation. At the time of writing this aim has not quite been reached. To support internationalisation, Perl extensions were provided for HTML output conforming to ISO-Latin 1, 2, 3, 4, 5, 6, and Unicode^{17} encodings. All of the above work was done by Ross Moore.

Additional document formats are now supported, these are IndicTEX, FoilTEX, and CWEB documents. You may use any of these packages to translate such documents together with L^{A}TEX2HTML, refer to the instructions in the various README files. Thanks go to Ross Moore for IndicTEX/HTML, to Boris Veytsman for FoilTEX/HTML and to Jens Lippmann for the CWEB to HTML translator.

Numerous discussions and efforts have been undertaken to get L^{A}TEX2HTML working independent from the underlying operating system. Yet all obstacles are not quite taken, but it is forseeable that we are OS independent very soon. This release has been reported to run on OS/2, DOS, and MacOS, besides Unix-like operating systems. A former version has also been ported to Amiga OS, but that results still need to be re-integrated into the source. Ports for Windows’95 and Windows NT exist, but are not yet integrated with the main distribution. Thanks go to Marcus Hennecke, Axel Ramge, Marek Rouchal and Uli Wortmann for fruitful and refreshening discussions about that Override.pm loading scheme (which finally made its way after enough chickens and eggs chased one another to death ◯_{¨⌣} ), and to Daniel Taupin for his successful efforts to get L^{A}TEX2HTML running on DOS. Thanks go also to Fabrice Popineau for his port to Windows NT, and Nikos Drakos for a Windows 95 port based on v96.1h (which is mentioned here at last, but not least).

We want to take the opportunity to thank Scott Nelson and the people at Lawrence Livermore National Laboratory who help to keep up the L^{A}TEX2HTML main archive and the mailing list, and to Achim Bohnet at the Max Planck Institut fuer extraterrestrische Physik, Garching for maintaining the list’s online archive. Finally thanks and greetings to all people that contributed to this release and have not been mentioned here...

You all showed spirit and favour. Thank you for your efforts!

… and don’t forget Jens and the LiPS team at Darmstadt!

17http://www.unicode.org/

1998 to 1999

During this period large parts of L^{A}TEX2HTML have been overhauled and compatibility with Perl 4 broken once and for all. The 99.2 release is the first known to work out of the box on several UNIX systems as well as on Windows 95, 98, NT and OS/2. The number of supported L^{A}TEX packages is bigger than ever.

Thanks to Adalbert Perbandt for testing every second alpha/beta release of 99.2 on OS/2 and ensuring that things work ok there.

2018

• SVG image generation

• Support for pdflatex

• Support for HTML 5

Proposals for Future Development:

Extended Capabilities in Web browsers

The following areas are the subject of active development within the Web community. Limited support is available within L^{A}TEX2HTML for some of these features, using the -html version 4.0 command-line switch.

style-sheets: proposals for a flexible mechanism to allow cascading (CSS) and DSSSL, within HTML 4.0.

XML: eXtensible Markup Language^{18}.

MathML: Mathematical Markup Language^{19}.

Fonts: further support for non-standard font encodings.

Icons: Alternative sets of icons for navigation buttons and other purposes.

18http://www.w3.org/pub/WWW/TR/WD-xml.html

19http://www.w3.org/pub/WWW/TR/WD-math-970515

General License Agreement and Lack of Warranty

This software is distributed in the hope that it will be useful but without any warranty. The author(s) do not accept responsibility to anyone for the consequences of using it or for whether it serves any particular purpose or works at all. No warranty is made about the software or its performance.

Use and copying of this software and the preparation of derivative works based on this software are permitted, so long as the following conditions are met:

• The copyright notice and this entire notice are included intact and prominently carried on all copies and supporting documentation.

• No fees or compensation are charged for use, copies, or access to this software. You may charge a nominal distribution fee for the physical act of transferring a copy, but you may not charge for the program itself.

• If you modify this software, you must cause the modified file(s) to carry prominent notices (a ChangeLog) describing the changes, who made the changes, and the date of those changes.

• Any work distributed or published that in whole or in part contains or is a derivative of this software or any part thereof is subject to the terms of this agreement. The aggregation of another unrelated program with this software or its derivative on a volume of storage or distribution medium does not bring the other program under the scope of these terms.

This software is made available as is, and is distributed without warranty of any kind, either expressed or implied. In no event will the author(s) or their institutions be liable to you for damages, including lost profits, lost monies, or other special, incidental or consequential damages arising out of or in connection with the use or inability to use (including but not limited to loss of data or data being rendered inaccurate or losses sustained by third parties or a failure of the program to operate as documented) the program, even if you have been advised of the possibility of such damages, or for any claim by any other party, whether in an action of contract, negligence, or other tortuous action.

The L^{A}TEX2HTML translator is written by Nikos Drakos, Computer Based Learning Unit, University of Leeds, Leeds, LS2 9JT. Copyright ©1993–1997. All rights reserved.

The v97.1, v98.1, v98.2 and v99.1 revisions of the L^{A}TEX2HTML translator and this manual were prepared by Ross Moore, Mathematics Department, Macquarie University, Sydney 2109, Australia. Copyright ©1996–1999. All rights reserved.

Year 2000 compliance: L^{A}TEX2HTML contains *no* executable software, per se. It consists entirely of scripts to run other pieces of software: Perl, L^{A}TEX, Ghostscript, netpbm, etc. and standard Unix utilities (e.g. cp, rm, make, ln, ... ) as well as the operating system shell.

These other pieces of software are to be obtained and installed independent from the L^{A}TEX2HTML scripts.

L^{A}TEX2HTML makes no reference to dates, apart from reading the current date from the operating system, and converting the resulting string data into a standard form. This may result in ‘00’ appearing in the year 2000. However this representation of the date is used for display only; it does not control any further processing.

References

[1] Leslie Lamport, L^{A}TEX,A Document Preparation System. User’s Guide & Reference Manual, 2nd edition. ISBN 0-201-52983-1, Paperback 256 pages, Addison–Wesley, 1994. Online information on TEX and L^{A}TEX is available at http://www.tug.org/ and http:

//www.dante.de/ .

[2] Michel Goossens, Frank Mittelbach, Alexander Samarin, The L^{A}TEX Companion ISBN 0-201-54199-8, Paperback 530 pages, Addison–Wesley, 1994.

[3] Michel Goossens, Sebastian Rahtz and Frank Mittelbach, The L^{A}TEX Graphics Companion. ISBN 0-201-85469-4, Softcover 608 pages, Addison–Wesley, 1997.

[4] Michel Goossens and Sebastian Rahtz, with E. Gurari, R. Moore & R. Sutor. The L^{A}TEX Web Companion. ISBN 0-201-43311-7, Addison–Wesley, 1999.

[5] Nikos Drakos, Text to Hypertext conversion with L^{A}TEX2HTML. Baskerville, December 1993, Vol. 3, No. 2, pp 12–15.

[6] Nikos Drakos, From Text to Hypertext: A Post-Hoc Rationalisation of L^{A}TEX2HTML. Published in “Proceedings of the 1st World Wide Web Conference”, May 1994, CERN, Geneva, Switzerland.

Glossary of variables and file-names

.dvipsrc initialisation file for dvips . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 9 .latex2html-init initialisation file for L^{A}TEX2HTML . . . . . 9, 58, 62, 68, 70, 73, 74, 81 $ACCENT IMAGES holds the L^{A}TEX style to be used when making images of accented char-

acters which are not present in the character encoding being used . . . . . . . . . . . 70 $ADDRESS contains information to be included as a signature at the bottom of each HTML

page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 $ALTERNATIVE ICONS contains the location of special icons, other than the standard ones

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 68 $ANTI ALIAS when set, specifies that anti-aliasing be used with generated images of figures

and included images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 65 $ANTI ALIAS TEXT when set (the default), specifies that anti-aliasing be used with gener-

ated images of maths and typeset text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 $ASCII MODE when set, all icons and images are replaced with text strings, suitable for

text-only browsers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 $AUTO NAVIGATION when set, a navigation panel is placed at the bottom of long pages,

unless overridden by $NO NAVIGATION being set also . . . . . . . . . . . . . . . . . . . 65, 71 $AUTO PREFIX if set, all files produced have a filename-prefix using the name of the L^{A}TEX

file being processed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 52, 59 $BODYTEXT string to be used within the <BODY...> tag on HTML pages . . . . . . . . . 71 $BOTTOM NAVIGATION when set, a navigation panel is placed at the bottom of every HTML

page, unless overridden by $NO NAVIGATION being set also . . . . . . . . . . . . . . . . . 65 $CHARSET specifies the character set used within the HTML pages, currently this defaults

to "iso 8859 1" (ISO–Latin–1) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 $CHILDLINE has piece of HTML code to be inserted between the child-links table and page

contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 $CONTENTS when navigation panels are created, this holds a hyperlink to the HTML file

that has the Table-of-Contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 $CONTENTS IN NAVIGATION when set, an iconic link to the Table-of-Contents is placed in

navigation panels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 $CRAYOLAFILE when color.perl has been loaded, holds the name of the file which con-

tains named color specifications for crayon colors . . . . . . . . . . . . . . . . . . . . . . . 70 $CUSTOM BUTTONS variable available to link to customised buttons in navigation panels 76 $CUSTOM TITLES when set, a customised subroutine is used to determine the filename for

each HTML page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 $DEBUG when non-zero, L^{A}TEX2HTML shows messages that would otherwise be ignored 67 $DESTDIR specifies the directory into which all the generated files will be placed . . 58 $DISCARD PS when set, specifies that the temporary PostScript files created by dvips be

discarded immediately after use . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 $DISP SCALE FACTOR extra magnification for images of displayed math environments; mul-

tiplies the $MATH SCALE FACTOR to give the total scaling . . . . . . . . . . . . 16, 18, 61, 71

$DONT INCLUDE holds a list of filenames and filename extensions to not include, even if

requested; filenames ending with .sty are automatically excluded . . . . . . . . 69, 77 $DO INCLUDE holds a list of filenames and filename extensions to be included, even when

the $DONT INCLUDE variable would normally exclude them . . . . . . . . . . . . . . . . . 69 $DVIPS contains the command to be used when the dvips program processes images.dvi

to create PostScript files describing the required images . . . . . . . . . . . . . . . . . . 68 $DVIPS MODE contains the Metafont mode to be used when the dvips program supports

automatic font-generation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 61, 69 $EXTERNAL BIBLIO specifies the URL for a bibliography page external to the current

document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 $EXTERNAL CONTENTS specifies the URL for the CONTENTS button in the navigation

panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 $EXTERNAL DOWN LINK specifies the URL for the NEXT button in the navigation panel 66 $EXTERNAL DOWN TITLE specifies the text for the NEXT button hyperlink in the navigation

panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 $EXTERNAL FILE prefix of a .aux file to be read for numbering and other information

provided by L^{A}TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 $EXTERNAL IMAGES when set, hyperlinks to the generated images are included on the HTML

pages, not <IMG> tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 $EXTERNAL INDEX specifies the URL for the INDEX button in the navigation panel 66 $EXTERNAL PREV LINK specifies the URL for the PREVIOUS button in the navigation

panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 $EXTERNAL PREV TITLE specifies the text for the PREVIOUS button hyperlink in the

navigation panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 $EXTERNAL UP LINK specifies the URL for the UP button in the navigation panel . . 66 $EXTERNAL UP TITLE specifies the text for the UP button hyperlink in the navigation

panel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 $EXTN holds the filname suffix for the HTML files to be produced, either .html or .htm 70 $EXTRA IMAGE SCALE when set, specifies that images should be created at one size but

displayed at another; it provides the ratio between these sizes . . . . . . . . . 18, 20, 71 $FIGURE SCALE FACTOR magnification factor to be used with images of whole figure and

table environments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18, 19, 21, 61, 71 $FILE holds the name of the HTML output of the top-most page in a document, when vital

links are to be established . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62, 70 $FOILCLASS The document class that is substituted for foils in the images.tex file

when using FoilTEX; default is 'article' . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 $FOILHEADLEVEL holds the sectioning level for the pages containing FoilTEX ‘foils’ gener-

ated using foilhtml.sty and foils.perl . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 $FOILOPTIONS Options for images.tex, obtained from processing a FoilTEX source; by

default ’10pt’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 $FONT SIZE gives font size used with images of math and other environments …. 61 $GS contains the command to call the Ghostscript program, used by pstoimg …. 72 $GS DEVICE contains the name of the device driver for the Ghostscript program, used by

pstoimg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 $GS LIB contains the directory path prefix to the Ghostscript program’s library of special

input files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 $HOME/.latex2html-init user’s default initialisation file for L^{A}TEX2HTML; if present, it

will be read unless another has already been found . . . . . . . . . . . . 9, 62, 73, 75, 77 $HOWTO INTERLACE GIF contains the name of the graphics utility package to use for cre-

ating interlaced GIF images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 $HOWTO TRANSPARENT GIF contains the name of the graphics utility package to use for

creating transparent GIF images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 $HTML VERSION specifies the version of HTML code to be generated . . . . . . . . . . . . . 60 $ICONSERVER contains the URL where icon-images are to be found . . . . . 8, 10, 62, 69 $IMAGES ONLY if set, this run simply processes the images.tex file generated on an earlier

run . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 $IMAGE TYPE specifies the graphics format for the images that may be generated; currently

either svg, png or gif . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 $INDEX when navigation panels are created, this holds a hyperlink to the HTML file that

has the Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 $INDEX IN NAVIGATION when set, an iconic link to the Index is placed in navigation

panels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 $INFO when set, contains Perl code to produce a replacement for the standard “About

this document” information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 63 $INTERLACE if set, this specifies that any generated images should be interlaced … 71 $LANGUAGE TITLES contains the titles for chapters, sections, etc. in the chosen language

(default: English) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 $LATEX holds the command to be used to invoke L^{A}TEX on the images.tex file, for type-

setting the contents of images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 $LATEX2HTMLDIR directory where L^{A}TEX2HTML is installed . . . . . . . . . . . . 8, 62, 68, 77 $LATEX2HTMLDIR/latex2html.config configuration-file for the local setup; this file is

read before any .latex2html-init files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73, 75

$LATEX2HTMLDIR/local.pm created by configure-pstoimg, this file holds information

about what graphics utilities are available on the local system . . . . . . . . . . . . . . 72 $LATEX2HTMLDIR/styles/german.perl adds Perl code to establish compatibility with

the german.sty L^{A}TEX 2ε package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 $LATEX2HTMLSTYLES contains the directory path to find the styles/ directory; usually a

subdirectory of $LATEX2HTMLDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68, 70 $LATEX2HTMLVERSIONS contains the directory path to find the versions/ directory; usu-

ally a subdirectory of $LATEX2HTMLDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 $LATEX DUMP if set, speeds up image processing during subsequent runs on your document

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 $LINE WIDTH holds the width of displayed equations, used for right-justification … 71 $LINKNAME name for the duplicate HTML file which enables automatic browser-linking from

a directory name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 70 $LINKPOINT specifies the name for the HTML file to be duplicated to enable automatic

browser-linking from a directory name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 70 $LOCAL ICONS when non-zero, causes a copy of the icons used to be placed within the

directory along with the HTML files and generated images . . . . . . . . . . . . . . . . 10, 62 $LONG TITLES when non-zero, specifies how many words from the section-title are used

to determine the filename for each HTML page . . . . . . . . . . . . . . . . . . . . . . . . . . 58 $MATH SCALE FACTOR magnification factor for images of inlined equations and displayed

math environments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16, 18, 19, 61, 63, 71 $MAX LINK DEPTH number of depth levels for which child-links are displayed, within the

mini table-of-contents on each page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 $MAX SPLIT DEPTH sectioning-level at (and below) which creating a new HTML page ceases

to occur . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 75 $METAFONT DPI dots per inch required when Metafont is called to generate a font . 69 $NAV BORDER when set, holds the width (in pt) of a border for navigation buttons . 70 $NETSCAPE HTML when set, will allow HTML extensions understood by Netscape to be used;

e.g. tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 $NEXT when navigation panels are created, this holds a hyperlink to the next HTML page 75 $NEXT GROUP when navigation panels are created, this holds a hyperlink to the next HTML

page at the same level or shallower . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 $NEXT GROUP TITLE when navigation panels are created, this holds the title of the next

HTML page at the same section level or shallower . . . . . . . . . . . . . . . . . . . . . . . . 75 $NEXT PAGE IN NAVIGATION when set, an iconic link to the next logical page is placed in

navigation panels, unless overridden by $NO NAVIGATION being set also . . . . . . . 66 $NEXT TITLE when navigation panels are created, this holds the title of the next HTML

page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 $NOFORK when set, prevent the translator to crash if your operating system does not

support forking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 $NOLATEX when set, no image generation is performed but a text-string is placed; this is

no longer needed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 $NO AUTO LINK if set, the second copy of the HTML file specified in $LINKPOINT is not

created, with the name given in $LINKNAME; this allows a URL to a directory name to reach the main HTML page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 70

$NO FOOTNODE if set, footnotes will appear at the bottom of the current HTML page, rather

than on a separate page for footnotes only . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 $NO IMAGES if set, the images.tex file is generated but not processed to create images 64 $NO NAVIGATION when set, no navigation panels are placed on HTML pages . . . . . . . 65 $NO SIMPLE MATH when set, inhibits the use of “simple math”, unless the ,math exten-

sion is loaded; in which case TEXlike parsing is used, creating images of sub-parts of expressions, as needed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15, 61

$NO SUBDIR when set, the HTML files and images will be placed in the current directory,

rather than a sub-directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 $NUMBERED FOOTNOTES numbers your footnotes when defined . . . . . . . . . . . . . . . . 62 $PAPERSIZE controls the maximum size for large images constructed by L^{A}TEX, on a single

page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71, 81 $PBMMAKE contains the command to call the pbmmake utility, used by pstoimg …. 72 $PBMPLUSDIR when set, contains where to find the pbmplus graphics utilities; (no longer

needed) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 $PK GENERATION when set, automatic font-generation is enabled for low-res bitmapped

fonts; not needed with anti-aliasing of hi-res bitmaps, or PostScript outlines 9, 61, 69 $PNMCAT contains the command to call the pnmcat utility, used by pstoimg . . . . . . 72 $PNMCROP contains the command to call the pnmcrop utility, used by pstoimg …. 72 $PNMFILE contains the command to call the pnmfile utility, used by pstoimg …. 72 $PNMFLIP contains the command to call the pnmflip utility, used by pstoimg …. 72 $PPMQUANT contains the command to call the ppmquant utility, used by pstoimg . . 72 $PPMTOGIF contains the command to call the ppmtogif utility, used by pstoimg . . 72 $PREFIX contains the filename-prefix to prepend to all HTML, image and data files created

for the current document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 $PREVIOUS when navigation panels are created, this holds a hyperlink to the previous

HTML page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 $PREVIOUS GROUP when navigation panels are created, this holds a hyperlink to the HTML

page starting the previous section at the same level or shallower . . . . . . . . . . . . 75 $PREVIOUS GROUP TITLE when navigation panels are created, this holds the title of the

previous HTML page at the same section level or shallower . . . . . . . . . . . . . . . . . 75 $PREVIOUS PAGE IN NAVIGATION when set, an iconic link to the previous logical page is

placed in navigation panels, unless overridden by $NO NAVIGATION being set also 66 $PREVIOUS TITLE when navigation panels are created, this holds the title of the previous

HTML page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 $PSTOIMG contains the command to run the pstoimg utility to convert PostScript files

into images in the required graphics format . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 $PS IMAGES when set, hyperlinks to PostScript files are included on the HTML pages, not

<IMG> tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 $REUSE settings affect whether images constructed on a previous run of L^{A}TEX2HTML are

re-used on the next run of the same document . . . . . . . . . . . . . . . . . . . . . . . . . 64 $RGBCOLORFILE when color.perl has been loaded, holds the name of the file which

contains color specifications and names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 $SCALABLE FONTS used when scalable fonts (e.g. PostScript) are available for use with

image-generation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 $SERVER used in conjunction with image-maps to specify the type of web-server which

will interpret the map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 $SHORTEXTN when set, uses .htm as the filename suffix, rather than .html … 58, 70 $SHORT INDEX when set, results in a more compact Index being produced . . . . . 30, 62 $SHOW SECTION NUMBERS when set, L^{A}TEX’s section numbering is included with section

titles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 $TEXDEFS when set, the texdefs.perl file is read, allowing some TEX constructions to

be interpreted . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 $TEXEXPAND contains the command to run the texexpand utility which combines all re-

quired portions of input files for a L^{A}TEX2HTML document . . . . . . . . . . . . . . . . . 68 $TEXINPUTS contains the paths for L^{A}TEX to search for styles and input files, when making

images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 69 $TITLE initially can be set to hold a title for the document, as a whole . . . . . . . . . 58 $TITLES LANGUAGE specifies the collection of strings to use as titles, for special document

sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 $TMP when set in local.pm, used by pstoimg as a temporary directory for image-

generation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 $TOC DEPTH sectioning-level down to which links occur on the Table-of-Contents page 60 $TOC STARS when set, star-sections are listed/linked on the Table-of-Contents page 60 $TOP NAVIGATION when set, a navigation panel is placed at the top of every HTML page,

unless overridden by $NO NAVIGATION being set also . . . . . . . . . . . . . . . . . . . . . 65 $TRANSPARENT FIGURES when set, this specifies that the background of generated images

should be made transparent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 $UNSEGMENT when set, L^{A}TEX2HTML translates a segmented document as a whole rather

than a single segment; use this for testing your segmented document . . . . . . . . . 60 $UP when navigation panels are created, this holds a hyperlink to the parent HTML page 75 $UP TITLE when navigation panels are created, this holds a title of the parent HTML

page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 $USENETPBM when set, indicates that the netpbm graphics utilities are available; (no longer

needed) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 $VERBOSITY has a numerical value which determines the messages to be printed on-screen

as L^{A}TEX2HTML processes a document . . . . . . . . . . . . . . . . . . . . . 57, 67, 68, 77, 79 $WHITE BACKGROUND when set, ensures that images of figure environments have a white

background . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 65 $WORDS IN NAVIGATION PANEL TITLES used to limit the number of words shown with the

navigation-panel buttons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71, 75 $WORDS IN PAGE contains the number of words required before a navigation-panel is placed

also at the bottom of a page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65, 71 $cb version when using the changebar package, this holds a version identifier set by the

\cbversion command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 $dd “directory delimiter”, set automatically to ‘/’ in Override.pm for Unix systems, and

to whatever else is appropriate on other platforms . . . . . . . . . . . . . . . . . . . . . . 68 $default language specifies the language whose international code should be specified

within the <!DOCTYPE...> tag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 $depth holds the value of the sectioning depth when the <ADDRESS> tag of an HTML page

is being created . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 $file holds the name of the HTML page when its <ADDRESS> tag is being created . . 62 $title holds the title of the current section when the <ADDRESS> tag of an HTML is being

created . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

ActivePerl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Aladdin GhostScript . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 amsfonts.perl contains code to recognise or implement features of the amsfonts package

for L^{A}TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 amstex.perl contains code to recognise or implement features of the amstex package for

L^{A}TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 Arena Web-browser^{20}; currently the only one which displays HTML 3.1 math markup 12 article.perl implements section-numbering conventions, etc. for the article document-

class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

bibtex bibliography data-base program, used with L^{A}TEX . . . . . . . . . . . . . . . . . . . 81 book.perl implements section-numbering conventions, etc. for the book document-class

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

CERN originally named “Conseil Europeen pour la Recherche Nucleaire”, now known as the “European Laboratory for Particle Physics”^{21}, located near Geneva; also where the World-Wide Web^{22} was originally conceived . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

changebar.sty L^{A}TEX 2ε package to mark changes in documentation; this supercedes an

earlier package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 28, 85 Changes text file, recording the changes made to L^{A}TEX2HTML since it was first released in

July 1993, until V96.1 and revisions during 1996 . . . . . . . . . . . . . . . . . . . . . . 7, 11 chmod Unix command for changing the access-privileges on files . . . . . . . . . . . . . . 77 color.perl contains Perl code to implement all the features of L^{A}TEX’s color package and

loads a list of named colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49, 70 configure-pstoimg Perl script to locate which Unix utilities are available on the local

system, for use in generating images . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 8, 72, 86 credits.tex input file for the L^{A}TEX2HTML manual; acknowledges the multitude of con-

tributors to the development of L^{A}TEX2HTML . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

DBM Unix Database Management utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 79 DCOM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 dn.gif image of a small downward-turning arrow . . . . . . . . . . . . . . . . . . . . . . . . 10 docs/ subdirectory of $LATEX2HTMLDIR containing the L^{A}TEX2HTML documentation and a

Makefile to control its processing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 10 docs/foilhtml/ subdirectory of $LATEX2HTMLDIR containing an implementation for the

FoilTEX system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 28 dot.latex2html-init example initialisation file for L^{A}TEX2HTML; usually to be found

in the $LATEX2HTMLDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 9 dvips dvi to PostScript conversion program . . . . . . . . . 4, 5, 9, 55, 58, 68, 69, 78, 81 dvipsk a variant of dvips . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

example/ subdirectory of $LATEX2HTMLDIR containing an example illustrating the document segmentation technique . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 features.tex input file for a section of the L^{A}TEX2HTML manual; describes special features

20http://www.w3.org/hypertext/WWW/Arena

21http://www.cern.ch

22http://www.w3.org/hypertext/WWW/TheProject.html

available with L^{A}TEX2HTML . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 floatfig.sty L^{A}TEX 2ε package for floating figures with text flowing around . . . . . 27 foilhtml.sty a L^{A}TEX package that adds structural markup to FoilTEX . . . . . . . . 28 foils.perl adds Perl code to implement the foils document-class of FoilTEX . . 25, 28 french.perl adds Perl code to be compatible with the french.sty L^{A}TEX package 26

GDBM GNU Database Management utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 79 german.perl adds Perl code to be compatible with german.sty L^{A}TEX package … 26 german.sty L^{A}TEX package for documents written in German . . . . . . . . . . . . . . . . 25 GhostScript . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Ghostscript program to render graphics written in the PostScript language, or minor

variations thereof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 5, 16, 87, 91 giftool utility to make transparent the backgrounds of GIF images, or to create inter-

laced GIF images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 giftrans utility to make transparent the backgrounds of GIF images; not needed if

netpbm is used . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 grep standard Unix utility, useful for finding specific strings within (groups of) files 79 gs commonly-used shortening of the ghostscript command . . . . . . . . . . . . . . . 5, 9 GS-Preview dvi-viewer, using Ghostscript . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 gs fonts.ps read by Ghostscript to locate the PostScript fonts available on the local

system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 GSWIN32C.EXE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

heqn.sty empty L^{A}TEX package to force loading of heqn.perl; moves equation numbering

to the left of images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 HTML “Hyper-Text Markup Language”, as used by Web-browsers . . . . . . . . i, 7, 8, 89 html.perl adds Perl code to be compatible with html.sty L^{A}TEX package …. 33, 48 html.sty symbolic link (see below) to a L^{A}TEX package allowing HTML links and conditional

code . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1, 5, 8, 10, 18, 23, 32–34, 44, 48, 81 htmllist.perl contains Perl code to implement the htmllist environment . . . . . . . 33 htmllist.sty L^{A}TEX package for the htmllist environment, which uses special icons in the

HTML version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 27, 33 hypextra.tex input file for a section of the L^{A}TEX2HTML manual; describes hypertext

extensions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

icons.gif/ subdirectory of $LATEX2HTMLDIR containing GIF format versions of the de-

fault icons used for navigation buttons and other purposes . . . . . . . . . . . . . . . . . 7 icons.png/ subdirectory of $LATEX2HTMLDIR containing PNG format versions of the

default icons used for navigation buttons and other purposes . . . . . . . . . . . . . . . . 7 icons/ subdirectory of $LATEX2HTMLDIR containing the default GIF icons used for navi-

gation buttons (no longer used, from v97.1 onwards) . . . . . . . . . . . . . . . . . . . . . 10 images.log the log-file produced by running L^{A}TEX on images.tex . . . . . . . . . 81, 82 images.pl data file produced by L^{A}TEX2HTML; contains information about the images in

your document, used for cyclic image generation . . . . . . . . . . . . . . . . . . . . . . . . 82 images.tex file created by L^{A}TEX2HTML which is passed to L^{A}TEX to typeset complicated

environments for conversion to images . . . . . . . . . . . . . . . . . . 28, 33, 43, 65, 82, 83 index.html on many systems a file of this name will be sent to the browser when a URL

to its parent directory has been specified . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 IndicTeX-HTML/ subdirectory of $LATEX2HTMLDIR containing L^{A}TEX and Perl extensions

to handle IndicTeX documents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 install-test tests the L^{A}TEX2HTML installation . . . . . . . . . . . . 4, 7, 8, 68, 69, 77, 79

justify.sty L^{A}TEX package which does nothing other than force loading of justify.perl by L^{A}TEX2HTML . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

l2hfiles.dat contains the descriptive data appearing in this glossary …. 10, 31, 32 l2hglo.ist configuration for makeindex to construct this glossary . . . . . . . . . . 10, 32 l2hidx.ist configuration for makeindex to construct the index . . . . . . . . . . . . . . 10 l2hman.sty style-file needed for typesetting the L^{A}TEX2HTML manual using L^{A}TEX . 10 labels.pl data file containing symbolic labels for references to locations within a HTML

document produced by L^{A}TEX2HTML . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46–48, 51 latex command to run to L^{A}TEX . . . . . . . . . . . . . . . . . . . . . . . 4, 9, 10, 55, 69, 81, 82 L^{A}TEX the markup language, interpreted by L^{A}TEX2HTML; images are processed by the L^{A}TEX

program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . i L^{A}TEX2ε the current version of L^{A}TEX, first introduced in 1993 . . . . . . . . . . . . . . . . . . 1 latex2html Unix command to execute the L^{A}TEX2HTML script for generating HTML pages

from a L^{A}TEX document . . . . . . . . . . . . . . . . . . . . 7, 8, 33, 55, 58, 67, 68, 73, 77–79 L^{A}TEX2HTML Perl script to convert L^{A}TEX documents into web-pages, by Nikos Drakos with

updates and extensions by others . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 latex2html-2019.tar.gz . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 latex2html.config configuration file for L^{A}TEX2HTML, normally to be found in the

$LATEX2HTMLDIR directory . . . . . . . . . . . . . . . . . 7–10, 13, 16, 59, 60, 62, 68, 69, 81 letter.perl implements section-numbering conventions, etc. for the letter document-

class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 libpng-0.89c.tar.gz implements the PNG image format routines, available via Archie 5 libz implements data compression routines, should be part of your operating system, if

not get it via Archie . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 licence.tex input file for the L^{A}TEX2HTML manual; contains the licence statement and

copyright notice . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 local.pm written by configure-pstoimg, this Perl file records the locations of the graph-

ics utilities on the local system used by pstoimg . . . . . . . . . . . . . . . . . . . . . . . . . 7, 8

lslide.sty L^{A}TEX package for producing colour-separations . . . . . . . . . . . . . . . . . 83 lynx character-based Web-browser, available for DOS and Unix systems . . . . . . 5, 64

make Unix utility used to track the time-dependency of multiple files in the development

of a document or computing project . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55, 58 Makefile contains the time-dependency data used by the make utility . 55, 58, 66, 67 makeidx.perl adds Perl code to interpret index-entries otherwise intended for the

makeindex program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30, 84 makeindex utility for sorting and filtering textual data; used primarily for making the

index or glossary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 31 makemap Perl script to generate image-maps . . . . . . . . . . . . . . . . . . . . . . . . . 7–9, 50 makeseg Perl script to generate Makefiles for segmented documents, so that you can

initiate the translation with a single make call. . . . . . . . . . . . . . . . . . . . . . . . 7, 85 MakeTeXPK script used by dvips, and other dvi-drivers, controlling generation of missing

fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9, 69 manhtml.tex input file containing definitions needed for segments of the HTML version of

the L^{A}TEX2HTML manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 manual.aux auxiliary file created by running L^{A}TEX on the L^{A}TEX2HTML manual … 10 manual.glo collects (unsorted) glossary entries for the L^{A}TEX version of the L^{A}TEX2HTML

manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30, 31 manual.gls file containing the sorted entries for the glossary in the L^{A}TEX2HTML manual;

created by running makeindex on manual.glo using l2hglo.ist . . . . . . . . . 10, 32 manual.ind file containing the sorted entries for the index in the L^{A}TEX2HTML manual;

created by running makeindex on manual.idx using l2hidx.ist . . . . . . . . . . . 10 manual.tex main L^{A}TEX source for the L^{A}TEX2HTML manual . . . . . . . . . . . . . . . . . 10 Metafont generates bitmapped fonts for TEX and L^{A}TEX . . . . . . . . . . . . . . . . . . . 69 misc/ subdirectory containing miscellaneous files used by Metafont, often the location for

modes.mf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 modes.mf used by Metafont to customise font bitmaps for different printers …. 9, 69 Mosaic commonly-used Web-browser, see NCSA Mosaic below . . . . . . . . . . . . . . 83 MS IE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 mybutton.gif file-name used in example code for an image as a button to a hyperlink 74

natbib.perl implements the commands in the natbib package, allowing flexible styles for

citations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 NCSA “National Center for Supercomputing Applications” located at University of Illinois

at Urbana–Champaign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 NCSA Mosaic Web-browser, freely available for most platforms . . . . . . . . . . . . . . . 56 NDBM alternative Database Management utility for Unix . . . . . . . . . . . . . . . . . . 4, 79 NetPBM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 netpbm suite of Unix graphics utilities^{23}; alternative to pbmplus . . . . . . . . . . 4, 5, 91 Netscape Navigator popular Web-browser, commercially available^{24} for most platforms

and free to academic institutions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 56

Override.pm Perl input file for operating system specific constructs like directory delim-

iter, getenv()/setenv(), and other system calls and quirks . . . . . . . . . . . . . . . . . 89 overview.tex input file for a section of the L^{A}TEX2HTML manual; contains a brief summary

of L^{A}TEX2HTML’s capabilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

pbmplus suite of Unix graphics utilities^{25} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 pdftocairo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Perl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Perl the Unix program^{26} which interprets L^{A}TEX2HTML and orchestrates the whole con-

version process . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . i PNG Portable Network Graphics format, devised as an alternative to GIF with extra

capabilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 86 pnmcrop Unix graphics utility which crops whole rows and/or columns of a single color,

from specified edges of a bitmapped graphic . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 5 pnmflip Unix graphics utility which rotates and reflects through multiples of 90^{∘}. . 19 pnmtopng Unix graphics utility which converts to PNG format. . . . . . . . . . . . . . . . . . 5 PostScript graphics programming language used in printers; interpreted by Ghostscript

. . . . . . . . . . . . . . . 5, 9, 10, 18, 57, 61, 64, 65, 69, 78, 80, 81, 84 problems.tex input file for the L^{A}TEX2HTML manual; contains a description of known

problems with L^{A}TEX2HTML, and how best to avoid or deal with them . . . . . . . . . 10 psfiles/ sub-directory of docs/ containing PostScript graphics used in the printed ver-

23ftp://ftp.x.org/contrib/utilities/

24http://home.netscape.com

25ftp://ftp.x.org/contrib/utilities/

26ftp://ftp.uu.net/languages/perl/

sion of the L^{A}TEX2HTML manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 pstogif Perl script which controls the conversion of images from PostScript to GIF

bitmaps; replaced by pstoimg for V97.1 . . . . . . . . . . . . . . . . . . . . . . . . 7, 8, 84, 86 pstoimg Perl script which controls the conversion of images from PostScript to a chosen

graphics format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 8, 68, 72, 77, 86 pstoppm.ps contains PostScript code needed for image conversion with versions of

Ghostscript earlier than 3.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

README found in $LATEX2HTMLDIR, this file contains simple information concerning the

latest release . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 report.map example data-file used as input for makemap to produce an image-map 51 report.perl implements section-numbering conventions, etc. for the report document-

class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 report.ps PostScript version produced by L^{A}TEX, for the report.tex example of a seg-

mented document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 report.tex example source for a segmented document . . . . . . . . . . . . . . . . . . 53–55 report/ subdirectory created when processing the example segmented document . 55

SDBM Unix Database Management utility distributed with Perl 5 . . . . . . . . . . . . . 79 sec1.tex one section of the example segmented document . . . . . . . . . . . . . . . . 53–55 sec2.html one HTML page produced with the example segmented document . . . . . 55 sec2.tex another section of the example segmented document . . . . . . . . . . . . . 53, 55 slides.perl implements section-numbering conventions, etc. for the slides document-

class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 styles/ subdirectory of $LATEX2HTMLDIR containing <file>.perl files corresponding to

L^{A}TEX packages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 13, 24, 25, 68 support.tex input file for a section of the L^{A}TEX2HTML manual; details installation and

some technical aspects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

tar utility for creating, and extracting files from, archives . . . . . . . . . . . . . . . . . . . 6 tests/ subdirectory of $LATEX2HTMLDIR with various test documents, in particular you

might be interested in running pages.tex to check the image generation . . . . . . . 7 texdefs.perl package providing ability to recognise and process some common TEX

macro constructions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 texexpand Perl script used with L^{A}TEX2HTML for combining the required portions of all

input-files into a single document . . . . . . . . . . . . . . . . . . . . . . . 7, 8, 68, 77, 84, 86 TEXINPUTS shell-variable which contains the usual paths used by L^{A}TEX to find packages

and other input-files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 9, 80, 81 texinputs/ subdirectory of $LATEX2HTMLDIR containing the L^{A}TEX packages needed to

successfully typeset its manual . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 8, 28 toshiba Metafont mode recommended for setting mathematics and text included within

automatically generated images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

up.gif image of a small upward-turning arrow . . . . . . . . . . . . . . . . . . . . . . . . . . 10 url.sty L^{A}TEX 2ε package for typesetting URLs across multiple lines . . . . . . . . 10, 85 urls.bib suggested name for a BibTEX input file, containing @string definitions for

URLs to be used within bibliographic references . . . . . . . . . . . . . . . . . . . . . . . . 45 userman.tex input file for the main section of the L^{A}TEX2HTML user-manual . . . . . 10

versions/ subdirectory of $LATEX2HTMLDIR containing files to adapt to different versions of HTML . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 60 WinZip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 wrapfig.sty L^{A}TEX 2ε package for wrapping text around non-floating figures …. 27

xv useful Unix graphics utility; allows interactive manipulation and format-conversion of bitmapped images . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

Index

Symbols

TEX Users Group (TUG), 11

A

About this document..., 62 accents, 3

foreign, 84 address

signature, 62 using a subroutine, 62

alltt environment, 26 AMS-L^{A}TEX

amsfonts package, 26 amsmath package, 26 amssymb package, 26 amstex package, 26

AMS-style environments listed, 35 old, use discouraged, 35

anchor tag

NAME attribute, 36 anti-aliasing, 20

blurred printing, 65 comparison, 16 in figures, 64 in math and text, 64 override defaults, 19 requires Ghostscript 4.03, 16 turned off, 65

arbitrary environments, 2, 22 Arena

math mark-up, 12 Arseneau, Donald, 85 ascii-mode, 63 auxiliary file, 57

which to read, 61

B

babel package, 26 backdrop

tiled with an image, 3

%begin{latexonly}, 42

Belinfante, Axel, 84 bibliography, 3, 45–46, 57

bibliographic database, 45 example using URLs, 45 natbib package, 45 @string commands, 45 @string commands, 45 using .bib file, 45

\bodytext, 49

no checking for valid attributes, 49 Bohnet, Achim, 89 Braams, Johannes, 85 browser

Arena, 12 character-based, 3, 5, 63 lynx, 5 NCSA Mosaic, 56 Netscape Navigator, 56

browser, 56

supports images, 3 bugs, 76

bug reports, 11

C

Cailliau, Robert, 84 Carroll, Eric, 84 CERN, 84

Arena, 12 image-map server, 9, 51 Michel Goossens, 85 World-Wide Web Project, 84, 85

change-bars, 28, 85

different versions, 28 changebar package, 26 character

tilde in URLs, 37 character set, 70

dtd isolanguage, 62 ISO–10646 (Unicode), 3, 12 ISO–8879 (ISO–Latin–1), 3, 12

"iso 8859 1", 70

character set, 3 chemsym package, 26 child-links table, 59

at bottom of page, 59 at top of page, 59

tableofchildlinks, 59

citations harvard package, 26 Harvard style

handled by natbib, 45 natbib package, 26, 45 nharvard package, 26 within external bibliographies, 47

class

foils, 26 class-option

loads file, 25 specific file, 25

color colordvi package, 26 coloured background, 3 coloured text, 3

color package, 26 columns

multicol package, 26 comment environment, 38 Common L^{A}TEX Commands

L^{A}TEX blue book, 3 Companion, see The L^{A}TEX Companion components

specify depth, 2, 59 Computer Based Learning Unit

Ana Maria Paiva, 84 Andrew Cole, 84 CBLU, 84 Jamil Sawar, 84 Roderick Williams, 84

Computer Science Education Project

CSEP, 84 Computer Based Learning Unit

University of Leeds, 84 Computer Science Education Project

CSEP, 21 conditional text, 3, 43

avoid using counters, 42 example, 42 HTML or L^{A}TEX version, 33 scoped variant, 41 shorthand notation, 42 without scope, 42

contents

from other segments, 52 copyright

Leeds, 91 Macquarie, 91

counter automatic, 3 dependency, 3

cross-references, see references current version of L^{A}TEX2HTML, 67 customised layout, 56

D

Daly, Patrick, 45, 85

DANTE, 87 Darmstadt

Fachbereich Informatik, 87 debugging, 67, 77

Perl debugger, 67 depth

child nodes revealed, 59 diagnostic information, 67 directory delimiter, 68 discussion group, 11 document

illustrative examples, 3 segments, 52 title, 58 document class, 24 loads file, 25 loads a .perl file, 24 options, 24

document segmentation, 57 documentation

auxiliaries, 10

Changes section, 11 .dvi version, 10 filename data, 10 graphics, 10 images, 10 index and glossary, 10 input files, 10 requirements, 10 style-files, 10 test of installation, 10

                                       using makeindex, 10

without makeindex, 10 draft mode, 64 Drakos, Nikos, 1, 87, 89, 92 dvips version, 4

E

electronic forms, 37

                                   %end{latexonly}, 42

enumerate package, 26

environment equation, 21 alltt, 26 arbitrary, 22 comment, 38 eqnarray, 21 figure, 21 floatingfigure, 27 hthtml, 26 htmllist, 26, 27 htmlonly, 41 imagesonly, 42 latexonly, 41 makeimage, 43 minipage, 23 rawhtml, 37 table, 23 table, 23 theorem-like, 3 wrapfigure, 27

environments, 11

old AMS-style, use discouraged, 35 epsbox package, 26 epsfig package, 26 eqnarray environment, 17, 21

example, 15 equation environment, 17 equations

alignment, 2, 19 array, 15, 21 inlined, 2 numbered, 2, 15 right-justified, 2, 15

error checking missing braces, 57 unmatched brace, 57 using L^{A}TEX, 57

extended markup

XML, 90 extension

options, 13 extensions, 60–76

definitions, 76 disabled, 60 examples, 25 hypertext, 32 in initialisation file, 72 in style-files, 3 languages, 12 new environments, 3 new theorems, 3 require understanding of Perl, 27 TEXdefinitions, 60

extensions

new commands, 3 external references, 46–47

\externalcite, 47

                                   \externallabels, 46
                                   \externalref, 46, 47

extra scaling of images, 18

F

Fachbereich Informatik, Darmstadt, 87 features

listing, 2 figure environment, 21 figures, 2, 21

arbitrarily scaled, 2, 19

captions, 21 fine control, 18 from other segments, 52 oriented, 2, 19 wrapfig package, 26

file suffix

                                       .tex or.doc, 56

optional, 56 filenames

.htm extension, 58 auto-prefix, 59 automatic link to directory name, 59

                                       custom title hook, 58

customised, 58

                                       index.html, 59

long names, 58

prefix, 58 finnish

finnish package, 26 fixes, 77 floatfig package, 26 floatingfigure environment, 27 floatflt package, 26 foils class, 26, 27 FoilTEX, 26, 27 font size

for image generation, 61

scalable, 61 font-size

in environments, 61 magnification, 61

fonts

non-standard encodings, 90 footnotes, 3

marker style, 62 numbering, 62 on same page, 62 on separate page, 62

forking

prevent, 62 Foster, Ian, 11, 84 frames package, 26 french

francais package, 26 french package, 26

G

german

germanb package, 26 german package, 26 getting started, 56 Ghostscript version, 4 GIF, see images

image conversion, 5 Glossary

HTML version, 30 integrated with Index, 30 printed version, 30

Goossens, Michel, 1, 85, 87 graphics

graphicx package, 26 XY-pic package, 26

graphics package, 26 Graphics Companion, see The L^{A}TEX

Graphics Companion

H

Hennecke, Marcus, 84, 86, 89 heqn package, 17, 26 heqn.sty style-file, 17 hthtml environment, 26

HTML

an evolving standard, 12 arbitrary markup, 37 basic version, 2.0 , 12 current version, 3.2 , 12 future version, 4.0 , 12 image-maps, 12 interactive forms, 3, 12 new HTML tags, 37 non-standard extensions, 60 raw HTML commands, 3, 37 SGML-like markup, 37 Version 2.0, 12 Version 2.1, 3 Version 2.1, now ‘i18n’ extension, 12 Version 3.0, 3, 19 Version 3.0, HTML-Math model, 12 Version 3.1, now ‘math’ extension, 12

Version 3.2, 50

\html, 42

html package, 18 HTML version, 60

html.sty

new commands defined, 33 new environments defined, 33

html.sty package

how to load it, 32 html.sty style-file, 10, 18, 32, 44, 48

location, 8

needed for special features, 5

\htmladdimg, 36
\htmladdnormallink, 36
\htmladdnormallinkfoot, 36
\htmladdtonavigation, 74
\htmlbase, 49
\htmlbody, 49
\htmlcite, 46

example, 45

\htmlhead, 53
\htmlimage, 18

overrides configuration, 21

\htmlinfo, 49
\htmlitemmark, 27

htmllist example, 27 item-marks, 27 prints as description, 27

htmllist environment, 26, 27

\htmlnohead, 53
\htmlonly, 41

htmlonly environment, 41

\htmlref, 44

easy to make links, 44 htmlrule

attributes to the <HR> tag, 48

    \htmlrule*, 48

variants, 48

\htmlrule, 48
\HTMLsetenv

alters a Perl variable, 50

\HTMLset

alters a Perl variable, 50 hyper-links, 36

\hypercite, 45

example, 45

\hyperref, 43

\pageref example, 44 hypertext active link, 36 arbitrary references, 36 extensions, 32 URL as footnote, 36

hypertext links, 3 commands to create, 33 Internet services, 3 multi-media resources, 3 news/sound/video, 3

I

icons, 90 alternative set, 62 copied to local directory, 61 customised images, 62

image

backdrop, 3 image conversion

PostScript to GIF, 5 PostScript to PNG, 5

image files not recycled, 64 not reused, 64 recycled, 64 reuse option, 64 shared by default, 64

image generation

using \htmlimage, 16

image-map CERN server, 51 example, 50, 51 map-file, 50 NCSA server, 51 user-map file, 50

image-reuse

interactive, 64 image-type, 63 images

alignment, 19 anti-alias, 19 anti-aliasing, 64 attributes for the <IMG> tag, 36 client-side image-map, 19 disabled, 64 discard PostScript files, 64 display scale-factor, 18 embedded image, 20 equations, 4 external, 19 extra scaling, 18, 20 figure scale-factor, 18 figures, 4, 18 flip option, 19 generated off-line, 64 hypertext link, 19 image-map, 19 image-maps, 21, 50–51 improved print quality, 18 inlined by default, 19 latex dump, 65 links to PostScript, 64 math scale-factor, 18 no anti-aliasing, 65 options, 19 order-insensitive, 21 order-sensitive, 21 recycling, 21 scale, 19 scale-factor, default 1, 18 server-side image-map, 19 sharing, 21 small images, 18, 21 specified width or height, 20 strange accents, 70 tables, 3, 4 thumbnail, 19, 21 transparent, 5, 19 via hypertext links, 64 white background, 65

images

equations, 3 images, 18 imagesonly environment, 42 index, 3, 28–30, 76

alphabetization, 29 blank lines, 29 codified links, 30, 62

for easier browsing, 30 top-most node, 30

cross-link, 28, 29 cross-link incorrect, 29 cumbersome, 29, 53 data from other segments, 52

|emph, 29

hierarchical, 28 integrated with Glossary, 30 labelled entries, 29 page-ranges, 28 quoted delimiters, 29 section-names, 28, 29

|see, 28

short prefixes preferred, 53

|<style>, 29

styled entries, 28

with <prefix>, 30

initialisation

-init file, 62

initialisation file

specified, 62 initialization file

incompatible with early versions, 9 per directory, 9 per user, 9

installation, 4 change configuration, 8 change defaults, 8 check path-names, 8 create manual, 10 environment variable, 8 filename-prefix, 9 fonts-path variable, 9 graphics utilities, 8 icons/ subdirectory, 9 initialization files, 9 input-path variable, 9 LaTeX packages, 8 latex2html, 8 local icons, 10 makemap script, 9 on Windows 95/98/NT, 5 system installation, 8 without Perl shell scripts, 8

installation variables

checked by install-test, 68

inserted by install-test, 68

\internal, 52

internals

labels from other segments, 52 Isani, Sidik, 85 ISO–10646, see character set

bidirectional languages, 12 Unicode, 12

ISO–8859–1

no longer used, 12 ISO–8879, see character set ISO–Latin–1, see character set

J

justify package, 26

L

labels external, 46 internal, 47 symbolic, 46

languages, 12 dtd isolanguage, 62 english, 13 finnsh, 13 french, 13 german, 13 spanish, 13

L^{A}TEX

L^{A}TEX3, 87

                                   \latex, 42

L^{A}TEX blue book, 3, 45 Leslie Lamport, 92

L^{A}TEX commands, 3 embedded, 3 not syntactically legal, 3 L^{A}TEX Companion, see The L^{A}TEX Companion

L^{A}TEX Graphics Companion, see The

L^{A}TEX Graphics Companion L^{A}TEX Web Companion, see The

L^{A}TEX Web Companion L^{A}TEX2HTML

command for its name, 50 latex2html-NG, 86

                                   \latexhtml, 42
                                   \latexonly, 38, 41

latexonly environment, 41 latexsym package, 26

                                   \latextohtml

gives L^{A}TEX2HTML, 50 levels

sectioning, 60 lgrind package, 26 link validation

done by hand, 47 symbolic labels, 47 Lippmann, Jens, 7, 85–87, 89 LiPS Design Team, 87 list of figures, 3, 52 list of tables, 3, 52 Little, Todd, 84 looping endlessly, 77

M

mailing list Argonne National Labs, 11, 84 subscribe, 11 unsubscribe, 11

makeidx package, 26 makeimage environment, 43

empty sub-environment, 43 inside a figure, 43

makemap, 50

manual

short on-line, 56 math, 60, 61 mathematics

markup, MathML, 90 MathML, 12

MathML

mathematics markup, 90 \mbox command with math

generates an image, 17 minipage environment, 23 Moore, Ross, 1, 11, 45, 49, 85–89

N

names of important packages, 50 navigation panel, 74–76

according to number of words, 65 at bottom of page, 65 at top of page, 65 buttons, 75 customisable, 2, 74 customised buttons, 74 disabled, 65 example, 75 iconic links, 75 link to index, 66 link to table-of-contents, 66 next logical page, 65 number of words, 75 optional, 2 order of buttons, 74 Perl subroutines, 75 previous logical page, 66 section titles, 75 textual links, 75

NCSA

image-map server, 9, 51 Nelson, Scott, 89 new commands

arguments remain, 72, 73 command ignored, 73 ignore arguments, 72 order-sensitive, 74

processed in L^{A}TEX, 74 processed in L^{A}TEX, 74 order-sensitive, 74

                                   \nonumber, 18

Noworolski, Mark, 85 numbered equations, 15

O

Ogawa, Art, 11 Olson, Bob, 11, 84 options, 58–68

command-line options, 56 defining temp directory, 67 environment variables, 58 extensions, 60 file-names, 58 print option listing, 67 sectioning, 58 set in initialisation file, 58 special features, 60 titles, 58

options, 58 order-sensitive

setting parameters, 74 wrap-deferred, 74

output default directory, 58 in current directory, 58 redirect to directory, 58

overview, 1

P

package amsfonts, 26 amsmath, 26 amssymb, 26 amstex, 26 babel, 26 changebar, 26 chemsym, 26 color, 26 colordvi, 26 enumerate, 26 epsbox, 26 epsfig, 26 finnish, 26 floatfig, 26 floatflt, 26 frames, 26 francais, 26 french, 26 german, 26 germanb, 26 graphics, 26 graphicx, 26 harvard, 26 heqn, 17, 26 html, 18 justify, 26 latexsym, 26 lgrind, 26 longtable, 26 makeidx, 26 multicol, 26 natbib, 26 nharvard, 26 seminar, 26 spanish, 26 supertabular, 26 texdefs, 26 verbatim, 26 verbatimfiles, 26 wrapfig, 26 xspace, 26 XY-pic, 26

packages, 25

\parbox command

generates an image, 17 Perbandt, Adalbert, 90 PNG

image conversion, 5 Popineau, Fabrice, 1, 5, 86, 89 portability

Unix systems, 86 printing, 57

hyperlink to typeset version, 57 problems, 76–83

R

Ramge, Axel, 87, 89

\rawhtml, 37

rawhtml environment, 37 rawhtml environment

example, 37 references

between documents, 3 bibliographic, 45 collaboration required, 48 document segments, 3 highlighted text, 43 iconic, 43 numeric, 43 symbolic, 3, 46 symbolic, 43 to external documents, 46, 47 without indirection, 43

references

hyper-links, 3 references, 43 requirements, 4–5

DataBase Management system, 4 for full graphics, 4 for segmentation feature, 5 for special features, 5 for transparent images, 5 minimal, 4 without images, 5

Rose, Kristoffer, 87 Rouchal, Marek, 1, 5, 7, 86, 87, 89

S

scale-factors

for math images, 16 sectioning levels, 60 sections

from other segments, 52 in separate files, 59 numbers not shown, 60 show section numbers, 60

\segment, 52

segmentation, 51–56 child segments, 52, 54, 55 child-links, 30 circular dependency, 55 codified index, 30 data about contents, 52 data about figures, 52 data about labels, 52 data about sections, 52 data about tables, 52 data for the index, 52 different sub-directories, 55 document segments, 52 example, 53 internal labels, 55 list of commands, 34 needs L^{A}TEX 2ε, 5 parent segment, 52, 54 relative path, 55 requires book-keeping, 52 same sub-directory, 55 shared references, 52 starting a segment, 53 symbolic links, 51 time-stamps, 55 unsegment, 60

use of Makefile, 55
\segmentcolor, 53
\segmentpagecolor, 53

seminar package, 26 simple math

cancelled, 61 source code, 6

compressed, 6 developer’s repository, 6 for Windows NT platform, 5 listing, 7

spanish package, 26 special, 32

characters, see character set features, 5

special considerations, 11 special features, 11

\startdocument, 53
\strikeout, 48

style sheets CSS, 90 DSSSL, 90

style-files, 25 support, 11

for german language, 25 for specific style-files, 25 mailing list, 11

Swan, Herb, 1 switches, see options symbolic labels, 43–48, 51

see also references, symbolic, 46

T

table environment, 23 table environment, 23 table of contents, 3

depth, 59 star-sections not listed, 60 star-sections shown, 60

\tableofchildlinks, 49

tables, 2, 22 arbitrarily scaled, 2 as HTML mark-up, 3 as images, 3, 4 captions, 21 from other segments, 52 longtablelongtable package, 26 oriented, 2, 19 supertabular package, 26

Taupin, Daniel, 1, 89 temporary files, 58 texdefs package, 26 Thau, Robert S., 7, 84 The L^{A}TEX Companion

Goossens–Mittelbach–Samarin, 92 The L^{A}TEX Graphics Companion

Goossens–Rahtz–Mittelbach, 92 The L^{A}TEX Web Companion

Rahtz–Goossens et al., 92 theorems

theorem-like environments, 3 thumbnail, 2, 19–21, 50

ignores scale-factors, 19 implies external, 19

tilde in URL, 37 time-stamp

used with segmentation, 55 transparent images

override defaults, 19 TUG

TUG: TEX Users Group, 11

U

Uffe Engberg, 47 Umar, Verena, 84 unending loop, 77 Unicode, see character set unrecognised commands, 76 URL

for CONTENTS button, 66 for external Bibliography, 66 for INDEX button, 66 for NEXT button, 66 for PREVIOUS button, 66 for UP button, 66 universal resource locator, 66 url package, 85

use-dvipng, 63 use-luadvi, 63 use-luatex, 63 use-pdftex, 63

                                   usemap, 50

verbatim

verbatimfiles package, 26 verbatim package, 26 verbosity level, 67 Veytsman, Boris, 7, 89 viewer, see browser visual separation

    using \htmlrule, 48

Vojik, Franz, 84

W

Web Companion, see The L^{A}TEX

Web Companion Wilck, Martin, 45, 49, 85 Wortmann, Uli, 1, 87, 89 wrapfigure environment, 27

X

XML

extended markup, 90 xspace package, 26 XY-pic package, 26